### v1

# Representations induced by positive or completely positive sesquilinear maps

Abstract.
The representations that can be defined starting from positive sesquilinear maps $\Phi$ with values in an ordered Banach module are essentially of two types: if $\Phi$ is positive, a natural generalization follows the path of the GNS construction; if $\Phi$ is completely positive, one can move toward a possible generalization of the Stinespring dilation theorem. In this paper both possibilities are studied.
Key words and phrases:
Representations, modules, positive sesquilinear maps, completely positive sesquilinear maps, quasi normed spaces
1991 Mathematics Subject Classification:
Primary 46K10, 47A07, 16D10.
1. Introduction
Representations of *-algebras, partial *-algebras, modules over a *-algebra, are a common tool for the study of the properties of these structures, in particular when they are endowed with some topology and therefore an interplay between the algebraic aspects and the topological ones occurs. In a recent paper [3], we have considered a GNS-like construction defined by positive sesquilinear maps on a quasi *-algebra taking its values in a $C^{*}$-algebra $\mathfrak{C}$. The result was a *-representation with values in a space of operators acting on a certain quasi normed $\mathfrak{C}$-module. In this paper, we go some steps further and we consider the possibility of replacing sesquilinear maps taking values in a $C^{*}$-algebra with sesquilinear maps with values in a ordered Banach module $\mathfrak{Y}$ over some *-algebra $\mathfrak{Y}_{0}$. In this situation, an extension of the result of [3] is possible, provided that the sesquilinear form $\Phi$, where we start from, satisfies a generalized Cauchy-Schwarz inequality. This discussion is developed in Section 3. Section 4 is devoted to the *-representations induced by completely positive sesquilinear maps.
It turns out that, in our framework, several definitions of complete positivity can be introduced. First, a definition of complete positivity based directly on the structure of $\mathfrak{Y}$ (Definition 4.1) is given; then, following an idea proposed in [7, Ch. 11] and adopted in [2] for partial *-algebras, we consider completely positive maps from a normed quasi *-algebra $\mathfrak{A}$ into a set of sesquilinear maps on a vector space which take values in an ordered Banach module $\mathfrak{Y}$. We explore them with the aim of providing some generalization of the celebrated Stinespring dilation theorem [8] (see also [6, 9]), concerning the representation of a completely positive map from a $C^{*}$-algebra into the $C^{*}$-algebra of bounded operators in Hilbert space. This section also includes several examples and some results concerning sesquilinear forms produced by traces on a von Neumann algebra.
2. Preliminaries
A quasi *-algebra $(\mathfrak{A},\mathfrak{A}_{0})$ is a pair consisting of a vector space $\mathfrak{A}$ and a *-algebra $\mathfrak{A}_{0}$ contained in $\mathfrak{A}$ as a subspace and such that
  • $\mathfrak{A}$ carries an involution $a\mapsto a^{*}$ extending the involution of $\mathfrak{A}_{0}$;
  • $\mathfrak{A}$ is a bimodule over $\mathfrak{A}_{0}$ and the module multiplications extend the multiplication of $\mathfrak{A}_{0}$. In particular, the following associative laws hold:
    $(ca)d=c(ad);\ \ a(cd)=(ac)d,\quad\forall\ a\in\mathfrak{A},\ c,d\in\mathfrak{A}_{0};$
  • $(ac)^{*}=c^{*}a^{*}$, for every $a\in\mathfrak{A}$ and $c\in\mathfrak{A}_{0}$.
The identity or unit element of $(\mathfrak{A},\mathfrak{A}_{0})$, if any, is a necessarily unique element ${\sf e}\in\mathfrak{A}_{0}$, such that $a{\sf e}=a={\sf e}a$, for all $a\in\mathfrak{A}$.
We will always suppose that
$ac=0,\;\forall c\in\mathfrak{A}_{0}\Rightarrow a=0$
$ac=0,\;\forall a\in\mathfrak{A}\Rightarrow c=0.$
Clearly, both these conditions are automatically satisfied if $(\mathfrak{A},\mathfrak{A}_{0})$ has an identity ${\sf e}$.
[…]
  • there exists $\gamma>0$ such that for every $a\in\mathfrak{A}$
    $\max\{\|ac\|,\|ca\|\}\leq\gamma\|a\|,\quad\forall c\in\mathfrak{A}_{0};$
  • $\|a^{*}\|=\|a\|,\;\forall a\in\mathfrak{A};$
  • $\mathfrak{A}_{0}$ is dense in $\mathfrak{A}[\|\cdot\|]$.
If the normed vector space $\mathfrak{A}[\|\cdot\|]$ is complete, then $(\mathfrak{A},\mathfrak{A}_{0})$ is called a Banach quasi *-algebra.
Definition 2.1.
Let $\mathfrak{Y}$ be a Banach bimodule over the *-algebra $\mathfrak{Y}_{0}$ (with $\mathfrak{Y}_{0}$ equipped with a not necessarily sub-multiplicative norm). We say that $\mathfrak{Y}$ is an ordered Banach bimodule over $\mathfrak{Y}_{0}$ if
  • (i)
    $\mathfrak{Y}$ is ordered as a vector space; that is, $\mathfrak{Y}$ contains a (positive) closed cone $\mathfrak{K}$, i.e., $\mathfrak{K}\subset\mathfrak{Y}$ is such that $\mathfrak{K}+\mathfrak{K}\subset\mathfrak{K}$, $\lambda\mathfrak{K}\subset\mathfrak{K}$ for $\lambda\geq 0$ and $\mathfrak{K}\cap(-\mathfrak{K})=\{0\}$;
  • (ii)
    $z^{*}\mathfrak{K}z\subset\mathfrak{K},\quad\forall z\in\mathfrak{Y}_{0}.$
As usual, we will write $y_{1}\leq y_{2}$ whenever $y_{2}-y_{1}\in\mathfrak{K}$, with $y_{1},y_{2}\in\mathfrak{Y}$.
Example 2.2.
Examples of ordered Banach bimodule over a Banach *-algebra are provided by:
  • the non commutative spaces $\mathfrak{Y}=L^{2}(\rho)$, $\mathfrak{Y}_{0}=L^{\infty}(\rho)$, with $\rho$ a finite trace;
[…]
  • $\mathfrak{Y}=\mathfrak{Y}_{0}=\ell_{2}(C(\Omega))$, with $\ell_{2}(C(\Omega))$ the space of squared summable sequences of functions in $C(\Omega)$. It is a Banach module over itself because it is a (non-unital) normed *-algebra.
The definitions of the positive cones are obvious in all these cases.
Let ${\mathfrak{X}}$ be a vector space and $\mathfrak{Y}$ an ordered Banach module over $\mathfrak{Y}_{0}$ with positive cone $\mathfrak{K}$. Let $\Phi$ be a $\mathfrak{Y}$-valued positive sesquilinear map on ${\mathfrak{X}}\times{\mathfrak{X}}$
$\Phi:(x_{1},x_{2})\in{\mathfrak{X}}\times{\mathfrak{X}}\to\Phi(x_{1},x_{2})\in\mathfrak{Y};$
i.e., a map with the properties
  • $i)$
    $\Phi(x_{1},x_{1})\in\mathfrak{K}$,
  • $ii)$
    $\Phi(\alpha x_{1}+\beta x_{2},\gamma x_{3})=\overline{\gamma}[\alpha\Phi(x_{1},x_{3})+\beta\Phi(x_{2},x_{3})]$,
with $x_{1},x_{2},x_{3}\in{\mathfrak{X}}$ and $\alpha,\beta,\gamma\in\mathbb{C}$.
The $\mathfrak{Y}$-valued positive sesquilinear map $\Phi$ is called faithful if
$\Phi(x,x)=0_{\mathfrak{Y}}\;\Rightarrow\;x=0.$
3. Representations induced by positive sesquilinear maps
It is quite apparent that the cornerstone of every extension of the GNS-construction requires a Cauchy-Schwarz inequality or, at least, some generalization of it. We begin with presenting some results in this direction.
Throughout this section we will assume that $\mathfrak{Y}$ is (an ordered) Banach bimodule over a *-algebra $\mathfrak{Y}_{0}$ (with $\mathfrak{Y}_{0}$ equipped with a not necessarily sub-multiplicative norm $\|\cdot\|_{\mathfrak{Y}_{0}}$) with the respective cones $\mathfrak{K}$ and $\mathfrak{K}_{0}$, with $\mathfrak{K}_{0}=\{\sum_{i=1}^{N}z_{i}^{*}z_{i},z_{i}\in\mathfrak{Y}_{0},N\in\mathbb{N}\}$.
  • (D1)
    Every $z\in\mathfrak{Y}_{0}$ can be written as $z=z_{1}-z_{2}+i(z_{3}-z_{4})$, where $z_{j}\in\mathfrak{K}_{0}$ and $\|z_{j}\|_{\mathfrak{Y}_{0}}\leq\|z\|_{\mathfrak{Y}_{0}}$ for all $j\in\{1,\dots 4\}.$
  • (D2)
    There exists a family $\mathcal{F}$ of sesquilinear forms $\varphi:\mathfrak{Y}\times\mathfrak{Y}_{0}\to\mathbb{C}$ enjoying the following properties:
    • i)
      For each $y^{+}\in\mathfrak{K}$ and $z^{+}\in\mathfrak{K}_{0}$ we have that $\varphi(y^{+},z^{+})\geq 0$ for all $\varphi\in\mathcal{F}.$
    • ii)
      For each $y\in\mathfrak{Y}$ and $z\in\mathfrak{Y}_{0}$ we have that
      $|\varphi(y,z)|\leq\|y\|_{\mathfrak{Y}}\|z\|_{\mathfrak{Y}_{0}},\quad\forall\varphi\in\mathcal{F}.$
    • iii)
      For each $y\in\mathfrak{Y}$ we have that the norm in $\mathfrak{Y}$
      $\|y\|_{\mathfrak{Y}}=\sup_{\begin{subarray}{c}z\in\mathfrak{Y}_{0}\\ \|z\|_{\mathfrak{Y}_{0}}\leq 1\end{subarray}}\sup_{\varphi\in\mathcal{F}}|\varphi(y,z)|.$
Under these assumptions, the following proposition holds.
Proposition 3.1.
Let ${\mathfrak{X}}$ be a complex vector space, $\mathfrak{Y}$ an ordered Banach bimodule over the *-algebra $\mathfrak{Y}_{0}$ with positive closed cone $\mathfrak{K}$, satisfying $(D1)-(D2)$. Let $\Phi:{\mathfrak{X}}\times{\mathfrak{X}}\to\mathfrak{Y}$ be a positive sesquilinear map. Then, for all $x_{1},x_{2}\in{\mathfrak{X}}$ we have that
$\|\Phi(x_{1},x_{2})\|_{\mathfrak{Y}}\leq 4\|\Phi(x_{1},x_{1})\|_{\mathfrak{Y}}^{1/2}\|\Phi(x_{2},x_{2})\|_{\mathfrak{Y}}^{1/2}.$
Proof.
Let $\mathcal{F}$ be the family in $(D2)$. For each fixed $\varphi\in\mathcal{F}$ and $z^{+}\in\mathfrak{K}_{0}$, let $\widetilde{\varphi}_{z^{+}}$ be the sesquilinear form on ${\mathfrak{X}}\times{\mathfrak{X}}$ given by
$\widetilde{\varphi}_{z^{+}}(x_{1},x_{2})=\varphi(\Phi(x_{1},x_{2}),z^{+}),\quad\forall x_{1},x_{2}\in{\mathfrak{X}}.$
Since $\Phi$ is positive, by $(D2)$ $i)$, we obtain that $\widetilde{\varphi}_{z^{+}}$ is positive. Hence, by $(D2)$ $ii)$, for all $x_{1},x_{2}\in{\mathfrak{X}}$ we get that
(3.1) $|\widetilde{\varphi}_{z^{+}}(x_{1},x_{1})||\widetilde{\varphi}_{z^{+}}(x_{2},x_{2})|\leq\|\Phi(x_{1},x_{1})\|_{\mathfrak{Y}}\|\Phi(x_{2},x_{2})\|_{\mathfrak{Y}}\|z^{+}\|_{\mathfrak{Y}_{0}}^{2}.$
Now, if we choose a general $z\in\mathfrak{Y}_{0}$, by assumption, we can find $z_{1}^{+},z_{2}^{+},z_{3}^{+},z_{4}^{+}\in\mathfrak{K}_{0}$ such that $z=z_{1}^{+}-z_{2}^{+}+i(z_{3}^{+}-z_{4}^{+})$ and $\|z_{j}^{+}\|_{\mathfrak{Y}_{0}}\leq\|z\|_{\mathfrak{Y}_{0}}$ for all $j\in\{1,...,4\}.$ Hence, by the Cauchy-Schwarz inequality applied to the positive sesquilinear forms $\widetilde{\varphi}_{z_{j}^{+}}$ and by (3.1) it is
$|\varphi(\Phi(x_{1},x_{2}),z)|$ $\leq\sum_{j=1}^{4}|\varphi(\Phi(x_{1},x_{2}),z_{j}^{+})|$
$\leq\sum_{j=1}^{4}\|\Phi(x_{1},x_{1})\|_{\mathfrak{Y}}^{1/2}\|\Phi(x_{2},x_{2})\|_{\mathfrak{Y}}^{1/2}\|z_{j}^{+}\|_{\mathfrak{Y}_{0}}$
$\leq 4\|\Phi(x_{1},x_{1})\|_{\mathfrak{Y}}^{1/2}\|\Phi(x_{2},x_{2})\|_{\mathfrak{Y}}^{1/2}\|z\|_{\mathfrak{Y}_{0}}$
for all $x_{1},x_{2}\in{\mathfrak{X}}$ and all $\varphi\in\mathcal{F}$. Therefore, for every $z\in\mathfrak{Y}_{0}$ with $\|z\|_{\mathfrak{Y}_{0}}\leq 1$ and all $x_{1},x_{2}\in{\mathfrak{X}}$ we have
$\sup_{\varphi\in\mathcal{F}}|\varphi(\Phi(x_{1},x_{2}),z)|\leq 4\|\Phi(x_{1},x_{1})\|_{\mathfrak{Y}}^{1/2}\|\Phi(x_{2},x_{2})\|_{\mathfrak{Y}}^{1/2}.$
However, by $(D2)$ $iii)$, we deduce that for all $x_{1},x_{2}\in{\mathfrak{X}}$
$\|\Phi(x_{1},x_{2})\|_{\mathfrak{Y}}$ $=\sup_{\begin{subarray}{c}z\in\mathfrak{Y}_{0}\\ \|z\|_{\mathfrak{Y}_{0}}\leq 1\end{subarray}}\sup_{\varphi\in\mathcal{F}}|\varphi(\Phi(x_{1},x_{2}),z)|$
$\leq 4\|\Phi(x_{1},x_{1})\|_{\mathfrak{Y}}^{1/2}\|\Phi(x_{2},x_{2})\|_{\mathfrak{Y}}^{1/2}.$
We now give a series of examples of ordered Banach bimodules over *-algebras $\mathfrak{Y}_{0}$ and of families $\mathcal{F}$ of forms on $\mathfrak{Y}\times\mathfrak{Y}_{0}$ for which conditions $(D1)-(D2)$ hold.
Example 3.2.
Let $\rho$ be a finite trace on a Von Neumann algebra $\mathfrak{M}$. Put $\mathfrak{Y}=L^{2}(\rho)$, $\mathfrak{Y}_{0}=L^{\infty}(\rho)$ and $\mathfrak{K}$ be the natural cone consisting of positive elements and $\mathfrak{K}_{0}=L^{\infty}(\rho)\cap\mathfrak{K}$. In this case, we will consider $\mathfrak{Y}_{0}$ as a *-algebra equipped with the norm $\|\cdot\|_{2}$ (instead of $\|\cdot\|_{\infty}$). Take as =$\mathcal{F}=\{\varphi\}$, a family with a single sesquilinear form $\varphi:\mathfrak{Y}\times\mathfrak{Y}_{0}\to\mathbb{C}$ given by
$\varphi(Y,W)=\rho(W^{*}Y),\quad\forall Y\in\mathfrak{Y},W\in\mathfrak{Y}_{0}.$
To see that the condition $(D2)$ $i)$ holds, just observe that if $W\in L^{\infty}(\rho)\cap\mathfrak{K}$, then $W^{\frac{1}{2}}\in L^{\infty}(\rho)\cap\mathfrak{K}\subseteq L^{2}(\rho)\cap\mathfrak{K}$, so
$\rho(YW)=\rho(W^{\frac{1}{2}}YW^{\frac{1}{2}})\geq 0,\quad\forall Y\in L^{2}(\rho)\cap\mathfrak{K}.$
Condition $(D2)$ $ii)$ holds by the Cauchy-Schwarz inequality for $\rho$. As for condition $(D2)$ $iii)$, let $Y\in L^{2}(\rho)$, if $Y=0$, then obviously
$\|Y\|_{2}=\sup_{\begin{subarray}{c}W\in L^{\infty}(\rho)\\ \|W\|_{2}\leq 1\end{subarray}}|\rho(W^{*}Y)|=0.$
Now, suppose that $Y\neq 0$ and choose some $\{W_{n}\}_{n}\subseteq L^{\infty}(\rho)$ such that $W_{n}\to Y$ in $L^{2}(\rho)$ as $n\to\infty$. Assume without loss of generality that $W_{n}\neq 0$, for all $n\in\mathbb{N}$. Then, since $\|W_{n}\|_{2}\to\|Y\|_{2}$ as $n\to\infty$, it is not hard to see that
$\dfrac{W_{n}^{*}}{\|W_{n}\|_{2}}\to\dfrac{Y^{*}}{\|Y\|_{2}},\quad\mbox{ as }n\to\infty$
since $\|W_{n}^{*}-Y^{*}\|_{2}=\|W_{n}-Y\|_{2}$. Hence,
$\rho\left(\dfrac{W_{n}^{*}}{\|W_{n}\|_{2}}Y\right)\to\|Y\|_{2},\quad\mbox{ as }n\to\infty,$
moreover,
$\dfrac{\|W_{n}^{*}\|_{2}}{\|W_{n}\|_{2}}=1,\quad\forall n\in\mathbb{N}$
since $\|W_{n}^{*}\|_{2}=\|W_{n}\|_{2}$. Thus,
$\|Y\|_{2}\leq\sup_{\begin{subarray}{c}W\in L^{\infty}(\rho)\\ \|W\|_{2}\leq 1\end{subarray}}|\rho(W^{*}Y)|\leq\sup_{\begin{subarray}{c}W\in L^{\infty}(\rho)\\ \|W\|_{2}\leq 1\end{subarray}}\|W\|_{2}\|Y\|_{2}\leq\|Y\|_{2},$
so
$\|Y\|_{2}=\sup_{\begin{subarray}{c}W\in L^{\infty}(\rho)\\ \|W\|_{2}\leq 1\end{subarray}}|\rho(W^{*}Y)|.$
Similar considerations apply to $(L^{2}([0,1]),C([0,1]))$.
Example 3.3.
Let $\mathfrak{Y}=M(\Omega)$, $\mathfrak{Y}_{0}=C(\Omega)$, $\mathfrak{K}$ be the natural cone of positive Radon measures on $\Omega$, $\mathfrak{K}_{0}=\{f\in C(\Omega)|f\geq 0\}$ and $\mathcal{F}$ consist of the only sesquilinear form $\varphi:M(\Omega)\times C(\Omega)\to\mathbb{C}$ given by
[…]
Example 3.4.
Let $\mathfrak{Y}=\mathfrak{Y}_{0}=\ell_{2}(C(\Omega))$,
$\mathfrak{K}=\{\{f_{n}\}_{n}\in\ell_{2}(C(\Omega))|f_{n}\geq 0,\,\forall n\in\mathbb{N}\}$
and =$\mathcal{F}=\{\varphi_{t}|t\in\Omega\}$, with $\varphi_{t}:\ell_{2}(C(\Omega))\times\ell_{2}(C(\Omega))\to\mathbb{C}$ defined by
$\varphi_{t}(\{f_{n}\}_{n},\{g_{n}\}_{n})=\langle{\{g_{n}\}_{n}}|{\{f_{n}\}_{n}}\rangle(t).$
Corollary 3.5.
Let $\mathfrak{Y}$ be an ordered Banach bimodule over a *-algebra $\mathfrak{Y}_{0}$ (with $\mathfrak{Y}_{0}$ equipped with a not necessarily sub-multiplicative norm $\|\cdot\|_{\mathfrak{Y}_{0}}$) with the respective cones $\mathfrak{K}$ and $\mathfrak{K}_{0}$ satisfying $(D1)-(D2)$. Let $\mathfrak{A}_{0}$ be a unital *-algebra and $\omega:\mathfrak{A}_{0}\to\mathfrak{Y}$ be a positive linear map. Then
$\|\omega(c)\|_{\mathfrak{Y}}^{2}\leq 16\|\omega(c^{*}c)\|_{\mathfrak{Y}}\|\omega({\sf e})\|_{\mathfrak{Y}},\quad\forall c\in\mathfrak{A}_{0}.$
Proof.
To prove the statement it is enough to consider $\Phi:\mathfrak{A}_{0}\times\mathfrak{A}_{0}\to\mathfrak{Y}$, $\Phi(c,d)=\omega(d^{*}c)$ and apply Proposition 3.1. ∎
Definition 3.6.
Let ${\mathfrak{X}}$ be a vector space. A $\mathfrak{Y}$-valued faithful positive sesquilinear map $\Phi$ on ${\mathfrak{X}}\times{\mathfrak{X}}$ is said to be a $\mathfrak{Y}$-valued quasi inner product and we often will write $\langle{x_{1}}|{x_{2}}\rangle_{\Phi}:=\Phi(x_{1},x_{2})$, $x_{1},x_{2}\in{\mathfrak{X}}$.
A $\mathfrak{Y}$-valued quasi inner product $\Phi:{\mathfrak{X}}\times{\mathfrak{X}}\to\mathfrak{Y}$ induces a quasi norm $\|\cdot\|_{\Phi}$ on ${\mathfrak{X}}$:
$\|x\|_{\Phi}:=\sqrt{\|\langle{x}|{x}\rangle_{\Phi}\|_{\mathfrak{Y}}}=\sqrt{\|\Phi(x,x)\|_{\mathfrak{Y}}},\quad x\in{\mathfrak{X}},$
since (see [3])
$\|x_{1}+x_{2}\|_{\Phi}\leq\sqrt{2}(\|x_{1}\|_{\Phi}+\|x_{2}\|_{\Phi}),\qquad\forall x_{1},x_{2}\in{\mathfrak{X}}.$
The space ${\mathfrak{X}}$ is then a quasi normed space w.r.to the quasi norm $\|\cdot\|_{\Phi}$.
Definition 3.7.
If the complex vector space ${\mathfrak{X}}$ is complete w.r.to the quasi norm $\|\cdot\|_{\Phi}$, then ${\mathfrak{X}}$ is called a quasi Banach space with $\mathfrak{Y}$-valued quasi inner product or for short a quasi $B_{\mathfrak{Y}}$-space.
If $\Phi$ is not faithful, we can consider the set
$\mathfrak{N}_{\Phi}=\{x_{1}\in{\mathfrak{X}}:\,\Phi(x_{1},x_{2})=0_{\mathfrak{Y}},\forall x_{2}\in{\mathfrak{X}}\}.$
which is a subspace of ${\mathfrak{X}}$ (see [3]). If a generalized Cauchy-Schwarz inequality holds, then
$\mathfrak{N}_{\Phi}=\{x\in{\mathfrak{X}}:\,\Phi(x,x)=0_{\mathfrak{Y}}\}.$
We denote by $\Lambda_{\Phi}(x)$ the coset containing $x\in{\mathfrak{X}}$; i.e., $\Lambda_{\Phi}(x)=x+\mathfrak{N}_{\Phi}$ and define a $\mathfrak{Y}$-valued positive sesquilinear map on ${\mathfrak{X}}/\mathfrak{N}_{\Phi}\times{\mathfrak{X}}/\mathfrak{N}_{\Phi}$ as follows:
(3.2) $\langle{\Lambda_{\Phi}(x_{1})}|{\Lambda_{\Phi}(x_{2})}\rangle_{\Phi}:=\Phi(x_{1},x_{2}),\quad x_{1},x_{2}\in{\mathfrak{X}}.$
The associated quasi norm is:
(3.3) $\|\Lambda_{\Phi}(x)\|_{\Phi}:=\sqrt{\|\Phi(x,x)\|_{\mathfrak{Y}}},\quad x\in{\mathfrak{X}}.$
The quotient space ${\mathfrak{X}}/\mathfrak{N}_{\Phi}=\Lambda_{\Phi}({\mathfrak{X}})$ is a quasi normed space (see [3]). We denote by $\widetilde{{\mathfrak{X}}}$ the completion of $({\mathfrak{X}}/\mathfrak{N}_{\Phi},\|\cdot\|_{\Phi})$.
Let $\mathcal{K}$ be a quasi $B_{\mathfrak{Y}}$-space and $D(T)$ a dense subspace of $\mathcal{K}$. A linear map $T:D(T)\to\mathcal{K}$ is said $\Phi$-adjointable if there exists a linear map $T^{*}$ defined on a subspace $D(T^{*})\subset\mathcal{K}$ such that
$\Phi(T\xi,\eta)=\Phi(\xi,T^{*}\eta),\quad\forall\xi\in D(T),\eta\in D(T^{*}).$
Let $\mathcal{D}$ be a dense subspace of $\mathcal{K}$ and let us consider the following families of linear operators acting on $\mathcal{D}$:
${{\mathcal{L}}^{\mbox{\scriptsize${\dagger}$}}(\mathcal{D},\mathcal{K})}$ =$=\{T\mbox{ $\Phi$-adjointable},D(T)=\mathcal{D};\;D(T^{*})\supset\mathcal{D}\}$
${{\mathcal{L}}^{\dagger}(\mathcal{D})}$ =$=\{T\in{\mathcal{L}}^{\mbox{\scriptsize${\dagger}$}}(\mathcal{D},\mathcal{K}):T\mathcal{D}\subset\mathcal{D};\;T^{*}\mathcal{D}\subset\mathcal{D}\}$
${{\mathcal{L}}^{\dagger}(\mathcal{D})_{b}}$ =$=\{T\in{\mathcal{L}}^{\dagger}(\mathcal{D}):\,T\mbox{ is bounded on $\mathcal{D}$}\}.$
The involution in ${\mathcal{L}}^{\mbox{\scriptsize${\dagger}$}}(\mathcal{D},\mathcal{K})$ is defined by $T^{\dagger}:=T^{*}\upharpoonright\mathcal{D}$, the restriction of $T^{*}$, the $\Phi$-adjoint of $T$, to $\mathcal{D}$.
The sets ${\mathcal{L}}^{\dagger}(\mathcal{D})$ and ${{\mathcal{L}}^{\dagger}(\mathcal{D})_{b}}$ are *-algebras.
Remark 3.8.
If $T\in{\mathcal{L}}^{\mbox{\scriptsize${\dagger}$}}(\mathcal{D},\mathcal{K})$ then $T$ is closable. By definition $T$ is $\Phi$-adjointable. We denote by $T^{*}$ its $\Phi$-adjoint and by $D(T^{*})$ its domain. $T^{*}$ is a closed operator.
Remark 3.9.
[…]
$\Phi(T_{2}\xi,T_{1}\eta)=\Phi(W\xi,\eta),\quad\forall\xi,\eta\in\mathcal{D}.$
Due to the density of $\mathcal{D}$ in $\mathcal{K}$, the element $W$, if it exists, is unique. We put $W=T_{1}\,{\scriptstyle\square}\,T_{2}$.
Definition 3.10.
[…]
$(\mathcal{Y},\mathcal{D})_{w}^{\prime}=\{B\in{\mathfrak{B}}(\mathcal{K}):\Phi(BY\xi,\eta)=\Phi(B\xi,Y^{\dagger}\eta),\,\,\forall Y\in\mathcal{Y},\,\xi,\eta\in\mathcal{D}\}.$
Definition 3.11.
Let $(\mathfrak{A},\mathfrak{A}_{0})$ be a quasi *-algebra with unit ${\sf e}$. Let $\mathcal{D}$ be a dense subspace of a certain quasi $B_{\mathfrak{Y}}$-space $\mathcal{K}$ with $\mathfrak{Y}$-valued quasi inner product $\langle{\cdot}|{\cdot}\rangle_{\mathcal{K}}$. A linear map $\pi$ from $\mathfrak{A}$ into ${\mathcal{L}}^{\mbox{\scriptsize${\dagger}$}}(\mathcal{D},\mathcal{K})$ is called a *–representation of $(\mathfrak{A},\mathfrak{A}_{0})$, if the following properties are fulfilled:
  • (i)
    $\pi(a^{*})=\pi(a)^{\dagger}:=\pi(a)^{*}\upharpoonright\mathcal{D},\quad\forall\ a\in\mathfrak{A}$;
  • (ii)
    for $a\in\mathfrak{A}$ and $c\in\mathfrak{A}_{0}$, $\pi(a)\,{\scriptstyle\square}\,\pi(c)$ is well–defined and $\pi(a)\,{\scriptstyle\square}\,\pi(c)=\pi(ac)$.
We assume that for every *-representation $\pi$ of $(\mathfrak{A},\mathfrak{A}_{0})$, $\pi({\sf e})={{\mathbb{I}}_{\mathcal{D}}}$, the identity operator on the space $\mathcal{D}$.
The *-representation $\pi$ is said to be
  • closable if there exists $\widetilde{\pi}$ closure of $\pi$ defined as $\widetilde{\pi}(a)=\overline{\pi(a)}\upharpoonright{\widetilde{\mathcal{D}}}$ where $\widetilde{\mathcal{D}}$ is the completion under the graph topology $t_{\pi}$ defined by the seminorms $\xi\in\mathcal{D}\to\|\xi\|_{\mathcal{K}}+\|\pi(a)\xi\|_{\mathcal{K}}$, $a\in\mathfrak{A}$, where $\|\cdot\|_{\mathcal{K}}$ is the norm induced by the inner product of $\mathcal{K}$;
  • closed if $\mathcal{D}[t_{\pi}]$ is complete;
  • cyclic if there exists $\xi\in\mathcal{D}$ such that $\pi(\mathfrak{A}_{0})\xi$ is dense in $\mathcal{K}$ in its quasi norm topology.
Definition 3.12.
We denote by ${\mathcal{I}}_{{\mathfrak{A}}_{0}}^{\,\mathfrak{Y}}(\mathfrak{A})$ the set of all $\mathfrak{Y}$-valued positive sesquilinear maps on $\mathfrak{A}\times\mathfrak{A}$ with the following properties:
  • (i)
    $\Lambda_{\Phi}(\mathfrak{A}_{0})=\mathfrak{A}_{0}/\mathfrak{N}_{\Phi}$ is dense in the completion $\widetilde{\mathfrak{A}}$ of $\mathfrak{A}$.
  • (ii)
    $\Phi(ac,d)=\Phi(c,a^{*}d),\quad\forall\ a\in\mathfrak{A},\ c,d\in\mathfrak{A}_{0}$ (left-invariant).
Theorem 3.13.
Let $(\mathfrak{A},\mathfrak{A}_{0})$ be a quasi *-algebra with unit ${\sf e}$ and $\Phi\in{\mathcal{Q}}_{{\mathfrak{A}}_{0}}^{\,\mathfrak{Y}}(\mathfrak{A})$. The following statements are equivalent:
  • (i)
    $\Phi\in{\mathcal{I}}_{{\mathfrak{A}}_{0}}^{\,\mathfrak{Y}}(\mathfrak{A})$;
  • (ii)
    there exist a quasi $B_{\mathfrak{Y}}$-space $\mathcal{K}_{\Phi}$ with $\mathfrak{Y}$-valued inner product $\langle{\cdot}|{\cdot}\rangle_{\mathcal{K}_{\Phi}}$, a dense subspace $\mathcal{D}_{\Phi}\subseteq\mathcal{K}_{\Phi}$ and a closed cyclic *-representation $\pi:\mathfrak{A}\to{\mathcal{L}}^{\dagger}(\mathcal{D}_{\Phi},\mathcal{K}_{\Phi})$ with cyclic vector $\xi_{\Phi}$ such that
    $\langle{\pi(a)\xi}|{\eta}\rangle_{\mathcal{K}_{\Phi}}=\langle{\xi}|{\pi(a^{*})\eta}\rangle_{\mathcal{K}_{\Phi}},\quad\forall\xi,\eta\in\mathcal{D}_{\Phi},a\in\mathfrak{A}$
    and such that
    $\Phi(a,b)=\langle{\pi(a)\xi_{\Phi}}|{\pi(b)\xi_{\Phi}}\rangle_{\mathcal{K}_{\Phi}},\quad\forall a,b\in\mathfrak{A}.$
Proof.
The proof proceeds along the lines of that one of [3, Theorem 3.2], again due to Cauchy-Schwarz-like inequality for positive $\mathfrak{Y}$ sesquilinear maps.∎
4. Representations induced by completely positive sesquilinear maps
In this section we will suppose that $\mathfrak{Y}$ is an Banach bimodule over the normed *-algebra $\mathfrak{Y}_{0}$ such that
(4.1) $\|z_{2}^{*}yz_{1}\|_{\mathfrak{Y}}\leq\|z_{2}\|_{\mathfrak{Y}_{0}}\|z_{1}\|_{\mathfrak{Y}_{0}}\|y\|_{\mathfrak{Y}},\quad\forall y\in\mathfrak{Y},z_{1},z_{2}\in\mathfrak{Y}_{0}.$
Examples of such $\mathfrak{Y}$ and $\mathfrak{Y}_{0}$ can be find in Example 2.2 (where $L^{\infty}(\rho)$ has to be considered with the operator norm).
Definition 4.1.
Let $\mathfrak{S}$ be a vector space over $\mathbb{C}$, $\mathfrak{Y}$ an ordered Banach module over the normed *-algebra $\mathfrak{Y}_{0}$ with positive cone $\mathfrak{K}$. The positive sesquilinear map $\Phi:\mathfrak{S}\times\mathfrak{S}\to\mathfrak{Y}$ is called completely positive if, for every $N\in{\mathbb{N}}$, $z_{i}\in\mathfrak{Y}_{0},s_{i}\in\mathfrak{S}$, $i\in\{1,\ldots N\}$,
$\sum_{i,j=1}^{N}z_{j}^{*}\Phi(s_{i},s_{j})z_{i}\in\mathfrak{K}.$
In the following we will take $\mathfrak{S}=\mathfrak{A}$, a normed quasi *-algebra on a $C^{*}$-algebra $\mathfrak{A}_{0}$.
Let now $\mathfrak{X}$ be a normed vector space, $\mathfrak{Y}$ an ordered Banach module over the normed *-algebra $\mathfrak{Y}_{0}$ with positive cone $\mathfrak{K}$ and $\mathcal{S}_{\mathfrak{Y}}(\mathfrak{X})$ be the space of all $\mathfrak{Y}$-valued sesquilinear maps on $\mathfrak{X}$ and let $(\mathfrak{A},\mathfrak{A}_{0})$ be a unital normed quasi *-algebra.
Definition 4.2.
The sesquilinear map $\Phi:\mathfrak{A}\times\mathfrak{A}\to\mathcal{S}_{\mathfrak{Y}}(\mathfrak{X})$ will be called
  • bounded if, there exists a constant $M>0$ such that
    $\|\Phi(a,b)(x_{1},x_{2})\|_{\mathfrak{Y}}\leq M\|a\|\|b\|\|x_{1}\|\|x_{2}\|,\quad\forall a,b\in\mathfrak{A},x_{1},x_{2}\in\mathfrak{X};$
  • left-invariant if $\Phi(ac,d)=\Phi(c,a^{*}d),\quad\forall a\in\mathfrak{A},c,d\in\mathfrak{A}_{0}$;
  • right-invariant if $\Phi(ca,d)=\Phi(c,da^{*}),\quad\forall a\in\mathfrak{A},c,d\in\mathfrak{A}_{0}$;
  • completely positive if for every $N\in{\mathbb{N}}$, $a_{1},\dots,a_{N}\in\mathfrak{A}$,
    $x_{1},\dots,x_{N}\in{\mathfrak{X}}$,
    $\sum_{i,j=1}^{N}\Phi(a_{i},a_{j})(x_{i},x_{j})\in\mathfrak{K}.$
The main result on completely positive maps is the Stinespring Theorem, which is, in a sense, an extension to completely positive maps of the well known GNS construction for states (see e.g. [9, Theorem 1.2.7]). In [2, Theorem 3.9], a further generalization of the Stinespring theorem for completely positive, conjugate bilinear, quasi-invariant maps on a partial *-algebra $\mathfrak{A}$, with values in a vector space $\mathfrak{X}$ has been proved. In the following we consider certain positive sesquilinear map taking values on the space of all $\mathfrak{Y}$-valued sesquilinear maps on a vector space. The proof goes along the lines of that of [2]. We give here the proof for convenience of the reader.
Consider the algebraic tensor product $\mathfrak{A}\otimes{\mathfrak{X}}$ and its subspace
$\mathcal{N}_{\Phi}=\left\{\sum_{i=1}^{n}a_{i}\otimes x_{i}\in\mathfrak{A}\otimes\mathfrak{X}|\,\sum_{i=1}^{n}\sum_{j=1}^{n}\Phi(a_{i},a_{j})(x_{i},x_{j})=0_{\mathfrak{Y}}\right\}.$
Proposition 4.3.
Let $\mathfrak{X}$ be a normed complex vector space, $\mathfrak{Y}$ an ordered Banach module over the normed *-algebra $\mathfrak{Y}_{0}$ with positive cone $\mathfrak{K}$, $(\mathfrak{A},\mathfrak{A}_{0})$ be a normed quasi *-algebra with unit ${\sf e}$ and $\Phi:\mathfrak{A}\times\mathfrak{A}\to\mathcal{S}_{\mathfrak{Y}}(\mathfrak{X})$ be a left-invariant positive sesquilinear map. Let $\Phi$ be bounded with bound $M>0$. Then, $\Phi$ is completely positive if and only if there exist a quasi Banach space $\mathcal{K}_{\Phi}$, whose quasi norm is induced by an $\mathfrak{Y}$-valued (quasi)-inner product $\langle{\cdot}|{\cdot}\rangle_{\Phi}$, a dense subspace $\mathcal{D}_{\Phi}$ of $\mathcal{K}_{\Phi}$, a closed *-representation $\pi$ of $\mathfrak{A}$ in ${\mathcal{L}}^{\mbox{\scriptsize${\dagger}$}}(\mathcal{D}_{\Phi},\mathcal{K}_{\Phi})$ and a bounded linear operator $V:\mathfrak{X}\to\mathcal{D}_{\Phi}$ such that $\pi(\mathfrak{A})V{\mathfrak{X}}=(\mathfrak{A}\otimes{\mathfrak{X}})/\mathcal{N}_{\Phi}$ and, for all $a,b\in\mathfrak{A}$ and $x_{1},x_{2}\in\mathfrak{X}$, it holds that
$\Phi(a,b)(x_{1},x_{2})=\langle{\pi(a)Vx_{1}}|{\pi(b)Vx_{2}}\rangle_{\mathcal{K}_{\Phi}}.$
In this case $\|V\|^{2}\leq M\|{\sf e}\|^{2}$.
Proof.
Suppose that $\Phi$ is completely positive. Consider the algebraic tensor product $\mathfrak{A}\otimes{\mathfrak{X}}$, its subspace $\mathcal{N}_{\Phi}=\{u\in\mathfrak{A}\otimes\mathfrak{X}|\,\langle{u}|{u}\rangle=0\}$ and the quotient $(\mathfrak{A}\otimes\mathfrak{X})/\mathcal{N}_{\Phi}$. To keep the notation lighter, when there is no fear of confusion, we will identify the coset $(\sum_{i=1}^{n}a_{i}\otimes x_{i})+\mathcal{N}_{\Phi}$ for each $n\in\mathbb{N}$ and $a_{1},\dots,a_{n}\in\mathfrak{A}$, $x_{1},\dots,x_{n}\in\mathfrak{X}$ with its representative $\sum_{i=1}^{n}a_{i}\otimes x_{i}$. Define the sesquilinear map on $(\mathfrak{A}\otimes\mathfrak{X})/\mathcal{N}_{\Phi}$ by
$\langle{\sum_{i=1}^{n}a_{i}\otimes x_{i}}|{\sum_{j=1}^{m}b_{j}\otimes x^{\prime}_{j}}\rangle=\sum_{i=1}^{n}\sum_{j=1}^{m}\Phi(a_{i},b_{j})(x_{i},x^{\prime}_{j}),$
with $i\in\{1,\dots,N\}$; it is positive since for each $N\in\mathbb{N}$ and $a_{1},\dots,a_{N}\in\mathfrak{A}$, $x_{1},\dots,x_{N}\in\mathfrak{X}$ it is $\sum_{i,j=1}^{N}\Phi(a_{i},a_{j})(x_{i},x_{j})\in\mathfrak{K}$.
Hence, it defines a quasi inner product on the space $(\mathfrak{A}\otimes\mathfrak{X})/\mathcal{N}_{\Phi}$. Let $\mathcal{K}_{\Phi}$ be the quasi Banach space completion of $(\mathfrak{A}\otimes\mathfrak{X})/\mathcal{N}_{\Phi}$ under the induced quasi norm. We claim that $(\mathfrak{A}_{0}\otimes{\mathfrak{X}})/\mathcal{N}_{\Phi}$ is dense in $(\mathfrak{A}\otimes\mathfrak{X})/\mathcal{N}_{\Phi}$, hence in $\mathcal{K}_{\Phi}$. Indeed, if $a_{1},\dots,a_{N}\in\mathfrak{A}$, then there exist $N$ sequences $\{a_{n}^{(1)}\}_{n},\dots,\{a_{n}^{(N)}\}_{n}\subseteq\mathfrak{A}_{0}$ such that $a_{n}^{(i)}\to a_{i}$ as $n\to\infty$ for every $i\in\{1,\dots,N\}$. Then, by the assumptions, for any $x_{1},\dots,x_{N}\in\mathfrak{X}$, we get
$\|\langle{\sum_{i=1}^{N}(a_{n}^{(i)}-a_{i})\otimes x_{i}}|{\sum_{j=1}^{N}(a_{n}^{(j)}-a_{j})\otimes x_{j}}\rangle\|_{\mathfrak{Y}}\\ =\|\sum_{i,j=1}^{N}\Phi(a_{n}^{(i)}-a_{i},a_{n}^{(j)}-a_{j})(x_{i},x_{j})\|_{\mathfrak{Y}}\\ \leq M\sum_{i,j=1}^{N}\|a_{n}^{(i)}-a_{i}\|\|a_{n}^{(j)}-a_{j}\|\|x_{i}\|\|x_{j}\|\to 0,\quad\mbox{ as }n\to\infty.$
For each $a\in\mathfrak{A}$, let $\pi_{0}(a)$ be the linear map of $(\mathfrak{A}_{0}\otimes{\mathfrak{X}})/\mathcal{N}_{\Phi}$ into $(\mathfrak{A}\otimes\mathfrak{X})/\mathcal{N}_{\Phi}$ given by
$\pi_{0}(a)\left(\sum_{i=1}^{n}a_{i}\otimes x_{i}\right):=\sum_{i=1}^{n}aa_{i}\otimes x_{i},\quad a_{i}\in\mathfrak{A},x_{i}\in{\mathfrak{X}},n\in\mathbb{N}.$
Then it is well-defined:
$\langle{\pi_{0}(a)\left(\sum_{i=1}^{n}a_{i}\otimes x_{i}\right)}|{\sum_{j=1}^{m}b_{j}\otimes x^{\prime}_{j}}\rangle$
$=\sum_{i=1}^{n}\sum_{j=1}^{m}\Phi(aa_{i},b_{j})(x_{i},x^{\prime}_{j})=\sum_{i=1}^{n}\sum_{j=1}^{m}\Phi(a_{i},a^{*}b_{j})(x_{i},x^{\prime}_{j})$
$=\langle{\sum_{i=1}^{n}a_{i}\otimes x_{i}}|{\pi_{0}(a^{*})\left(\sum_{j=1}^{m}b_{j}\otimes x^{\prime}_{j}\right)}\rangle,$
for all $\sum_{i=1}^{n}a_{i}\otimes x_{i},\,\sum_{j=1}^{m}b_{j}\otimes x^{\prime}_{j}\in(\mathfrak{A}_{0}\otimes{\mathfrak{X}})/\mathcal{N}_{\Phi}$. Hence, if $\sum_{i}a_{i}\otimes x_{i}\in\mathcal{N}_{\Phi}$, then
$\langle{\pi_{0}(a)\left(\sum_{i=1}^{n}a_{i}\otimes x_{i}\right)}|{\sum_{j=1}^{m}b_{j}\otimes x^{\prime}_{j}}\rangle\\ \leq 4\left\|\sum_{i=1}^{n}a_{i}\otimes x_{i}\right\|\left\|\pi_{0}(a^{*})\left(\sum_{j=1}^{m}b_{j}\otimes x^{\prime}_{j}\right)\right\|=0,$
for every $\sum_{j=1}^{m}b_{j}\otimes x^{\prime}_{j}\in(\mathfrak{A}_{0}\otimes{\mathfrak{X}})/\mathcal{N}_{\Phi}$; by the density of $(\mathfrak{A}_{0}\otimes{\mathfrak{X}})/\mathcal{N}_{\Phi}$ in $\mathcal{K}_{\Phi}$ and by the joint continuity of the quasi inner product, it follows that $\left\|\pi_{0}(a)\left(\sum_{i=1}^{n}a_{i}\otimes x_{i}\right)\right\|=0$, thus $\pi_{0}(a)\left(\sum_{i=1}^{n}a_{i}\otimes x_{i}\right)\in\mathcal{N}_{\Phi}$. Therefore, $\pi_{0}(a)$ maps $\mathcal{N}_{\Phi}$ into itself, so $\pi_{0}(a)$ is a well-defined linear map on $(\mathfrak{A}_{0}\otimes{\mathfrak{X}})/\mathcal{N}_{\Phi}$. The map $\pi_{0}$ is a *-representation of $\mathfrak{A}$ in ${\mathcal{L}}^{\mbox{\scriptsize${\dagger}$}}((\mathfrak{A}_{0}\otimes{\mathfrak{X}})/\mathcal{N}_{\Phi},\mathcal{K}_{\Phi})$. By the density of $(\mathfrak{A}_{0}\otimes{\mathfrak{X}})/\mathcal{N}_{\Phi}$ and Remark 3.8, $\pi_{0}(a)$ is closable. Let $\pi$ be the closure of $\pi_{0}$ and let $\mathcal{D}_{\Phi}$ be its domain. Let now $V:\mathfrak{X}\to(\mathfrak{A}_{0}\otimes\mathfrak{X})/\mathcal{N}_{\Phi}\subset\mathcal{D}_{\Phi}$ be the operator given by
$Vx:={\sf e}\otimes x+\mathcal{N}_{\Phi},\quad x\in\mathfrak{X}.$
It is such that $\pi(\mathfrak{A})V{\mathfrak{X}}=(\mathfrak{A}\otimes{\mathfrak{X}})/\mathcal{N}_{\Phi}$, indeed, from $\pi(a)Vx=a\otimes x+\mathcal{N}_{\Phi}$, $a\in\mathfrak{A}$, $x\in{\mathfrak{X}}$. Then,
$\|Vx\|^{2}=\|\Phi({\sf e},{\sf e})(x,x)\|\leq M\|{\sf e}\|^{2}\|x\|^{2},\quad\forall x\in\mathfrak{X}$
thus $V$ is bounded with $\|V\|^{2}\leq M\|{\sf e}\|^{2}$. Finally, for all $a,b\in\mathfrak{A}$ and $x_{1},x_{2}\in\mathfrak{X}$, it holds that
$\Phi(a,b)(x_{1},x_{2})=\langle{\pi(a)Vx_{1}}|{\pi(b)Vx_{2}}\rangle.$
Conversely, suppose that for all $a,b\in\mathfrak{A}$ and $x_{1},x_{2}\in{\mathfrak{X}}$, we have that $\Phi(a,b)(x_{1},x_{2})=\langle{\pi(a)Vx_{1}}|{\pi(b)Vx_{2}}\rangle$. Then, for every $N\in\mathbb{N}$ and $a_{1},\dots,a_{N}\in\mathfrak{A}$, $x_{1},\dots,x_{N}\in{\mathfrak{X}}$, we get that
$\begin{split}\sum_{i,j=1}^{N}\Phi(a_{i},a_{j})(x_{i},x_{j})&=\sum_{i,j=1}^{N}\langle{\pi(a_{i})Vx_{i}}|{\pi(a_{j})Vx_{j}}\rangle\\ &=\langle{\sum_{i=1}^{N}\pi(a_{i})Vx_{i}}|{\sum_{j=1}^{N}\pi(a_{j})Vx_{j}}\rangle\in\mathfrak{K}.\end{split}$
In the following we will refer to a triple $(\pi,V,\mathcal{D}_{\Phi})$ as in Proposition 4.3 as a Stinespring triple decomposing $\Phi$.
Remark 4.4.
A similar result holds true for $\Phi:\mathfrak{A}\times\mathfrak{A}\to\mathcal{S}_{\mathfrak{Y}}(\mathfrak{X})$ bounded, right-invariant, completely positive sesquilinear map and, in this case, $\pi(a)\left(\sum_{i}a_{i}\otimes x_{i}\right)=\sum_{i}a_{i}a\otimes x_{i}$, $a_{i}\in\mathfrak{A}$, $x_{i}\in{\mathfrak{X}}$.
In analogy to Corollary 3.10 in [2], we can prove the following
Proposition 4.5.
Let $\mathfrak{X}$ be a normed complex vector space, $\mathfrak{Y}$ an ordered Banach module over the normed *-algebra $\mathfrak{Y}_{0}$ with positive cone $\mathfrak{K}$, $(\mathfrak{A},\mathfrak{A}_{0})$ be a normed quasi *-algebra with unit ${\sf e}$ and $\Phi:\mathfrak{A}\times\mathfrak{A}\to\mathcal{S}_{\mathfrak{Y}}(\mathfrak{X})$ be a bounded left-invariant positive sesquilinear map. If $\Phi$ is completely positive then the couple $(\pi,V)$ of the triple $(\pi,V,\mathcal{D}_{\Phi})$ decomposing $\Phi$ is uniquely determined up to unitary equivalence.
Remark 4.6.
Let $\mathfrak{Y}$ be an Banach bimodule over a normed *-algebra $\mathfrak{Y}_{0}$, satisfying (4.1). If $\Phi:\mathfrak{A}\times\mathfrak{A}\to\mathfrak{Y}$ is completely positive in the sense of Definition 4.1, then $\Phi$ induces a completely positive sesquilinear map $\widetilde{\Phi}:\mathfrak{A}\times\mathfrak{A}\to\mathcal{S}_{\mathfrak{Y}}(\mathfrak{Y}_{0})$, given by:
$\widetilde{\Phi}(a,b)(z_{1},z_{2})=z_{2}^{*}\Phi(a,b)z_{1},\quad\forall a,b\in\mathfrak{A},\,\,z_{1},z_{2}\in\mathfrak{Y}_{0}.$
Hence, by Proposition 4.3, there exists a Stinespring triple $(\pi,V,\mathcal{D}_{\Phi})$ decomposing $\widetilde{\Phi}$ with $V:\mathfrak{Y}_{0}\to\mathcal{D}_{\Phi}$ and $\mathcal{K}_{\Phi}=\widetilde{\mathcal{D}_{\Phi}}$ such that, for every $a,b\in\mathfrak{A}$, $z_{1},z_{2}\in\mathfrak{Y}_{0},$
$\widetilde{\Phi}(a,b)(z_{1},z_{2})=\langle{\pi(a)Vz_{1}}|{\pi(b)Vz_{2}}\rangle_{\mathcal{K}_{\Phi}}=z_{2}^{*}\Phi(a,b)z_{1}.$
In this case, we have that $\|Vz\|_{\mathcal{K}_{\Phi}}^{2}=\|z^{*}\Phi({\sf e},{\sf e})z\|_{\mathfrak{Y}}\leq\|z\|_{\mathfrak{Y}_{0}}^{2}\|\Phi({\sf e},{\sf e})\|_{\mathfrak{Y}}$, for all $z\in\mathfrak{Y}_{0}$.
In order to give a variant of the Radon-Nikodym Theorem for our framework in the case in which $\mathfrak{Y}=\mathbb{C}$ (Proposition 4.9 below), we need some preliminary lemmas.
Lemma 4.7.
Let $\mathfrak{X}$ be a normed complex vector space, $(\mathfrak{A},\mathfrak{A}_{0})$ be a normed quasi *-algebra and $\Phi:\mathfrak{A}\times\mathfrak{A}\to\mathcal{S}_{\mathbb{C}}(\mathfrak{X})$ be a bounded left-invariant completely positive sesquilinear map. Then there exists a Stinespring triple $(\pi,V,\mathcal{D}_{\Phi})$ decomposing $\Phi$, such that $\overline{\pi(\mathfrak{A})V{\mathfrak{X}}}=\mathcal{K}_{\Phi}$, with $\mathcal{K}_{\Phi}$ the Hilbert space completion of $\mathcal{D}_{\Phi}$.
Proof.
By Proposition 4.3, there exists a Stinespring triple $(\pi_{1},V_{1},\mathcal{D}_{\Phi})$ decomposing $\Phi$. Put $\mathcal{H}_{\Phi}$ the Hilbert space completion of $\mathcal{D}_{\Phi}$ (since in this case $\mathfrak{Y}=\mathbb{C}$). Let $P$ be the projection of $\mathcal{H}_{\Phi}$ onto the closure $\overline{\pi_{1}(\mathfrak{A})\mathcal{D}_{\Phi}}$. Then $P$ belongs to $(\pi_{1}(\mathfrak{A}),\mathcal{D}_{\Phi})_{w}^{\prime}$, the weak bounded commutant of $\pi_{1}(\mathfrak{A})$, in fact, $P$ is a bounded operator on $\mathcal{H}_{\Phi}$ and
$\begin{split}\langle{P\pi_{1}(a)\xi}|{\eta}\rangle_{\mathcal{H}_{\Phi}}&=\langle{\pi_{1}(a)\xi}|{P\eta}\rangle_{\mathcal{H}_{\Phi}}=\langle{P\xi}|{\pi_{1}(a^{*})\eta}\rangle_{\mathcal{H}_{\Phi}}\\ &=\langle{\xi}|{P\pi_{1}(a^{*})\eta}\rangle_{\mathcal{H}_{\Phi}},\qquad\forall a\in\mathfrak{A},\xi,\eta\in\mathcal{D}_{\Phi}.\end{split}$
This implies that $P\eta\in D(\pi_{1}(a)^{*})$ and $\pi_{1}(a)^{*}P\eta=P\pi_{1}(a)^{*}\eta$, hence $P\,{\scriptstyle\square}\,\pi_{1}(a)$ is well defined. Analogously we get
$\langle{\xi}|{P\pi_{1}(a^{*})\eta}\rangle_{\mathcal{H}_{\Phi}}=\langle{P\pi_{1}(a)\xi}|{\eta}\rangle_{\mathcal{H}_{\Phi}}$
for all $a\in\mathfrak{A},\xi,\eta\in\mathcal{D}_{\Phi}$, hence $\pi_{1}(a)\,{\scriptstyle\square}\,P$ is well defined too. Moreover, $\pi_{1}(a)\,{\scriptstyle\square}\,P=P\,{\scriptstyle\square}\,\pi_{1}(a)$ for all $a\in\mathfrak{A}$.
Let $\mathcal{K}_{\Phi}=P\mathcal{H}_{\Phi}$, $\pi=P\,{\scriptstyle\square}\,\pi_{1}$ and $V=PV_{1}$. Observe that the projection $P$ can be applied to a vector $V_{1}x$, $x\in{\mathfrak{X}}$, since $V_{1}x\in\mathcal{D}_{\Phi}$. The map $\pi=P\,{\scriptstyle\square}\,\pi_{1}$ is a *-representation. The linearity is obvious. Moreover, $\pi(a^{*})=\pi(a)^{\dagger}$, in fact:
$\pi(a^{*})=P\,{\scriptstyle\square}\,\pi_{1}(a^{*})=\pi_{1}(a^{*})\,{\scriptstyle\square}\,P=(P\,{\scriptstyle\square}\,\pi_{1}(a))^{\dagger}=\pi(a)^{\dagger},\quad\forall a\in\mathfrak{A}.$
Recalling that $P\pi_{1}(a)\eta=\pi_{1}(a)\eta$, for all $a\in\mathfrak{A}$ and $\eta\in\mathcal{D}_{\Phi}$, the equality $\pi(ac)=\pi(a)\,{\scriptstyle\square}\,\pi(c)$ holds, for every $a\in\mathfrak{A}$ and $c\in\mathfrak{A}_{0}$. Indeed, if $\xi,\eta\in\mathcal{D}_{\Phi}$, then
$\langle{\pi(ac)\xi}|{\eta}\rangle_{\mathcal{K}_{\Phi}}$ $=\langle{P\pi_{1}(ac)\xi}|{\eta}\rangle_{\mathcal{K}_{\Phi}}=\langle{P(\pi_{1}(a)\,{\scriptstyle\square}\,\pi_{1}(c))\xi}|{\eta}\rangle_{\mathcal{K}_{\Phi}}$
$=\langle{P\pi_{1}(c)\xi}|{\pi_{1}(a^{*})\eta}\rangle_{\mathcal{K}_{\Phi}}=\langle{P\pi_{1}(c)\xi}|{P\pi_{1}(a^{*})\eta}\rangle_{\mathcal{K}_{\Phi}}$
$=\langle{P\pi_{1}(c)\xi}|{\pi_{1}(a^{*})\,{\scriptstyle\square}\,P\eta}\rangle_{\mathcal{K}_{\Phi}}.$
hence $(P\,{\scriptstyle\square}\,\pi_{1}(a))\,{\scriptstyle\square}\,(P\pi_{1}(x))$ is well-defined and
$\pi(ac)=\pi(a)\,{\scriptstyle\square}\,\pi(c),\quad\forall a\in\mathfrak{A},c\in\mathfrak{A}_{0}.$
Finally, the domain of $\pi$ is $\mathcal{D}_{\Phi}$, since the weak commutation holds. In addition,
$\begin{split}\overline{\pi(\mathfrak{A})V{\mathfrak{X}}}&=\overline{P\pi_{1}(\mathfrak{A})PV_{1}{\mathfrak{X}}}=\overline{P^{2}\pi_{1}(\mathfrak{A})V_{1}{\mathfrak{X}}}\\ &=\overline{P\pi_{1}(\mathfrak{A})V_{1}{\mathfrak{X}}}=P(\overline{\pi_{1}(\mathfrak{A})V_{1}{\mathfrak{X}}})=\mathcal{K}_{\Phi}\end{split}$
and for all $a,b\in\mathfrak{A}$ and $x_{1},x_{2}\in\mathfrak{X}$
$\langle{\pi(a)Vx_{1}}|{\pi(b)Vx_{2}}\rangle_{\mathcal{K}_{\Phi}}$ $=\langle{P\pi_{1}(a)PV_{1}x_{1}}|{P\pi_{1}(b)PV_{1}x_{2}}\rangle_{\mathcal{K}_{\Phi}}$
$=\langle{P(\pi_{1}(a)\,{\scriptstyle\square}\,P)V_{1}x_{1}}|{P(\pi_{1}(b)\,{\scriptstyle\square}\,P)V_{1}x_{2}}\rangle_{\mathcal{K}_{\Phi}}$
$=\langle{P(P\,{\scriptstyle\square}\,\pi_{1}(a))V_{1}x_{1}}|{P(P\,{\scriptstyle\square}\,\pi_{1}(b))V_{1}x_{2}}\rangle_{\mathcal{K}_{\Phi}}$
$=\langle{P\pi_{1}(a)V_{1}x_{1}}|{P\pi_{1}(b)V_{1}x_{2}}\rangle_{\mathcal{K}_{\Phi}}$
$=\langle{\pi_{1}(a)V_{1}x_{1}}|{\pi_{1}(b)V_{1}x_{2}}\rangle_{\mathcal{H}_{\Phi}}=\Phi(a,b)(x_{1},x_{2}).$
Lemma 4.8.
Let $\mathfrak{X}$ be a normed complex vector space, $(\mathfrak{A},\mathfrak{A}_{0})$ be a normed quasi *-algebra with unit ${\sf e}$ and $\Psi,\Phi:\mathfrak{A}\times\mathfrak{A}\to\mathcal{S}_{\mathbb{C}}(\mathfrak{X})$ be bounded left-invariant completely positive sesquilinear maps. Suppose that there exists $\gamma>0$ such that, for every $n\in\mathbb{N}$, for all $a_{i},b_{j}\in\mathfrak{A},x_{i},y_{j}\in{\mathfrak{X}},i,j\in\{1,...,n\}$
(4.2) $\sum\Psi(a_{i},b_{j})(x_{i},y_{j})\leq\gamma\sum\Phi(a_{i},b_{j})(x_{i},y_{j}).$
Let $(\pi_{1},V_{1},\mathcal{D}_{\Psi})$ and $(\pi_{2},V_{2},\mathcal{D}_{\Phi})$ be two Stinespring triples decomposing respectively $\Psi$ and $\Phi$ such that the closures $\overline{\pi_{1}(\mathfrak{A})V_{1}{\mathfrak{X}}}=\mathcal{K}_{\Psi}$ and $\overline{\pi_{2}(\mathfrak{A})V_{2}{\mathfrak{X}}}=\mathcal{K}_{\Phi}$. Then there exists an operator $T:\mathcal{K}_{\Phi}\to\mathcal{K}_{\Psi}$ with $\|T\|\leq\gamma$ such that:
  • (i)
    $TV_{2}=V_{1}$.
  • (ii)
    $T\,{\scriptstyle\square}\,\pi_{2}(a)=\pi_{1}(a)\,{\scriptstyle\square}\,T$, for every $a\in\mathfrak{A}$.
Proof.
Let $N\in\mathbb{N}$ and $a_{1},\dots,a_{N}\in\mathfrak{A}$ and $x_{1},\dots,x_{N}\in{\mathfrak{X}}$ then
$\begin{split}\left\|\sum_{i=1}^{N}\pi_{1}(a_{i})V_{1}x_{i}\right\|^{2}&=\langle{\sum_{i=1}^{N}\pi_{1}(a_{i})V_{1}x_{i}}|{\sum_{j=1}^{N}\pi_{1}(a_{j})V_{1}x_{j}}\rangle\\ &=\sum_{i,j=1}^{N}\Psi(a_{i},a_{j})(x_{i},x_{j})\\ &\leq\gamma\sum_{i,j=1}^{N}\Phi(a_{i},a_{j})(x_{i},x_{j})\\ &=\gamma\left\|\sum_{i=1}^{N}\pi_{2}(a_{i})V_{2}x_{i}\right\|^{2}.\end{split}$
If $\sum_{i=1}^{N}\pi_{2}(a_{i})V_{2}x_{i}=0$ then also $\sum_{i=1}^{N}\pi_{1}(a_{i})V_{1}x_{i}=0$, hence, we can define a unique bounded linear operator $T$ by
(4.3) $T\sum_{i=1}^{N}\pi_{2}(a_{i})V_{2}x_{i}=\sum_{i=1}^{N}\pi_{1}(a_{i})V_{1}x_{i},\quad\forall a_{i}\in\mathfrak{A},\,x_{i}\in{\mathfrak{X}},i\in\{1,\dots,N\}.$
and extend it to the closure $\mathcal{K}_{\Phi}=\overline{\pi_{2}(\mathfrak{A})V_{2}{\mathfrak{X}}}$ into $\mathcal{K}_{\Psi}=\overline{\pi_{1}(\mathfrak{A})V_{1}{\mathfrak{X}}}$. Moreover, by taking $a={\sf e}\in\mathfrak{A}_{0}$, since $\pi_{2}({\sf e})=I_{\mathcal{D}_{\Phi}}$ and $\pi_{1}({\sf e})=I_{\mathcal{D}_{\Psi}}$ then $TV_{2}=V_{1}$.
Then $T\,{\scriptstyle\square}\,\pi_{2}(a)=\pi_{1}(a)\,{\scriptstyle\square}\,T$, for every $a\in\mathfrak{A}$. If now $a\in\mathfrak{A}$, $c\in\mathfrak{A}_{0}$ $x\in{\mathfrak{X}}$, by (4.3)
$\begin{split}T\,{\scriptstyle\square}\,(\pi_{2}(a)\,{\scriptstyle\square}\,\pi_{2}(c))V_{2}x&=T\pi_{2}(ac)V_{2}x=\pi_{1}(ac)V_{1}x=\pi_{1}(a)\,{\scriptstyle\square}\,\pi_{1}(c)V_{1}x\\ &=\pi_{1}(a)\,{\scriptstyle\square}\,T\pi_{2}(c)V_{2}x=(\pi_{1}(a)\,{\scriptstyle\square}\,T)\,{\scriptstyle\square}\,\pi_{2}(c)V_{2}x.\end{split}$
the set $\pi_{2}(\mathfrak{A}_{0})V_{2}{\mathfrak{X}}=(\mathfrak{A}_{0}\otimes{\mathfrak{X}})/\mathcal{N}_{\Phi}$ is dense in $\mathcal{K}_{\Phi}$ (see the proof of Proposition 4.3), hence we conclude that $T\pi_{2}(a)=\pi_{1}(a)\,{\scriptstyle\square}\,T$ for every $a\in\mathfrak{A}$. ∎
Proposition 4.9.
Let $\mathfrak{X}$ be a normed vector space, $(\mathfrak{A},\mathfrak{A}_{0})$ be a normed quasi *-algebra with unit ${\sf e}$ and $\Phi,\Psi:\mathfrak{A}\times\mathfrak{A}\to\mathcal{S}_{\mathbb{C}}(\mathfrak{X})$ be bounded left-invariant completely positive sesquilinear maps and such that (4.2) holds for some $\gamma>0$. Let $(\pi,V,\mathcal{D}_{\Phi})$ be a Stinespring triple decomposing $\Phi$ such that $\overline{\pi(\mathfrak{A})V{\mathfrak{X}}}=\mathcal{K}_{\Phi}$. Then there exists a unique operator $W\in(\pi(\mathfrak{A}),\mathcal{D}_{\Phi})_{w}^{\prime}$, the weak commutant of $\pi(\mathfrak{A})$, with $0\leq W\leq\gamma^{2}I$ such that for all $a,b\in\mathfrak{A}$ and $x_{1},x_{2}\in\mathfrak{X}$, it holds that
$\Psi(a,b)(x_{1},x_{2})=\langle{W\pi(a)Vx_{1}}|{\pi(b)Vx_{2}}\rangle_{\mathcal{K}_{\Phi}}.$
Proof.
By Lemma 4.7, there exists a triple $(\pi,V,\mathcal{D}_{\Phi})$ decomposing $\Phi$ such that $\mathcal{K}_{\Phi}=\overline{\pi(\mathfrak{A})V{\mathfrak{X}}}$. Consider $W\in(\pi(\mathfrak{A}),\mathcal{D}_{\Phi})_{w}^{\prime}$, with $0\leq W\leq\gamma^{2}I$, and the sesquilinear map $\Phi_{W}:\mathfrak{A}\times\mathfrak{A}\to\mathcal{S}_{\mathbb{C}}(\mathfrak{X})$
$\Phi_{W}(a,b)(x_{1},x_{2}):=\langle{W\pi(a)Vx_{1}}|{\pi(b)Vx_{2}}\rangle_{\mathcal{K}_{\Phi}},\quad a,b\in\mathfrak{A},x_{1},x_{2}\in\mathfrak{X}.$
We want to prove that $\Psi=\Phi_{W}$ for some $W\in(\pi(\mathfrak{A}),\mathcal{D}_{\Phi})_{w}^{\prime}$ and $0\leq W\leq\gamma^{2}I$. By Lemma 4.7, there exists a triple $(\pi_{1},V_{1},\mathcal{D}_{\Psi})$ decomposing $\Psi$ such that $\mathcal{K}_{\Psi}=\overline{\pi_{1}(\mathfrak{A})V_{1}{\mathfrak{X}}}$. By Lemma 4.8 there exists an operator $T:\mathcal{K}_{\Phi}\to\mathcal{K}_{\Psi}$ with $\|T\|\leq\gamma$ such that $TV=V_{1}$ and $T\,{\scriptstyle\square}\,\pi(a)=\pi_{1}(a)\,{\scriptstyle\square}\,T$, $a\in\mathfrak{A}$. Hence it is also $\pi(a)\,{\scriptstyle\square}\,T^{*}=T^{*}\,{\scriptstyle\square}\,\pi_{1}(a)$, $a\in\mathfrak{A}$. Put $W=T^{*}T$, then $0\leq W\leq\gamma^{2}I$. Moreover $W\in(\pi(\mathfrak{A}),\mathcal{D}_{\Phi})_{w}^{\prime}$, in fact:
$\begin{split}\langle{W\pi(a)\xi}|{\eta}\rangle_{\mathcal{K}_{\Phi}}&=\langle{T^{*}T\pi(a)\xi}|{\eta}\rangle_{\mathcal{K}_{\Phi}}\\ &=\langle{T^{*}\pi_{1}(a)\,{\scriptstyle\square}\,T\xi}|{\eta}\rangle_{\mathcal{K}_{\Phi}}\\ &=\langle{\pi(a)\,{\scriptstyle\square}\,T^{*}T\xi}|{\eta}\rangle_{\mathcal{K}_{\Phi}}\\ &=\langle{W\xi}|{\pi(a^{*})\eta}\rangle_{\mathcal{K}_{\Phi}},\qquad\forall a\in\mathfrak{A},\xi,\eta\in\mathcal{D}_{\Phi}.\end{split}$
This implies also that $W^{\dagger}\eta=W\eta\in D(\pi(a)^{*})$ and $\langle{\xi}|{\pi(a)^{*}W\eta}\rangle_{\mathcal{K}_{\Phi}}=\langle{\xi}|{W\pi(a)^{*}\eta}\rangle_{\mathcal{K}_{\Phi}}$, hence $W\,{\scriptstyle\square}\,\pi(a)$ is well defined. Analogously we get $\langle{\xi}|{W\pi(a^{*})\eta}\rangle_{\mathcal{K}_{\Phi}}=\langle{W\pi(a)\xi}|{\eta}\rangle_{\mathcal{K}_{\Phi}}$ for all $a\in\mathfrak{A},\xi,\eta\in\mathcal{D}_{\Phi}$, hence $\pi(a)\,{\scriptstyle\square}\,W$ is well defined too. Moreover, $\pi(a)\,{\scriptstyle\square}\,W=W\,{\scriptstyle\square}\,\pi(a)$ for all $a\in\mathfrak{A}$.
Finally, for every $a,b\in\mathfrak{A}$, $x_{1},x_{2}\in{\mathfrak{X}}$
$\begin{split}\Phi_{W}(a,b)(x_{1},x_{2})&=\langle{W\pi(a)Vx_{1}}|{\pi(b)Vx_{2}}\rangle_{\mathcal{K}_{\Phi}}\\ &=\langle{T^{*}T\pi(a)Vx_{1}}|{\pi(b)Vx_{2}}\rangle_{\mathcal{K}_{\Phi}}\\ &=\langle{T\pi(a)Vx_{1}}|{T\pi(b)Vx_{2}}\rangle_{\mathcal{K}_{\Phi}}\\ &=\langle{(\pi_{1}(a)\,{\scriptstyle\square}\,T)Vx_{1}}|{(\pi_{1}(b)\,{\scriptstyle\square}\,T)Vx_{2}}\rangle_{\mathcal{K}_{\Phi}}\\ &=\langle{\pi_{1}(a)V_{1}x_{1}}|{\pi_{1}(b)V_{1}x_{2}}\rangle_{\mathcal{K}_{\Phi}}=\Psi(a,b)(x_{1},x_{2}).\end{split}$
Now we prove the uniqueness of $W$. Consider the linear map $W\to\Phi_{W}$, with $\Phi_{W}$ defined as before. If $\Phi_{W}=0$, then, for all $a,b\in\mathfrak{A}$ and $x_{1},x_{2}\in\mathfrak{X}$
$\Phi_{W}(a,b)(x_{1},x_{2})=\langle{W\pi(a)Vx_{1}}|{\pi(b)Vx_{2}}\rangle_{\mathcal{K}_{\Phi}}=0$
and recalling that $\mathcal{K}_{\Phi}=\overline{\pi(\mathfrak{A})V{\mathfrak{X}}}$ we get that $W=0$. ∎
Definition 4.10.
Let $(\mathfrak{A},\mathfrak{A}_{0})$ be a quasi *-algebra, $\mathfrak{Y}$ an ordered Banach module over the normed *-algebra $\mathfrak{Y}_{0}$ with positive cone $\mathfrak{K}$. Let $\omega:\mathfrak{A}\to\mathfrak{Y}$ be a bounded linear map. We shall say that $\omega$ is completely positive if for every $N\in{\mathbb{N}}$, $z_{i},\in\mathfrak{Y}_{0},c_{i}\in\mathfrak{\mathfrak{A}}_{0}$, $i\in{1,\ldots N}$
$\sum_{i,j=1}^{N}z_{j}^{*}\omega(c_{j}^{*}c_{i})z_{i}\in\mathfrak{K}$
(i.e., it is positive in $\mathfrak{Y}$).
Proposition 4.11.
Let $(\mathfrak{A},\mathfrak{A}_{0})$ be a normed quasi *-algebra, $\mathfrak{Y}$ an ordered Banach module over the normed *-algebra $\mathfrak{Y}_{0}$ with positive cone $\mathfrak{K}$. Let $\omega:\mathfrak{A}\to\mathfrak{Y}$ be a bounded completely positive linear map and let $M>0$ be such that
$\|\omega(c_{2}^{*}c_{1})\|_{\mathfrak{Y}}\leq M\|c_{1}\|_{\mathfrak{A}}\|c_{2}\|_{\mathfrak{A}},\quad c_{1},c_{2}\in\mathfrak{A}_{0}$
then there exists a $B_{\mathfrak{Y}}$-space $\mathcal{K}_{\omega}$ with a dense domain $\mathcal{D}_{\omega}\subseteq\mathcal{K}_{\omega}$, a closed *-representation $\pi$ of $\mathfrak{A}$ on $\mathcal{D}_{\omega}$ and a bounded linear operator $V:\mathfrak{Y}_{0}\to\mathcal{D}_{\omega}$ such that
$z_{2}^{*}\omega(a)z_{1}=\langle{\pi(a)Vz_{1}}|{Vz_{2}}\rangle_{\mathcal{K}_{\omega}},\quad\forall a\in\mathfrak{A},\,\,z_{1},z_{2}\in\mathfrak{Y}_{0}.$
In particular, if ${\sf e}_{\mathfrak{Y}_{0}}$ is the unit of $\mathfrak{Y}_{0}$, then
$\omega(a)=\langle{\pi(a)V{\sf e}_{\mathfrak{Y}_{0}}}|{V{\sf e}_{\mathfrak{Y}_{0}}}\rangle_{\mathcal{K}_{\omega}},\quad\forall a\in\mathfrak{A}.$
Proof.
Define $\Phi_{0}:\mathfrak{A}_{0}\times\mathfrak{A}_{0}\to\mathfrak{Y}$ by
$\Phi_{0}(c_{1},c_{2})=\omega(c_{2}^{*}c_{1}),\quad\forall c_{1},c_{2}\in\mathfrak{A}_{0}.$
Then, by the same argument as in [3, Corollary 3.13], $\Phi$ extends by continuity to a bounded, left-invariant sesquilinear map $\Phi:\mathfrak{A}\times\mathfrak{A}\to\mathfrak{Y}$. Now, given $N\in\mathbb{N}$ and $a_{1},\dots,a_{N}\in\mathfrak{A}$, choose for each $i\in\{1,\dots,N\}$ sequences $\{c_{n}^{1}\}_{n}$, …$\{c_{n}^{N}\}_{n}\subset\mathfrak{A}_{0}$ such that $c_{n}^{i}\to a_{i}$ as $n\to\infty$, for all $i\in\{1,\dots,N\}$. Since $\Phi_{0}(c_{n}^{i},c_{n}^{j})\to\Phi(a_{i},a_{j})$ as $n\to\infty$, for all $i,j\in\{1,\dots,N\}$ and
$\|z_{2}^{*}yz_{1}\|_{\mathfrak{Y}}\leq\|z_{1}\|_{\mathfrak{Y}_{0}}\|z_{2}\|_{\mathfrak{Y}_{0}}\|y\|_{\mathfrak{Y}},\quad\forall y\in\mathfrak{Y},\,\,z_{1},z_{2}\in\mathfrak{Y}_{0},$
it follows that for every $N\in\mathbb{N}$ and $a_{1},\dots,a_{N}\in\mathfrak{A}$, $z_{1},\dots,z_{N}\in\mathfrak{Y}_{0}$
$\sum_{i,j=1}^{N}z_{i}^{*}\Phi_{0}(c_{n}^{j},c_{n}^{i})z_{j}\to\sum_{i,j=1}^{N}z_{i}^{*}\Phi(a_{j},a_{i})z_{j}.$
Now, for each $n\in\mathbb{N}$, since $\Phi_{0}(c_{n}^{j},c_{n}^{i})=\omega((c_{n}^{i})^{*}c_{n}^{j})$ with $i,j\in\{1,\dots,N\}$ we have that $\sum_{i,j=1}^{N}z_{i}^{*}\Phi_{0}(c_{n}^{j},c_{n}^{i})z_{j}\in\mathfrak{K}$ for every $n\in\mathbb{N}$. Therefore, $\sum_{i,j=1}^{N}z_{j}^{*}\Phi(a_{i},a_{j})z_{i}$ is positive too. Then, we use that $\omega(a)=\Phi(a,{\sf e})$, for all $a\in\mathfrak{A}$ which can be proved by the same argument as in [3, Corollary 3.13] and apply Remark 4.6. ∎
Remark 4.12.
Every bounded left-invariant, completely positive, sesquilinear map $\Phi:\mathfrak{A}\times\mathfrak{A}\to\mathfrak{Y}$ induces a completely positive linear map $\omega$ satisfying the assumptions of Proposition 4.11, it suffices to take $\omega(a)=\Phi(a,{\sf e})$, for all $a\in\mathfrak{A}$. Then, if $c_{1},c_{2}\in\mathfrak{A}_{0}$, since $\Phi$ is left-invariant and bounded, we get
$\|\omega(c_{2}^{*}c_{1})\|$ $=\|\Phi(c_{2}^{*}c_{1},{\sf e})\|=\|\Phi(c_{1},c_{2})\|$
$\leq\|\Phi\|\|c_{2}\|\|c_{1}\|;$
similarly, we get, for every $N\in\mathbb{N}$
$\sum_{i,j=1}^{N}z_{j}^{*}\omega(c_{j}^{*}c_{i})z_{i}=\sum_{i,j=1}^{N}z_{j}^{*}\Phi(c_{i},c_{j})z_{i}\in\mathfrak{K},.$
for every $c_{1},\dots,c_{N}\in\mathfrak{A}_{0}$ and $z_{1},\dots,z_{N}\in\mathfrak{Y}_{0}$.
Example 4.13.
Let $\rho$ be a finite trace, $W\in L^{\infty}(\rho)$ and $\omega:L^{\infty}(\rho)\to L^{2}(\rho)$ be given by $\omega(X)=W^{*}XW$.
Example 4.14.
Let $\rho$ be a finite trace, $V\in L^{4}(\rho)$ and $\omega:L^{\infty}(\rho)\to L^{2}(\rho)$ be given by $\omega(A)=V^{*}AV$.
Example 4.15.
Let $\mu$ be the Lebesgue measure on $[0,1]$, $w$ be a positive continuous function on $[0,1]$, $\alpha$ be a homeomorphism of $[0,1]$ and $\omega:L^{2}([0,1])\to M([0,1])$ be given by $d\omega(f)=w(f\circ\alpha)d\mu$.
Example 4.16.
Let $(\mathfrak{A},\mathfrak{A}_{0})=(L^{2}([0,1]),C([0,1]))$, $\alpha$ be a homeomorphism of $[0,1]$ $w$ be a positive continuous function on $[0,1]$. If $w(f)=w(f\circ\alpha)$ for all $f\in L^{2}([0,1])$, then $w$ is completely positive.
Now we further generalize some previous results.
Definition 4.17.
Let $(\mathfrak{A},\mathfrak{A}_{0})$ be a quasi *-algebra, $\mathfrak{Y}$ an ordered Banach module over the normed *-algebra $\mathfrak{Y}_{0}$ with positive cone $\mathfrak{K}$, and $\varphi\in\mathfrak{K}^{*}$ (i.e., $\varphi(y^{+})\geq 0$, for all $y^{+}\in\mathfrak{K}$, where $\varphi\in\mathfrak{Y}^{*}$, the dual of $\mathfrak{Y}$). A bounded, left-invariant, positive sesquilinear map $\Phi:\mathfrak{A}\times\mathfrak{A}\to\mathfrak{Y}$ will be called completely positive with respect to $\varphi$ if, for every $N\in\mathbf{N}$ and each $a_{j}\in\mathfrak{A}$, $y_{i}\in\mathfrak{Y}$, $i,j\in\{1,...,N\}$, it is
$\varphi\left(\sum_{i,j=1}^{N}y_{j}^{*}\Phi(a_{i},a_{j})y_{i}\right)\geq 0.$
Remark 4.18.
If $\Phi$ is completely positive with respect to $\varphi$, it induces $\widetilde{\Phi}:\mathfrak{A}\times\mathfrak{A}\to\mathcal{S}_{\mathbb{C}}(\mathfrak{Y}_{0})$, given by:
(4.4) $\widetilde{\Phi}(a,b)(z_{1},z_{2})=\varphi(z_{2}^{*}\Phi(a,b)z_{1}),\quad a,b\in\mathfrak{A},z_{1},z_{2}\in\mathfrak{Y}_{0}$
which is clearly completely positive in the sense of Definition 4.2. If $\Phi$ is positive and invariant, so is $\widetilde{\Phi}$.
If $\Phi$ is bounded, then
(4.5) $\begin{split}|\varphi(z_{2}^{*}\Phi(a,b)z_{1})|&\leq\|\varphi\|\|\Phi(a,b)\|_{\mathfrak{Y}}\|z_{2}\|_{\mathfrak{Y}_{0}}\|z_{1}\|_{\mathfrak{Y}_{0}}\\ &\leq\|\varphi\|\|\Phi\|\|a\|\|b\|\|z_{2}\|_{\mathfrak{Y}_{0}}\|z_{1}\|_{\mathfrak{Y}_{0}},\end{split}$
thus $\widetilde{\Phi}$ is bounded too. By Proposition 4.3 we deduce that there exist a Hilbert space $\mathcal{H}_{\Phi}$ (since in this case $\mathfrak{Y}=\mathbb{C}$), a closed *-representation $\pi$ of $\mathfrak{A}$ defined on a dense domain $\mathcal{D}_{\Phi}\subseteq\mathcal{H}_{\Phi}$, and a bounded linear operator $V:\mathfrak{Y}_{0}\to\mathcal{D}_{\Phi}$ such that for all $a,b\in\mathfrak{A}$ and $z_{1},z_{2}\in\mathfrak{Y}_{0}$ we have
$\langle{\pi(a)Vz_{1}}|{\pi(b)Vz_{2}}\rangle_{\mathcal{H}_{\Phi}}=\widetilde{\Phi}(a,b)(z_{1},z_{2})=\varphi(z_{2}^{*}\Phi(a,b)z_{1}).$
By the very construction of $V$ it is
$\begin{split}\|Vz\|^{2}&=\widetilde{\Phi}({\sf e},{\sf e})(z,z)=\varphi(z^{*}\Phi({\sf e},{\sf e})z)\\ &\leq\|\Phi({\sf e},{\sf e})\|_{\mathfrak{Y}}\|z\|_{\mathfrak{Y}_{0}}^{2}\|\varphi\|,\qquad\qquad\forall z\in\mathfrak{Y}_{0};\end{split}$
hence, if $\|\varphi\|\leq 1$, then $\|V\|^{2}\leq\|\Phi({\sf e},{\sf e})\|_{\mathfrak{Y}}$.
Remark 4.19.
Special cases are
  • $\mathfrak{Y}=L^{2}(\rho)$ and $\mathfrak{Y}_{0}=L^{\infty}(\rho)$, $\varphi=\rho$, with $\rho(I)=1$;
  • $\mathfrak{Y}=M(\Omega)$ and $\mathfrak{Y}_{0}=C(\Omega)$, $\varphi(\mu)=\mu(\Omega)$, for all $\mu\in M(\Omega)$.
Remark 4.20.
If $\varphi\in\mathfrak{K}^{*}$, the dual of $\mathfrak{K}$, then by the previous considerations, we deduce that there exist a Hilbert space $\mathcal{H}_{\Phi}$ a closed *-representation $\pi$ of $\mathfrak{A}$ defined on a dense domain $\mathcal{D}_{\Phi}\subseteq\mathcal{H}_{\Phi}$ and a bounded linear operator $V:\mathfrak{Y}_{0}\to\mathcal{D}_{\Phi}$ such that for all $a\in\mathfrak{A}$ and $z_{1},z_{2}\in\mathfrak{Y}_{0},$ we have
$\langle{\pi(a)Vz_{1}}|{Vz_{2}}\rangle_{\mathcal{H}_{\Phi}}=\varphi(z_{2}^{*}\Phi(a,{\sf e})z_{1})=\varphi(z_{2}^{*}\omega(a)z_{1}).$
The following two corollaries are motivated by [9, Theorem 1.3.1].
Corollary 4.21.
Let $(\mathfrak{A},\mathfrak{A}_{0})$ be a quasi *-algebra and let $\Phi:\mathfrak{A}\times\mathfrak{A}\to L^{2}(\rho)$ be a completely positive sesquilinear map. If $\|\Phi({\sf e},{\sf e})\|_{2}\leq 1$, then, for all $a\in\mathfrak{A}$, we have
$|\rho(\Phi(a,{\sf e}))|^{2}\leq\rho(\Phi(a,a)).$
Proof.
Following Remark 4.18, let $V$ be the operator constructed in Proposition 4.3. For every $a\in\mathfrak{A}$ it is
$|\rho(\Phi(a,{\sf e}))|^{2}$ $\leq|\langle{\pi(a)V(I)}|{\pi({\sf e})V(I)}\rangle|^{2}$
$\leq\|\pi(a)V(I)\|^{2}\|V(I)\|^{2}$
$\leq\|\pi(a)V(I)\|^{2}\|V\|^{2}$
$\leq\|\pi(a)V(I)\|^{2}\|\Phi({\sf e},{\sf e})\|_{2}$
$\leq\|\pi(a)V(I)\|^{2}=\rho(I\Phi(a,a)I)=\rho(\Phi(a,a)),$
by recalling that $\|V\|^{2}\leq\|\Phi({\sf e},{\sf e})\|_{2}^{2}$ by construction. ∎
Corollary 4.22.
Let $(\mathfrak{A},\mathfrak{A}_{0})$ be a quasi *-algebra and let $\Phi:\mathfrak{A}\times\mathfrak{A}\to M(\Omega)$ be a $M(\Omega)$-valued, completely positive sesquilinear map. Then, for all $a\in\mathfrak{A}$, we have
$\|\Phi(a,a)\|\|\Phi({\sf e},{\sf e})\|\geq|\Phi({\sf e},{\sf e})(\Omega)|^{2}.$
Proof.
Following Remark 4.18, let $V$ be the operator constructed in Proposition 4.3. For every $a\in\mathfrak{A}$, we have that
[…]
$=\Phi(a,a)(\Omega)\Phi({\sf e},{\sf e})(\Omega).$
4.1. Examples
We give here some examples of bounded, positive and completely positive sesquilinear maps.
Example 4.23.
Let $\mathfrak{Y}$ be either a $C^{*}$-algebra, or $L^{2}(\rho)$ or $M(\Omega)$ or $\ell_{2}(C(\Omega))$ and let $\widetilde{\Phi}:\mathfrak{A}\times\mathfrak{A}\to\mathfrak{Y}$ be a bounded left-invariant (resp. right-invariant) positive sesquilinear map. Define
$\Phi:\mathfrak{A}\times\mathfrak{A}\to\mathcal{S}_{\mathfrak{Y}}(\mathfrak{A}_{0})$
by
$\Phi(a,b)(c,d)=\widetilde{\Phi}(ac,bd),\quad\forall a,b\in\mathfrak{A},c,d\in\mathfrak{A}_{0};$
(resp. $\Phi(a,b)(c,d)=\widetilde{\Phi}(ca,db),\quad\forall a,b\in\mathfrak{A},c,d\in\mathfrak{A}_{0}$). Then $\Phi$ is completely positive.
Example 4.24.
Let $(\mathfrak{A},\mathfrak{A}_{0})$ be a quasi *-algebra. Let $\mathfrak{Y}=L^{2}(\rho)$ where $\rho$ is a finite trace on a von Neumann algebra ${\mathfrak{M}}=L^{\infty}(\rho)$. Define
$\Phi_{1}:L^{4}(\rho)\times L^{4}(\rho)\to\mathcal{S}_{L^{2}(\rho)}(L^{\infty}(\rho))$
by
$\Phi_{1}(A,B)(X,Y)=Y^{*}B^{*}AX,\quad A,B\in L^{4}(\rho),X,Y\in L^{\infty}(\rho).$
Then $\Phi$ is completely positive.
We can, in particular, fix $V\in L^{4}(\rho)$ and define
$\Phi_{2}:L^{\infty}(\rho)\times L^{\infty}(\rho)\to\mathcal{S}_{L^{2}(\rho)}(L^{\infty}(\rho))$
by
$\Phi_{2}(A,B)(X,Y)=Y^{*}V^{*}B^{*}AVX,\quad A,B,X,Y\in L^{\infty}(\rho).$
Then $\Phi_{2}$ is completely positive.
In following example we construct some $L^{2}(\rho)$-valued positive sesquilinear maps.
Example 4.25.
Let $W\in L^{\infty}(\rho)$, with $W\geq 0$. Let $k\in C([0,\|W\|]\times[0,\|W\|])$ such that $k\geq 0$. Then, for each $x\in[0,\|W\|]$, the function $\eta_{x}:[0,\|W\|]\to\mathbb{C}$ defined by $\eta_{x}(t)=k(x,t)$ is a continuous, positive function on $[0,\|W\|]$. Therefore, by the functional calculus, $\eta_{x}(W)$ defines a positive operator in $L^{\infty}(\rho)$. Let us define, for each $x\in[0,\|W\|]$ and $X,Y\in L^{2}(\rho)$,
(4.6) $\varphi(X,Y)(x)=\rho(X\eta_{x}(W)Y^{*}).$
Then, $\varphi(X,Y)\in C([0,\|W\|])$ for all $X,Y\in L^{2}(\rho)$. Moreover, $\varphi:L^{2}(\rho)\times L^{2}(\rho)\to C([0,\|W\|])$ is a bounded, left-invariant, positive sesquilinear map (see [3]).
Let $T\in L^{\infty}(\rho)$ and consider the map
$\widetilde{\varphi}:L^{2}(\rho)\times L^{2}(\rho)\to L^{\infty}(\rho)$
given by $\widetilde{\varphi}(X,Y)=T(\varphi(X,Y)(W))T^{*},\quad\forall X,Y\in L^{2}(\rho)$ by functional calculus it is also a bounded, left-invariant positive sesquilinear map.
Let $\varphi:L^{p}(\rho)\times L^{p}(\rho)\to C([0,\|W\|])$ be the sesquilinear map defined in by (4.6). We construct the following map:
$\widetilde{\Phi}:L^{p}(\rho)\times L^{p}(\rho)\to M([0,\|W\|])$
defined by
[…]
that is $d\widetilde{\Phi}(X,Y)=\varphi(X,Y)d\mu$ for all $X,Y\in L^{p}(\rho)$. Then $\widetilde{\Phi}$ is a bounded, left-invariant positive, sesquilinear map.
Some more construction can be made in the same environment: let, in fact, $\{k_{n}\}_{n\in\mathbb{N}}$ be a uniformly bounded sequence of positive functions on $[0,1]\times[0,1]$, $\{W_{n}\}_{n\in\mathbb{N}}\subset{\mathfrak{M}}$, with $W_{n}\geq 0$ and $\|W_{n}\|=1$ for all $n$. Let us define, for every $n\in\mathbb{N}$ and $X,Y\in L^{2}(\rho)$,
$\varphi_{n}(X,Y)(x)=\rho(X\eta_{x}^{(n)}(W_{n})Y^{*}),\quad x\in[0,1].$
Let $y=\{y_{n}\}_{n\in\mathbb{N}}\subseteq\ell_{2}(C([0,1]))$ such that $y_{n}\geq 0$ for all $n\in\mathbb{N}$. Define $\Phi:L^{2}(\rho)\times L^{2}(\rho)\to\ell_{2}(C([0,1]))$ by
$\Phi(X,Y)=\{\varphi_{n}(X,Y)y_{n}\}_{n\in\mathbb{N}}.$
It is straightforward to check that the map $\Phi$ is bounded, left-invariant, positive and sesquilinear.
Example 4.26.
Consider $L^{2}([0,1])$ and $M([0,1])$. Let $w$ be a positive continuous function on $[0,1]$, $\alpha$ be a homeomorphism of $[0,1]$, $\mu$ be the Lebesgue measure on $[0,1]$ and define $\Phi:L^{2}([0,1])\times L^{2}([0,1])\to M([0,1])$ by
[…]
i.e., $d\Phi(f,g)=w((\overline{g}f)\circ\alpha)d\mu$, for all $f,g\in C([0,1])$. Then $\Phi$ is completely positive.
Example 4.27.
Let $\Omega$ be a compact Hausdorff space. The space $\mathfrak{A}$ of all uniformly bounded sequences $\{f_{j}\}_{j\in\mathbb{Z}}$ in $C(\Omega)$ is a unital normed *-algebra. In fact, choose some $\{w_{j}\}_{j\in\mathbb{Z}}\in\mathfrak{A}$ such that $0\leq w_{j}\leq\frac{1}{2^{|j|}}$, for all $j\in\mathbb{Z}$ and define $\langle{\cdot}|{\cdot}\rangle:\mathfrak{A}\times\mathfrak{A}\to C(\Omega)$ by
$\langle{\{f_{j}\}_{j\in\mathbb{Z}}}|{\{f_{j}\}_{j\in\mathbb{Z}}}\rangle=\sum_{j\in\mathbb{Z}}w_{j}f_{j}\overline{g_{j}}.$
[…]
$\|\{f_{j}\}_{j\in\mathbb{Z}}\|_{\mathfrak{A}}=\|\langle{\{f_{j}\}_{j\in\mathbb{Z}}}|{\{f_{j}\}_{j\in\mathbb{Z}}}\rangle\|_{\infty}^{1/2},\quad\{f_{j}\}_{j\in\mathbb{Z}}\in\mathfrak{A}.$
Moreover, ${\mathfrak{X}}$ is a unital *-algebra with unit $(1,1,\dots,1,\dots)$, with $1$ the unit of $C(\Omega)$. Let now $\alpha$ be a homeomorphism of $\Omega$ and choose some $\{v_{j}\}_{j\in\mathbb{Z}}\in\mathfrak{A}$ such that $v_{j}\circ\alpha^{-1}\leq w_{j-1}$, for all $j\in\mathbb{Z}$. Let
$\Phi:\mathfrak{A}\times\mathfrak{A}\to\ell_{2}(C(\Omega))$
be given by
$\Phi(\{f_{j}\}_{j\in\mathbb{Z}},\{g_{j}\}_{j\in\mathbb{Z}})=\{v_{j}((f_{j-1}\overline{g_{j-1}})\circ\alpha)\}_{j\in\mathbb{Z}}=\{v_{j}(f_{j-1}\circ\alpha)(\overline{g_{j-1}}\circ\alpha)\}_{j\in\mathbb{Z}}.$
Obviously, $\Phi$ is a well-defined invariant sesquilinear map (recall that $\mathfrak{A}=\mathfrak{A}_{0}$ is a unital normed *-algebra). Moreover, $\Phi$ is bounded and completely positive. Now, notice also that $\omega:\mathfrak{A}\to\ell_{2}(C(\Omega))$ given by
$\omega(\{f_{j}\}_{j\in\mathbb{Z}}):=\{\{v_{j}(f_{j-1}\circ\alpha)\}_{j\in\mathbb{Z}},\quad\{f_{j}\}_{j\in\mathbb{Z}}\in\mathfrak{A}$
satisfies the conditions of Proposition 4.11.
Example 4.28.
Let ${\mathfrak{M}}$ be a von-Neumann algebra with a finite trace $\rho$ and ${\mathfrak{X}}$ be the space of all uniformly bounded sequences in ${\mathfrak{M}}.$ Choose some $\{W_{j}\}_{j}\in\ell_{2}({\mathfrak{M}})$ and define $\langle{\cdot}|{\cdot}\rangle:{\mathfrak{X}}\times{\mathfrak{X}}\to{\mathfrak{M}}$ by
$\langle{\{T_{j}\}_{j}}|{\{S_{j}\}_{j}}\rangle=\sum_{j}W_{j}^{*}S_{j}^{*}T_{j}W_{j}.$
[…]
$\Phi(\{T_{j}\},\{S_{j}\})=V^{*}\langle{\{T_{j}\}}|{\{S_{j}\}}\rangle V,$ then $\Phi$ is bounded, positive sesquilinear, left-invariant $L^{2}(\rho)-$valued map on $X.$
Next, choose some $\{\widetilde{W}_{j}\}_{j}\subseteq L^{4}(\rho)$ such that $\sum_{j}\|\widetilde{W}_{j}\|_{4}^{2}<\infty.$ Let $\tilde{\Phi}:{\mathfrak{X}}\times{\mathfrak{X}}\to L^{2}(\rho)$ be given by $\tilde{\Phi}(\{{T}_{j}\},\{{S}_{j}\})=\sum_{j}\widetilde{W}_{j}^{*}S_{j}^{*}T_{j}\widetilde{W}_{j}.$ Then $\sum_{j}\widetilde{W}_{j}{S}_{j}^{*}T_{j}W_{j}\in L^{2}(\rho)$. Thus, $\tilde{\Phi}$ is a bounded, left-invariant sesquilinear positive $L^{2}(\rho)-$valued map on ${\mathfrak{X}}$.
Example 4.29.
If $\mathfrak{C}$ is a unital $C^{*}$-algebra and ${\mathfrak{X}}_{\mathfrak{C}}$ is the unital Banach *-algebra of all uniformly bounded sequences in $\mathfrak{C}$ equipped with the sup norm, then if $\{W_{j}\}_{j}\in{\mathfrak{X}}_{\mathfrak{C}}$ and $B^{a}(\ell_{2}(\mathfrak{C}))$ denotes the $C^{*}$-algebra of all bounded, adjointable $\mathfrak{C}$-linear operators on $\ell_{2}(\mathfrak{C})$, then we can define $\tilde{\tilde{\Phi}}:{\mathfrak{X}}_{\mathfrak{C}}\times{\mathfrak{X}}_{\mathfrak{C}}\to B^{a}(\ell_{2}(\mathfrak{C}))$ by
$\tilde{\tilde{\Phi}}(\{T_{j}\}_{j},\{S_{j}\}_{j})(\{E_{j}\})=(\{W_{j}^{*}S_{j}^{*}T_{j}W_{j}E_{j}\}_{j}).$
For a sesquilinear map $\Theta:\mathfrak{A}\times\mathfrak{A}\to B^{a}(\ell_{2}(\mathfrak{C}))$ where $\mathfrak{A}$ is a unital normed quasi *-algebra, we shall say that $\Theta$ is completely positive if for each $N\in\mathbb{N}$ and all $a_{1},\dots,a_{N}\in\mathfrak{A}$, $x_{1},\dots,x_{N}\in\ell_{2}(\mathfrak{C})$ we have that $\sum_{i,j=1}^{N}\langle{\Theta(a_{j},a_{i})x_{j}}|{x_{i}}\rangle\in\mathfrak{C}^{+}.$ Then, $\tilde{\tilde{\Phi}}$ is completely positive.
In fact, this is a more general result, see Proposition 4.30.
Recall (see [3]) that, given a $C^{*}$-algebra $\mathfrak{C}$, a vector space ${\mathfrak{X}}_{\Theta}$, domain of a $\mathfrak{C}$-valued faithful positive sesquilinear map $\Theta:{\mathfrak{X}}_{\Theta}\times{\mathfrak{X}}_{\Theta}\to\mathfrak{C}$, is a $B_{\mathfrak{C}}$-space over a $C^{*}$-algebra $\mathfrak{C}$ if it is complete with respect to the quasi norm induced on it by the quasi inner product $\langle{a}|{b}\rangle_{\Theta}:=\Theta(a,b)$, $a,b\in{\mathfrak{X}}_{\Theta}$.
Proposition 4.30.
Let $\mathfrak{C}$ be a $C^{*}$-algebra, ${\mathfrak{X}}_{\Theta}$ be a $B_{\mathfrak{C}}$-space over a $C^{*}$-algebra $\mathfrak{C}$, $\mathcal{E}_{\Theta}$ be a dense subset of ${\mathfrak{X}}_{\Theta}$. Let $(\mathfrak{A},\mathfrak{A}_{0})$ be a quasi *-algebra and $\Phi:\mathfrak{A}\times\mathfrak{A}\to\mathcal{L}^{{\dagger}}(\mathcal{E}_{\Theta},{\mathfrak{X}}_{\Theta})$ be a bounded, left-invariant sesquilinear map. Then $\Phi$ is completely positive if and only if there exists another $B_{\mathfrak{C}}$-space $\mathcal{K}_{\Phi}$ with a dense domain $\mathcal{D}_{\Phi}$, a representation $\pi$ of $\mathfrak{A}$ in $\mathcal{L}^{{\dagger}}(\mathcal{D}_{\Phi},\mathcal{K}_{\Phi})$ and a bounded linear operation $V:\mathcal{E}_{\Theta}\to\mathcal{D}_{\Phi}$ such that for all $a,b\in\mathfrak{A}$ and $\xi,\eta\in\mathcal{E}_{\Theta}$ we have that
$\langle{\pi(a)V\xi}|{\pi(b)V\eta}\rangle_{\mathcal{K}_{\Phi}}=\langle{\Phi(a,b)\xi}|{\eta}\rangle_{\Theta}.$
Proof.
Consider the map $\tilde{\Phi}:\mathfrak{A}\times\mathfrak{A}\to S_{\mathfrak{C}}(\mathcal{E}_{\Theta})$, given by
$\tilde{\Phi}(a,b)(\xi,\eta):=\langle{\Phi(a,b)\xi}|{\eta}\rangle_{\Theta},\quad a,b\in\mathfrak{A},\xi,\eta\in\mathcal{E}_{\Theta}.$
Then $\tilde{\Phi}$ is left-invariant and sesquilinear since $\Phi$ is so, and moreover,
$\|\tilde{\Phi}(a,b)(\xi,\eta)\|\leq 4\|\Phi(a,b)\xi\|\|\eta\|\leq 4\|\Phi\|\|a\|\|b\|\|\xi\|\|\eta\|,$
so $\tilde{\Phi}$ is bounded. By Proposition 4.3, the statement follows. ∎
Remark 4.31.
If $\Phi$ in Proposition 4.30 is not bounded, but only satisfies that whenever $a,b\in\mathfrak{A}$ and $\{a_{n}\}_{n},\{b_{n}\}_{n}\subseteq\mathfrak{A}_{0}$ such that $a_{n}\to a$ and $b_{n}\to b$, then $\Phi(a_{n}-a,b_{n}-b)(\xi,\eta)\to 0$ as $n\to\infty$ for all $\xi,\eta\in\mathcal{K}$, then the Proposition 4.30 still holds because in Proposition 4.3 it is neither needed that $\Phi$ is bounded, just that the same condition holds as in [2]. Moreover, Proposition 4.30 also holds if we consider $B^{a}(\ell_{2}(\mathfrak{C})))$ instead of $\mathcal{L}^{\dagger}(\mathcal{E}_{\Theta},{\mathfrak{X}}_{\Theta})$ (where ${\mathfrak{X}}_{\Theta}$ is $B_{\mathfrak{C}}$-space and $\mathcal{E}_{\Theta}\subset{\mathfrak{X}}_{\Theta}$ densely) because $\ell_{2}(\mathfrak{C})$ is also a particular $B_{\mathfrak{C}}$-space.
4.2. Completely positive trace valued sesquilinear maps
Let ${\mathfrak{M}}$ be a von Neumann algebra and $FT({\mathfrak{M}})$ denote the space of all linear combinations of finite traces on ${\mathfrak{M}}$.
Definition 4.32.
The positive sesquilinear map $\Phi:\mathfrak{A}\times\mathfrak{A}\to FT({\mathfrak{M}})$ will be called
  • right-invariant if $\Phi(ca,d)=\Phi(c,da^{*}),\quad\forall a\in\mathfrak{A},c,d\in\mathfrak{A}_{0}$;
  • bounded if there exists a constant $M>0$ such that
    $|\Phi(a,b)(X)|\leq M\|a\|\|b\|\|X\|,\quad\forall a,b\in\mathfrak{A},X\in{\mathfrak{M}}.$
  • completely positive if for each $N\in\mathbb{N}$ and $a_{1},\dots,a_{N}\in\mathfrak{A}$, $X_{1},\dots,X_{N}\in{\mathfrak{M}}$ it is $\sum_{i,j=1}^{N}\Phi(a_{i},a_{j})(X_{j}^{*}X_{i})\geq 0$.
Proposition 4.33.
Let $(\mathfrak{A},\mathfrak{A}_{0})$ be a quasi *-algebra with unit ${\sf e}$ and $\Phi:\mathfrak{A}\times\mathfrak{A}\to FT({\mathfrak{M}})$ be a bounded right-invariant, completely positive sesquilinear map. Then, there exist a dense subspace $\mathcal{D}_{\Phi}$ of a Hilbert space $\mathcal{H}_{\Phi}$, a closed *-representation $\pi$ of $\mathfrak{A}$ in ${\mathcal{L}}^{\mbox{\scriptsize${\dagger}$}}(\mathcal{D}_{\Phi},\mathcal{H}_{\Phi})$ and a bounded linear operator $V:{\mathfrak{M}}\to\mathcal{D}_{\Phi}$ with $\|V\|^{2}\leq\Phi({\sf e},{\sf e})(I)$ such that for all $a,b\in\mathfrak{A}$ we have
$\Phi(a,b)(Y^{*}X)=\langle{\pi(a)VX}|{\pi(b)VY}\rangle_{\mathcal{H}_{\Phi}},\quad\forall a,b\in\mathfrak{A},\,\,X,Y\in{\mathfrak{M}}.$
Proof.
Let $\mathcal{S}_{\mathbb{C}}({\mathfrak{M}})$ be the space of all sesquilinear forms on ${\mathfrak{M}}$ and $\widetilde{\Phi}:\mathfrak{A}\times\mathfrak{A}\to\mathcal{S}_{\mathbb{C}}({\mathfrak{M}})$ be given by:
$\widetilde{\Phi}(a,b)(X,Y)=\Phi(a,b)(Y^{*}X),\quad\forall a,b\in\mathfrak{A},\,\,X,Y\in{\mathfrak{M}}.$
Then, $\widetilde{\Phi}$ is a right-invariant, completely positive sesquilinear form on $\mathbb{C}$) since inherits these properties from $\Phi$. Moreover $\widetilde{\Phi}$ is bounded. By Proposition 4.3, there exist a Hilbert space $\mathcal{H}_{\widetilde{\Phi}}$ (since in this case $\mathfrak{Y}=\mathbb{C}$), a dense subspace $\mathcal{D}_{\widetilde{\Phi}}$ of $\mathcal{H}_{\widetilde{\Phi}}$, a bounded linear operator $V:{\mathfrak{M}}\to\mathcal{D}_{\widetilde{\Phi}}$ and a closed *-representation $\pi$ of $\mathfrak{A}$ such that for all $a,b\in\mathfrak{A}$ we have
$\Phi(a,b)(Y^{*}X)=\langle{\pi(a)VX}|{\pi(b)VY}\rangle_{\mathcal{H}_{\widetilde{\Phi}}},\quad\forall a,b\in\mathfrak{A},\,\,X,Y\in{\mathfrak{M}}.$
Moreover, by construction, it is
$\|VX\|^{2}\leq|\widetilde{\Phi}({\sf e},{\sf e})(X,X)|=|\Phi({\sf e},{\sf e})(X^{*}X)|\leq\Phi({\sf e},{\sf e})(I)\|X\|^{2}$
for all $X\in{\mathfrak{M}}$, hence $\|V\|^{2}\leq\Phi({\sf e},{\sf e})(I)$. ∎
Corollary 4.34.
Let $(\mathfrak{A},\mathfrak{A}_{0})$ be a quasi *-algebra with unit ${\sf e}$ and $\Phi:\mathfrak{A}\times\mathfrak{A}\to FT({\mathfrak{M}})$ be a bounded right-invariant, completely positive sesquilinear map. Then, for all $a\in\mathfrak{A}$,
$|\Phi(a,{\sf e})(I)|^{2}\leq\Phi(a,a)(I)\Phi({\sf e},{\sf e})(I).$
Example 4.35.
Let $\rho$ be a finite trace on the Von Neumann algebra ${\mathfrak{M}}=L^{\infty}(\rho)$ and let $\mathfrak{A}=L^{2}(\rho)$. For all $A,B\in L^{2}(\rho)$ define
$\Phi(A,B)(X):=\rho(B^{*}XA),\forall X\in L^{\infty}(\rho).$
By simple calculations it can be shown that $\Phi$ is a right-invariant, bounded, completely positive sesquilinear map and, considered the dense subspace $\mathcal{D}=L^{\infty}(\rho)\subseteq\mathcal{H}=L^{2}(\rho)$, and $V=I_{\mathcal{D}}$, the *-representation $\pi$ of $L^{2}(\rho)$ is $\pi(A)(Y)=YA$, for all $A\in L^{2}(\rho)$, $Y\in L^{\infty}(\rho)$.
Acknowledgements: GB and CT acknowledge that this work has been carried out within the activities of Gruppo UMI Teoria dell’Appros-simazione e Applicazioni and of GNAMPA of the INdAM. SI is supported by the Ministry of Science, Technological Development and Innovations, Republic of Serbia, grant no. 451-03-66/2024-03/200029.