### v3

# Representations induced by positive and completely positive sesquilinear maps

Abstract.
The representations that can be defined starting from positive sesquilinear maps $\Phi$ with values in an ordered Banach bimodule are essentially of two types: if $\Phi$ is positive, a natural generalization follows the path of the GNS construction; if $\Phi$ is completely positive, one can move toward a possible generalization of the Stinespring dilation theorem. Both possibilities are explored in this paper providing representations of positive and completely positive ordered Banach bimodule-valued sesquilinear maps on a quasi *-algebra.
Key words and phrases:
Representations, modules, positive sesquilinear maps, completely positive sesquilinear maps, quasi normed spaces.
1991 Mathematics Subject Classification:
Primary 46K10, 47A07, 16D10.
1. Introduction
Positive and completely positive maps play an important role in the theory of operator algebras and in the related fields of dilation theory, random matrix theory, free probability, semidefinite optimization theory and quantum information. Another area where positive maps play an important role is the one of operator spaces, see [17, Introduction]. These facts motivate a study of representations involving positive and completely positive maps. A classical result in representation theory is the famous Stinespring theorem [16] (see also [2, 13, 17] which gives a representation of completely positive maps from a $C^{*}$-algebra into the space of bounded operators in Hilbert space. Quite recently, in [5, Corollary 3.10] and [6, Corollary 3.3], a representation of general positive $C^{*}$-valued maps on arbitrary *-algebras has also been provided. However, it is now long time that the $C^{*}$-algebraic approach to quantum theories has been considered as a too rigid framework where casting all objects of physical interest. Therefore, several possible generalizations have been proposed: quasi *-algebras, partial *-algebras and so on. Further, quasi *-algebras, and in particular Banach quasi *-algebras (a special case of Banach modules over *-algebras), are also interesting from the mathematical point of view because several beautiful and important mathematical structures (such as commutative and noncommutative $L^{p}$-spaces) are examples of Banach quasi *-algebras. Hilbert $C^{*}$-modules [11, 12] and the dual of a unital $C^{*}$-algebra provide other examples of interest for applications. All these facts motivated us to study positive and completely positive maps from a Banach quasi *-algebra into an (ordered) Banach bimodule over a *-algebra, and the main aim of this paper is therefore obtaining representations of such maps. In a recent paper [5], we considered a GNS-like construction defined by positive sesquilinear maps on a quasi *-algebra taking their values in a $C^{*}$-algebra. In this paper we go some steps further and consider the possibility of replacing sesquilinear maps taking values in a $C^{*}$-algebra with sesquilinear maps with values in an ordered Banach bimodule $\mathfrak{Y}$ over some *-algebra $\mathfrak{Y}_{0}$. In this situation, an extension of the result of [5] is possible, provided that the sesquilinear form $\Phi$, where we start from, satisfies a generalized Cauchy-Schwarz inequality. This discussion is developed in Section 3 where we characterize a class of ordered Banach bimodules $\mathfrak{Y}$ with the property that every $\mathfrak{Y}$-valued positive sesquilinear map satisfies a generalized Cauchy-Schwarz inequality (concrete examples of which are noncommutative $L^{2}$-spaces and duals of unital $C^{*}$-algebras) and then obtain a representation of positive sesquilinear maps from a quasi *-algebra into such an ordered Banach bimodule.
Section 4 is devoted to the *-representations induced by completely positive sesquilinear maps. This is actually a subject for which several approaches have been proposed in view of a possible generalization of the Stinespring theorem to different environments (see, for instance, [3, 8, 9, 10, 14]). In the very same way, in our framework, several definitions of complete positivity can be introduced. First, motivated by [14, Proposition 3.1], following an idea proposed in [15, Ch. 11] and adopted in [4] for partial *-algebras, we consider completely positive sesquilinear maps from a normed quasi *-algebra $\mathfrak{A}$ into a set of sesquilinear maps on a vector space which take values in an ordered Banach bimodule $\mathfrak{Y}$ and provide a generalization of the classical Stinespring theorem. Further, we provide a version of Radon-Nikodym theorem for these maps as well as for ordered Banach bimodule-valued bounded positive sesquilinear maps on a quasi *-algebra, thus generalizing, in this setting, the results from [17, Section 3.5]. At the end of the section, we apply our results to obtain representations of some new classes of completely positive maps; in particular, we provide a representation of a certain class of bounded sesquilinear maps from a quasi *-algebra into an ordered Banach bimodule, then we give a representation of bounded completely positive sesquilinear maps on a quasi *-algebra with values in the space of ordered Banach bimodule-valued, bounded, linear operators. Moreover, as a further application of the results of Section 3, we obtain some inequalities of these maps and give some examples.
2. Preliminaries
A quasi *-algebra $(\mathfrak{A},\mathfrak{A}_{0})$ is a pair consisting of a vector space $\mathfrak{A}$ and a *-algebra $\mathfrak{A}_{0}$ contained in $\mathfrak{A}$ as a subspace and such that
  • $\mathfrak{A}$ carries an involution $a\mapsto a^{*}$ extending the involution of $\mathfrak{A}_{0}$;
  • $\mathfrak{A}$ is a bimodule over $\mathfrak{A}_{0}$ and the module multiplications extend the multiplication of $\mathfrak{A}_{0}$. In particular, the following associative laws hold:
    $(ca)d=c(ad);\ \ a(cd)=(ac)d,\quad\forall\ a\in\mathfrak{A},\ c,d\in\mathfrak{A}_{0};$
  • $(ac)^{*}=c^{*}a^{*}$, for every $a\in\mathfrak{A}$ and $c\in\mathfrak{A}_{0}$.
The identity or unit element of $(\mathfrak{A},\mathfrak{A}_{0})$, if any, is a necessarily unique element ${\sf e}\in\mathfrak{A}_{0}$, such that $a{\sf e}=a={\sf e}a$, for all $a\in\mathfrak{A}$.
We will always suppose that
$ac=0,\;\forall c\in\mathfrak{A}_{0}\Rightarrow a=0$
$ac=0,\;\forall a\in\mathfrak{A}\Rightarrow c=0.$
Clearly, both these conditions are automatically satisfied if $(\mathfrak{A},\mathfrak{A}_{0})$ has an identity ${\sf e}$.
[…]
  • there exists $\gamma>0$ such that for every $a\in\mathfrak{A}$
    $\max\{\|ac\|,\|ca\|\}\leq\gamma\|a\|,\quad\forall c\in\mathfrak{A}_{0};$
  • $\|a^{*}\|=\|a\|,\;\forall a\in\mathfrak{A};$
  • $\mathfrak{A}_{0}$ is dense in $\mathfrak{A}[\|\cdot\|]$.
If the normed vector space $\mathfrak{A}[\|\cdot\|]$ is complete, then $(\mathfrak{A},\mathfrak{A}_{0})$ is called a Banach quasi *-algebra. We refer to [7] for further details.
Definition 2.1.
Let $\mathfrak{Y}$ be a Banach bimodule over the *-algebra $\mathfrak{Y}_{0}$. We say that $\mathfrak{Y}$ is an ordered Banach bimodule over $\mathfrak{Y}_{0}$ if
  • (i)
    $\mathfrak{Y}$ is ordered as a vector space; that is, $\mathfrak{Y}$ contains a (positive) closed cone $\mathfrak{K}$, i.e., $\mathfrak{K}\subset\mathfrak{Y}$ is such that $\mathfrak{K}+\mathfrak{K}\subset\mathfrak{K}$, $\lambda\mathfrak{K}\subset\mathfrak{K}$ for $\lambda\geq 0$ and $\mathfrak{K}\cap(-\mathfrak{K})=\{0\}$;
  • (ii)
    $z^{*}\mathfrak{K}z\subset\mathfrak{K},\quad\forall z\in\mathfrak{Y}_{0}.$
As usual, we will write $y_{1}\leq y_{2}$ whenever $y_{2}-y_{1}\in\mathfrak{K}$, with $y_{1},y_{2}\in\mathfrak{Y}$ and we will sometimes suppose that $\mathfrak{Y}$ has an order-preserving norm in the sense that if $y_{1}\leq y_{2}$ with $y_{1},y_{2}\in\mathfrak{Y}$, then also $\|y_{1}\|_{\mathfrak{Y}}\leq\|y_{2}\|_{\mathfrak{Y}}$.
Example 2.2.
Examples of ordered Banach bimodules over a Banach *-algebra are provided by:
  • the non commutative spaces $\mathfrak{Y}=L^{2}(\rho)$, $\mathfrak{Y}_{0}=L^{\infty}(\rho)$, with $\rho$ a finite trace on a von Neumann algebra ${\mathfrak{M}}$;
[…]
  • $\mathfrak{Y}=\mathfrak{Y}_{0}=\ell_{2}(C(\Omega))$, with $\ell_{2}(C(\Omega))$ the space of squared summable sequences of functions in $C(\Omega)$ with $\Omega$ a compact Hausdorff space. It is a Banach bimodule over itself because it is a (non-unital) normed *-algebra.
The definitions of the positive cones are obvious in all these cases.
Let ${\mathfrak{X}}$ be a vector space and $\mathfrak{Y}$ an ordered Banach bimodule over the *-algebra $\mathfrak{Y}_{0}$, with positive cone $\mathfrak{K}$. Let $\Phi$ be a $\mathfrak{Y}$-valued positive sesquilinear map on ${\mathfrak{X}}\times{\mathfrak{X}}$
$\Phi:(x_{1},x_{2})\in{\mathfrak{X}}\times{\mathfrak{X}}\to\Phi(x_{1},x_{2})\in\mathfrak{Y}$
i.e., a map with the properties
  • $i)$
    $\Phi(x_{1},x_{1})\in\mathfrak{K}$,
  • $ii)$
    $\Phi(\alpha x_{1}+\beta x_{2},\gamma x_{3})=\overline{\gamma}[\alpha\Phi(x_{1},x_{3})+\beta\Phi(x_{2},x_{3})]$,
with $x_{1},x_{2},x_{3}\in{\mathfrak{X}}$ and $\alpha,\beta,\gamma\in\mathbb{C}$.
The $\mathfrak{Y}$-valued positive sesquilinear map $\Phi$ is called faithful if
$\Phi(x,x)=0_{\mathfrak{Y}}\;\Rightarrow\;x=0.$
3. Representations induced by positive sesquilinear maps
It is quite apparent that the cornerstone of every extension of the GNS-construction requires a Cauchy-Schwarz inequality or, at least, some generalization of it. We begin with presenting some results in this direction.
Throughout this section we will assume that $\mathfrak{Y}$ is (an ordered) Banach bimodule over a *-algebra $\mathfrak{Y}_{0}$ (with $\mathfrak{Y}_{0}$ equipped with a not necessarily sub-multiplicative norm $\|\cdot\|_{\mathfrak{Y}_{0}}$) with the respective cones $\mathfrak{K}$ and $\mathfrak{K}_{0}$, with $\mathfrak{K}_{0}=\{\sum_{i=1}^{N}z_{i}^{*}z_{i},\,z_{i}\in\mathfrak{Y}_{0},i=1,\dots,N;N\in\mathbb{N}\}$ and satisfying the following properties:
  • $(D1)$
    Every $z\in\mathfrak{Y}_{0}$ can be written as $z=z_{1}-z_{2}+i(z_{3}-z_{4})$, where $z_{j}\in\mathfrak{K}_{0}$ and $\|z_{j}\|_{\mathfrak{Y}_{0}}\leq\|z\|_{\mathfrak{Y}_{0}}$ for all $j\in\{1,\dots,4\}.$
  • $(D2)$
    There exists a family $\mathcal{F}$ of sesquilinear forms $\varphi:\mathfrak{Y}\times\mathfrak{Y}_{0}\to\mathbb{C}$ enjoying the following properties:
    • i)
      for every $y^{+}\in\mathfrak{K}$ and $z^{+}\in\mathfrak{K}_{0}$, $\varphi(y^{+},z^{+})\geq 0$ for all $\varphi\in\mathcal{F}$;
    • ii)
      for every $y\in\mathfrak{Y}$ and $z\in\mathfrak{Y}_{0}$,
      $|\varphi(y,z)|\leq\|y\|_{\mathfrak{Y}}\|z\|_{\mathfrak{Y}_{0}},\quad\forall\varphi\in\mathcal{F};$
    • iii)
      for every $y\in\mathfrak{Y}$, the norm in $\mathfrak{Y}$
      $\|y\|_{\mathfrak{Y}}=\sup_{\begin{subarray}{c}z\in\mathfrak{Y}_{0}\\ \|z\|_{\mathfrak{Y}_{0}}\leq 1\end{subarray}}\sup_{\varphi\in\mathcal{F}}|\varphi(y,z)|.$
Under these assumptions, we can prove that the following Cauchy-Schwarz-like inequality for positive $\mathfrak{Y}$-valued sesquilinear maps holds.
Proposition 3.1.
Let ${\mathfrak{X}}$ be a complex vector space, $\mathfrak{Y}$ an ordered Banach bimodule over the *-algebra $\mathfrak{Y}_{0}$, with positive closed cone $\mathfrak{K}$, satisfying $(D1)-(D2)$. Let $\Phi:{\mathfrak{X}}\times{\mathfrak{X}}\to\mathfrak{Y}$ be a positive sesquilinear map. Then, for all $x_{1},x_{2}\in{\mathfrak{X}}$ we have that
(3.1) $\|\Phi(x_{1},x_{2})\|_{\mathfrak{Y}}\leq 4\|\Phi(x_{1},x_{1})\|_{\mathfrak{Y}}^{1/2}\|\Phi(x_{2},x_{2})\|_{\mathfrak{Y}}^{1/2}.$
Proof.
Let $\mathcal{F}$ be the family in $(D2)$. For each fixed $\varphi\in\mathcal{F}$ and $z^{+}\in\mathfrak{K}_{0}$, let $\widetilde{\varphi}_{z^{+}}$ be the sesquilinear form on ${\mathfrak{X}}\times{\mathfrak{X}}$ given by
$\widetilde{\varphi}_{z^{+}}(x_{1},x_{2})=\varphi(\Phi(x_{1},x_{2}),z^{+}),\quad\forall x_{1},x_{2}\in{\mathfrak{X}}.$
Since $\Phi$ is positive, by $(D2)$ $i)$, we obtain that $\widetilde{\varphi}_{z^{+}}$ is positive. Hence, by $(D2)$ $ii)$, for all $x_{1},x_{2}\in{\mathfrak{X}}$ we get that
(3.2) $|\widetilde{\varphi}_{z^{+}}(x_{1},x_{1})||\widetilde{\varphi}_{z^{+}}(x_{2},x_{2})|\leq\|\Phi(x_{1},x_{1})\|_{\mathfrak{Y}}\|\Phi(x_{2},x_{2})\|_{\mathfrak{Y}}\|z^{+}\|_{\mathfrak{Y}_{0}}^{2}.$
Now, if we choose $z\in\mathfrak{Y}_{0}$, by $(D1)$, we can find $z_{1}^{+},z_{2}^{+},z_{3}^{+},z_{4}^{+}\in\mathfrak{K}_{0}$ such that $z=z_{1}^{+}-z_{2}^{+}+i(z_{3}^{+}-z_{4}^{+})$ and $\|z_{j}^{+}\|_{\mathfrak{Y}_{0}}\leq\|z\|_{\mathfrak{Y}_{0}}$ for all $j\in\{1,...,4\}.$ Hence, by the Cauchy-Schwarz inequality applied to the positive sesquilinear forms $\widetilde{\varphi}_{z_{j}^{+}}$ and by (3.2), it is
$|\varphi(\Phi(x_{1},x_{2}),z)|$ $\leq\sum_{j=1}^{4}|\varphi(\Phi(x_{1},x_{2}),z_{j}^{+})|$
$\leq\sum_{j=1}^{4}\|\Phi(x_{1},x_{1})\|_{\mathfrak{Y}}^{1/2}\|\Phi(x_{2},x_{2})\|_{\mathfrak{Y}}^{1/2}\|z_{j}^{+}\|_{\mathfrak{Y}_{0}}$
$\leq 4\|\Phi(x_{1},x_{1})\|_{\mathfrak{Y}}^{1/2}\|\Phi(x_{2},x_{2})\|_{\mathfrak{Y}}^{1/2}\|z\|_{\mathfrak{Y}_{0}}$
for all $x_{1},x_{2}\in{\mathfrak{X}}$ and all $\varphi\in\mathcal{F}$. Therefore, for every $z\in\mathfrak{Y}_{0}$, with $\|z\|_{\mathfrak{Y}_{0}}\leq 1$ and all $x_{1},x_{2}\in{\mathfrak{X}}$, we have
$\sup_{\varphi\in\mathcal{F}}|\varphi(\Phi(x_{1},x_{2}),z)|\leq 4\|\Phi(x_{1},x_{1})\|_{\mathfrak{Y}}^{1/2}\|\Phi(x_{2},x_{2})\|_{\mathfrak{Y}}^{1/2}.$
However, by $(D2)$ $iii)$, we deduce that for all $x_{1},x_{2}\in{\mathfrak{X}}$
$\|\Phi(x_{1},x_{2})\|_{\mathfrak{Y}}$ $=\sup_{\begin{subarray}{c}z\in\mathfrak{Y}_{0}\\ \|z\|_{\mathfrak{Y}_{0}}\leq 1\end{subarray}}\sup_{\varphi\in\mathcal{F}}|\varphi(\Phi(x_{1},x_{2}),z)|$
$\leq 4\|\Phi(x_{1},x_{1})\|_{\mathfrak{Y}}^{1/2}\|\Phi(x_{2},x_{2})\|_{\mathfrak{Y}}^{1/2}.$
We now give a series of examples of ordered Banach bimodules $\mathfrak{Y}$ over *-algebras $\mathfrak{Y}_{0}$ and of families $\mathcal{F}$ of sesquilinear forms on $\mathfrak{Y}\times\mathfrak{Y}_{0}$ for which conditions $(D1)-(D2)$ hold.
Example 3.2.
Let $\rho$ be a finite trace on a von Neumann algebra $\mathfrak{M}$. Put $\mathfrak{Y}=L^{2}(\rho)$, $\mathfrak{Y}_{0}=L^{\infty}(\rho)$ and let $\mathfrak{K}$ be the natural cone of $\mathfrak{Y}$ consisting of positive elements and $\mathfrak{K}_{0}=L^{\infty}(\rho)\cap\mathfrak{K}$. In this case, we will consider $\mathfrak{Y}_{0}$ as a *-algebra equipped with the norm $\|\cdot\|_{2}$ (instead of $\|\cdot\|_{\infty}$). Let $\varphi:\mathfrak{Y}\times\mathfrak{Y}_{0}\to\mathbb{C}$ be defined by
$\varphi(Y,W)=\rho(W^{*}Y),\quad Y\in\mathfrak{Y},W\in\mathfrak{Y}_{0}$
and define =$\mathcal{F}=\{\varphi\}$, a family with the single sesquilinear form $\varphi$. By some calculations it is not hard to check that $(D1)-(D2)$ are satisfied. If $\rho$ is not finite, the statement of Proposition 3.1 remains true; indeed, the condition $(D2)$ $i)$ is fulfilled by using Lemma 4.8 in [6]; the other assumptions do not depend on the finiteness of the trace.
Example 3.3.
Let $\mathfrak{Y}_{0}$ be a unital $C^{*}$-algebra and $\mathfrak{Y}:=\mathfrak{Y}_{0}^{*}$ be its dual. If $\phi\in\mathfrak{Y}$ and $z\in\mathfrak{Y}_{0}$, the module multiplication can be defined as
$(\phi\cdot z)(w)=\phi(zw)\mbox{ and }(z\cdot\phi)(w)=\phi(wz),\quad w\in\mathfrak{Y}_{0}.$
Let $\mathfrak{K}_{0}$ be the natural cone in $\mathfrak{Y}_{0}$ consisting of the positive elements in $\mathfrak{Y}_{0}$ and the cone $\mathfrak{K}$ in $\mathfrak{Y}$ be the dual cone i.e., the set consisting of all $\phi\in\mathfrak{Y}$ such that $\phi(z)\geq 0$ whenever $z\in\mathfrak{K}_{0}$. We can take as
$\mathcal{F}$ the set consisting of the single sesquilinear form $\varphi:\mathfrak{Y}\times\mathfrak{Y}_{0}\to\mathbb{C}$ given by
$\varphi(\phi,z)=\phi(z^{*}),\quad\phi\in\mathfrak{Y},z\in\mathfrak{Y}_{0}.$
Example 3.4.
Let $\mathfrak{Y}=\mathfrak{Y}_{0}=\ell_{2}(C(\Omega))$ (see Example 2.2), we can take as
$\mathfrak{K}=\{\{f_{n}\}_{n}\in\ell_{2}(C(\Omega))\,|\,f_{n}\geq 0,\,\forall n\in\mathbb{N}\}$
and =$\mathcal{F}=\{\varphi_{t}|\,t\in\Omega\}$, with $\varphi_{t}:\ell_{2}(C(\Omega))\times\ell_{2}(C(\Omega))\to\mathbb{C}$ defined by
$\varphi_{t}(\{f_{n}\}_{n},\{g_{n}\}_{n})=\langle{\{g_{n}\}_{n}}|{\{f_{n}\}_{n}}\rangle_{2}(t).$
Remark 3.5.
In some cases, given a positive sesquilinear map $\Phi:{\mathfrak{X}}\times{\mathfrak{X}}\to\mathfrak{Y}$, the inequality
(3.3) $\|\Phi(x_{1},x_{2})\|_{\mathfrak{Y}}\leq\beta\|\Phi(x_{1},x_{1})\|_{\mathfrak{Y}}^{1/2}\|\Phi(x_{2},x_{2})\|_{\mathfrak{Y}}^{1/2},$
with $x_{1},x_{2}\in{\mathfrak{X}}$, $\beta>0$, holds with $\beta$ smaller than $4$ as in (3.1). Indeed, following the proof of Proposition 3.1,
  • if $\mathfrak{Y}=L^{2}(\rho)$, $\mathfrak{Y}_{0}=L^{\infty}(\rho)$, with $\rho$ a (semi)finite trace on a von Neumann algebra ${\mathfrak{M}}$, then we can take $\beta=\sqrt{2}$, this can be proved by using the decomposition of every element $Z$ of a von Neumann algebra into $Z=(Z_{+}^{(1)}-Z_{-}^{(1)})+i(Z_{+}^{(2)}-Z_{-}^{(2)}),$ with $Z_{+}^{(j)},Z_{-}^{(j)}\geq 0$ and $Z_{+}^{(j)}Z_{-}^{(j)}=Z_{-}^{(j)}Z_{+}^{(j)}=0$, $j=1,2$ and the properties of the trace. Indeed, in this case, $\rho(Z^{*}Z)=\sum_{j=1}^{2}[\rho((Z_{+}^{(j)})^{2})+\rho((Z_{-}^{(j)})^{2})],$ hence $\sum_{j=1}^{2}(\|Z_{+}^{(j)}+Z_{-}^{(j)}\|_{2})\leq\sqrt{2}\|Z\|_{2}$. Therefore, by making some calculations using the positive sesquilinear form $\varphi(\Phi(x_{1},x_{2}),Z)=\rho(Z^{*}\Phi(x_{1},x_{2}))$, $x_{1},x_{2}\in{\mathfrak{X}}$, $Z\in L^{2}(\rho)\cap L^{\infty}(\rho)$, $Z\geq 0$, we obtain the constant $\beta=\sqrt{2}$ in (3.3).
  • if $\mathfrak{Y}=L^{2}(\Omega)$, with $\Omega$ a compact or locally compact Hausdorff space, we can take $\beta=1$ i.e., the Cauchy-Schwarz inequality holds by considering the integral of both sides of the Cauchy-Schwarz inequality for the positive sesquilinear form $\varphi_{t}(x_{1},x_{2})=\Phi(x_{1},x_{2})(t)$, $x_{1},x_{2}\in{\mathfrak{X}}$, $t\in\Omega$ and then applying the Cauchy-Schwarz to the inner product in $L^{2}(\Omega)$.
  • if $\mathfrak{Y}_{0}$ is a unital $C^{*}$-algebra and $\mathfrak{Y}=\mathfrak{Y}_{0}^{*}$ is its dual, then we can take $\beta=2$; this is due to the fact that every element $z\in\mathfrak{Y}_{0}$ can be decomposed as $z=(z_{1}-z_{2})+i(z_{3}-z_{4}),$ with $z_{i}\geq 0$ and $z_{i}z_{j}=z_{j}z_{i}=0$, $\|z_{i}-z_{j}\|=\|z_{j}+z_{i}\|$ when $i,j=1,2$ and $i,j=3,4$, and by applying property $(D2)$ $ii)$ to $\Phi(x_{1},x_{2})(z)$, with $x_{1},x_{2}\in{\mathfrak{X}}$ and $z\in\mathfrak{Y}_{0}$. Moreover, in the commutative case, then $\mathfrak{Y}_{0}=C(\Omega)$ and $\mathfrak{Y}=M(\Omega)$ we can take $\beta=1$ since we can apply Cauchy-Schwarz inequality to every positive sesquilinear form $\varphi_{E}(X,Y)=\Phi(X,Y)(E)$, with $E$ a given Borel subset of $\Omega$, and therefore by considering a disjoint partition $\{E_{n}\}_{n=1}^{N}$ of $\Omega$ we obtain
    $\sum_{n=1}^{N}|\Phi(X,Y)(E_{n})|\leq$
    $\left(\sum_{n=1}^{N}|\Phi(X,X)(E_{n})|\right)^{1/2}\left(\sum_{n=1}^{N}|\Phi(Y,Y)(E_{n})|\right)^{1/2};$
    hence by taking the supremum of both sides we get the Cauchy-Schwarz inequality;
  • if $\mathfrak{Y}=\mathfrak{Y}_{0}=\ell_{2}(C(\Omega))$, with $\Omega$ a compact Hausdorff space, we can take $\beta=1$, indeed we can apply the Cauchy-Schwarz inequality to all the positive sesquilinear forms $\varphi_{n,t}(x_{1},x_{2})=(\Phi(x_{1},x_{2}))_{n}(t)$ with $(\Phi(x_{1},x_{2}))_{n}$ the n-th component of the sequence $\Phi(x_{1},x_{2})$ and then applying the Cauchy-Schwarz inequality for the inner product in $\ell_{2}(C(\Omega))$.
Definition 3.6.
Let ${\mathfrak{X}}$ be a vector space. A $\mathfrak{Y}$-valued faithful positive sesquilinear map $\Phi$ on ${\mathfrak{X}}\times{\mathfrak{X}}$ is said to be a $\mathfrak{Y}$-valued quasi inner product on ${\mathfrak{X}}$ and we often write $\langle{x_{1}}|{x_{2}}\rangle_{\Phi}:=\Phi(x_{1},x_{2})$, $x_{1},x_{2}\in{\mathfrak{X}}$.
A $\mathfrak{Y}$-valued quasi inner product on ${\mathfrak{X}}$ $\Phi:{\mathfrak{X}}\times{\mathfrak{X}}\to\mathfrak{Y}$ induces a quasi norm $\|\cdot\|_{\Phi}$ on ${\mathfrak{X}}$:
$\|x\|_{\Phi}:=\sqrt{\|\langle{x}|{x}\rangle_{\Phi}\|_{\mathfrak{Y}}}=\sqrt{\|\Phi(x,x)\|_{\mathfrak{Y}}},\quad x\in{\mathfrak{X}},$
since
$\|x_{1}+x_{2}\|_{\Phi}\leq 2(\|x_{1}\|_{\Phi}+\|x_{2}\|_{\Phi}),\qquad\forall x_{1},x_{2}\in{\mathfrak{X}}$
(this can be shown as in [5]).
The space ${\mathfrak{X}}$ is then a quasi normed space w.r.to the quasi norm $\|\cdot\|_{\Phi}$.
Definition 3.7.
Let ${\mathfrak{X}}$ be a complex vector space and $\Phi$ be a $\mathfrak{Y}$-valued quasi inner product on ${\mathfrak{X}}$. If ${\mathfrak{X}}$ is complete w.r. to the quasi norm $\|\cdot\|_{\Phi}$, then ${\mathfrak{X}}$ is called a quasi Banach space with $\mathfrak{Y}$-valued quasi inner product or for short a quasi $B_{\mathfrak{Y}}$-space.
If $\Phi$ is not faithful, we can consider the subspace of ${\mathfrak{X}}$
$\mathfrak{N}_{\Phi}=\{x_{1}\in{\mathfrak{X}}:\,\Phi(x_{1},x_{2})=0_{\mathfrak{Y}},\forall x_{2}\in{\mathfrak{X}}\}.$
If a generalized Cauchy-Schwarz inequality holds, then
$\mathfrak{N}_{\Phi}=\{x\in{\mathfrak{X}}:\,\Phi(x,x)=0_{\mathfrak{Y}}\}.$
We denote by $\Lambda_{\Phi}(x)$ the coset of ${\mathfrak{X}}/\mathfrak{N}_{\Phi}$ containing $x\in{\mathfrak{X}}$; i.e., $\Lambda_{\Phi}(x)=x+\mathfrak{N}_{\Phi}$ and define a $\mathfrak{Y}$-valued quasi inner product on ${\mathfrak{X}}/\mathfrak{N}_{\Phi}$ as follows:
$\langle{\Lambda_{\Phi}(x_{1})}|{\Lambda_{\Phi}(x_{2})}\rangle_{\Phi}:=\Phi(x_{1},x_{2}),\quad x_{1},x_{2}\in{\mathfrak{X}}.$
The associated quasi norm is:
$\|\Lambda_{\Phi}(x)\|_{\Phi}:=\sqrt{\|\Phi(x,x)\|_{\mathfrak{Y}}},\quad x\in{\mathfrak{X}}.$
The quotient space ${\mathfrak{X}}/\mathfrak{N}_{\Phi}=\Lambda_{\Phi}({\mathfrak{X}})$ is a quasi normed space (see [5]).
Remark 3.8.
In the cases when $\beta$ in (3.3) can be taken as $1$, the space ${\mathfrak{X}}$ (or $\Lambda_{\Phi}({\mathfrak{X}})$ if $\Phi$ is not faithful) is a normed space, hence, if it is also complete, it is a Banach space.
Let $\mathcal{K}$ be a quasi $B_{\mathfrak{Y}}$-space and $D(T)$ be a dense subspace of $\mathcal{K}$. A linear map $T:D(T)\to\mathcal{K}$ is said $\Phi$-adjointable if there exists a linear map $T^{*}$ defined on a subspace $D(T^{*})\subset\mathcal{K}$ such that
$\Phi(T\xi,\eta)=\Phi(\xi,T^{*}\eta),\quad\forall\xi\in D(T),\eta\in D(T^{*}).$
Let $\mathcal{D}$ be a dense subspace of $\mathcal{K}$ and let us consider the following families of linear operators acting on $\mathcal{D}$:
${{\mathcal{L}}^{\mbox{\scriptsize${\dagger}$}}(\mathcal{D},\mathcal{K})}$ =$=\{T\mbox{ $\Phi$-adjointable},D(T)=\mathcal{D};\;D(T^{*})\supset\mathcal{D}\}$
${{\mathcal{L}}^{\dagger}(\mathcal{D})}$ =$=\{T\in{\mathcal{L}}^{\mbox{\scriptsize${\dagger}$}}(\mathcal{D},\mathcal{K}):T\mathcal{D}\subset\mathcal{D};\;T^{*}\mathcal{D}\subset\mathcal{D}\}.$
The involution in ${\mathcal{L}}^{\mbox{\scriptsize${\dagger}$}}(\mathcal{D},\mathcal{K})$ is defined by $T^{\dagger}:=T^{*}\upharpoonright\mathcal{D}$, the restriction of $T^{*}$, the $\Phi$-adjoint of $T$, to $\mathcal{D}$. The set ${\mathcal{L}}^{\dagger}(\mathcal{D})$ is a *-algebra.
Remark 3.9.
[…]
$\Phi(T_{2}\xi,T_{1}^{\dagger}\eta)=\Phi(W\xi,\eta),\quad\forall\xi,\eta\in\mathcal{D}.$
Due to the density of $\mathcal{D}$ in $\mathcal{K}$, the element $W$, if it exists, is unique. We put $W=T_{1}\,{\scriptstyle\square}\,T_{2}$.
Definition 3.10.
Let $(\mathfrak{A},\mathfrak{A}_{0})$ be a quasi *-algebra with unit ${\sf e}$. Let $\mathcal{D}$ be a dense subspace of a certain quasi $B_{\mathfrak{Y}}$-space $\mathcal{K}$ with $\mathfrak{Y}$-valued quasi inner product $\langle{\cdot}|{\cdot}\rangle_{\mathcal{K}}$. A linear map $\pi$ from $\mathfrak{A}$ into ${\mathcal{L}}^{\mbox{\scriptsize${\dagger}$}}(\mathcal{D},\mathcal{K})$ is called a *-representation of $(\mathfrak{A},\mathfrak{A}_{0})$, if the following properties are fulfilled:
  • (i)
    $\pi(a^{*})=\pi(a)^{\dagger}:=\pi(a)^{*}\upharpoonright\mathcal{D},\quad\forall\ a\in\mathfrak{A}$;
  • (ii)
    for $a\in\mathfrak{A}$ and $c\in\mathfrak{A}_{0}$, $\pi(a)\,{\scriptstyle\square}\,\pi(c)$ is well-defined and $\pi(a)\,{\scriptstyle\square}\,\pi(c)=\pi(ac)$.
We assume that for every *-representation $\pi$ of $(\mathfrak{A},\mathfrak{A}_{0})$, $\pi({\sf e})={{\mathbb{I}}_{\mathcal{D}}}$, the identity operator on the space $\mathcal{D}$.
The *-representation $\pi$ is said to be
  • closable if there exists $\widetilde{\pi}$ the closure of $\pi$, defined as $\widetilde{\pi}(a)=\overline{\pi(a)}\upharpoonright{\widetilde{\mathcal{D}}}$ where $\widetilde{\mathcal{D}}$ is the completion under the graph topology $t_{\pi}$ defined by the seminorms $\xi\in\mathcal{D}\to\|\xi\|_{\mathcal{K}}+\|\pi(a)\xi\|_{\mathcal{K}}$, $a\in\mathfrak{A}$, with $\|\cdot\|_{\mathcal{K}}$ the quasi norm induced by the quasi inner product on $\mathcal{K}$;
  • closed if $\mathcal{D}[t_{\pi}]$ is complete;
  • cyclic if there exists $\xi\in\mathcal{D}$ such that $\pi(\mathfrak{A}_{0})\xi$ is dense in $\mathcal{K}$ in its quasi norm topology. In this case $\xi$ is called cyclic vector.
Definition 3.11.
Let $(\mathfrak{A},\mathfrak{A}_{0})$ be a quasi *-algebra. We denote by ${\mathcal{I}}_{{\mathfrak{A}}_{0}}^{\,\mathfrak{Y}}(\mathfrak{A})$ the set of all $\mathfrak{Y}$-valued positive sesquilinear maps on $\mathfrak{A}\times\mathfrak{A}$ with the following properties:
  • (i)
    $\Lambda_{\Phi}(\mathfrak{A}_{0})=\mathfrak{A}_{0}/\mathfrak{N}_{\Phi}$ is dense in the completion $\widetilde{\mathfrak{A}}$ of $\mathfrak{A}$ w.r. to the quasi norm $\|\cdot\|_{\Phi}$;
  • (ii)
    $\Phi(ac,d)=\Phi(c,a^{*}d),\quad\forall\ a\in\mathfrak{A},\ c,d\in\mathfrak{A}_{0}$ (left-invariant).
Proposition 3.12.
Let $(\mathfrak{A},\mathfrak{A}_{0})$ be a quasi *-algebra with unit ${\sf e}$ and $\Phi$ be a $\mathfrak{Y}$-valued left-invariant positive sesquilinear map on $\mathfrak{A}\times\mathfrak{A}$. The following statements are equivalent:
  • (i)
    $\Phi\in{\mathcal{I}}_{{\mathfrak{A}}_{0}}^{\,\mathfrak{Y}}(\mathfrak{A})$;
  • (ii)
    there exist a quasi $B_{\mathfrak{Y}}$-space $\mathcal{K}_{\Phi}$ with $\mathfrak{Y}$-valued quasi inner product $\langle{\cdot}|{\cdot}\rangle_{\mathcal{K}_{\Phi}}$, a dense subspace $\mathcal{D}_{\Phi}\subseteq\mathcal{K}_{\Phi}$ and a closed cyclic *-representation $\pi:\mathfrak{A}\to{\mathcal{L}}^{\dagger}(\mathcal{D}_{\Phi},\mathcal{K}_{\Phi})$ with cyclic vector $\xi_{\Phi}$ such that
    $\langle{\pi(a)\xi}|{\eta}\rangle_{\mathcal{K}_{\Phi}}=\langle{\xi}|{\pi(a^{*})\eta}\rangle_{\mathcal{K}_{\Phi}},\quad\forall\xi,\eta\in\mathcal{D}_{\Phi},a\in\mathfrak{A}$
    and such that
    $\Phi(a,b)=\langle{\pi(a)\xi_{\Phi}}|{\pi(b)\xi_{\Phi}}\rangle_{\mathcal{K}_{\Phi}},\quad\forall a,b\in\mathfrak{A}.$
Proof.
The proof proceeds along the lines of that one of [5, Theorem 3.2], again due to Cauchy-Schwarz-like inequality for positive $\mathfrak{Y}$-valued sesquilinear maps (see Proposition 3.1). ∎
Remark 3.13.
By the same arguments as in [5, Corollary 3.5] one can show that every $\mathfrak{Y}$-valued, bounded, left-invariant, positive sesquilinear map on a unital normed quasi *-algebra belongs to ${\mathcal{I}}_{{\mathfrak{A}}_{0}}^{\,\mathfrak{Y}}(\mathfrak{A})$.
In the following example we construct an $L^{2}(\rho)$-valued, bounded, left-invariant, positive sesquilinear maps.
Example 3.14.
Let $\rho$ be a finite trace on a von Neumann algebra ${\mathfrak{M}}$. Let $W\in L^{\infty}(\rho)$, with $W\geq 0$. Let $k\in C([0,\|W\|]\times[0,\|W\|])$ be such that $k\geq 0$. Then, for each $x\in[0,\|W\|]$, the function $\eta_{x}:[0,\|W\|]\to\mathbb{C}$ defined by $\eta_{x}(t)=k(x,t)$ is a continuous, positive function on $[0,\|W\|]$. Therefore, by the functional calculus, $\eta_{x}(W)$ defines a positive operator in $L^{\infty}(\rho)$. Let us define, for every $x\in[0,\|W\|]$ and $X,Y\in L^{2}(\rho)$,
$\varphi(X,Y)(x)=\rho(X\eta_{x}(W)Y^{*}).$
Then, $\varphi(X,Y)\in C([0,\|W\|])$ for all $X,Y\in L^{2}(\rho)$. Moreover, $\varphi:L^{2}(\rho)\times L^{2}(\rho)\to C([0,\|W\|])$ is a bounded, left-invariant, positive sesquilinear map (see [5]).
Let $T\in L^{4}(\rho)$ and consider the map $\widetilde{\varphi}:L^{2}(\rho)\times L^{2}(\rho)\to L^{2}(\rho)$ given by
$\widetilde{\varphi}(X,Y)=T(\varphi(X,Y)(W))T^{*},\quad\forall X,Y\in L^{2}(\rho);$
by functional calculus $\widetilde{\varphi}$ is a bounded, left-invariant, positive sesquilinear map too. Let $f:L^{2}([0,1])\to[0,\|W\|]$. For every $X,Y\in L^{2}(\rho)$, let $\psi:L^{2}(\rho)\to L^{2}([0,1])$ defined by $\psi(X,Y)=\phi(X,Y)\circ f$. Then one can check that $\psi$ is a bounded, left-invariant, positive sesquilinear map.
Every bounded, left-invariant positive sesquilinear map $\Phi$ gives a rise to a bounded positive linear map $\omega$ by letting $\omega(a)=\Phi(a,{\sf e})$, $a\in\mathfrak{A}$. Conversely, for all bounded positive linear maps $\omega$ with the property that
$\|\omega(d^{*}c)\|\leq M\|c\|\|d\|,\forall c,d\in\mathfrak{A}_{0},$
we have that $\omega(a)=\Phi(a,{\sf e})$ for some bounded, left-invariant, positive sesquilinear map $\Phi$ (see the proof of [5, Corollary 3.13]).
4. Representations induced by completely positive sesquilinear maps
In [5] we have considered just positive (but not completely positive) $C^{*}$-valued sesquilinear maps. In this section, we generalize those results by considering more general $\mathfrak{Y}$-valued completely positive sesquilinear maps, with $\mathfrak{Y}$ an ordered Banach bimodule over a *-algebra, satisfying conditions $(D1)-(D2)$ and (4.1) below.
In this section $\mathfrak{A}$ is a normed quasi *-algebra on a *-algebra $\mathfrak{A}_{0}$ and $\mathfrak{Y}$ is an Banach bimodule over a *-algebra $\mathfrak{Y}_{0}$ with a norm and such that the norms of $\mathfrak{Y}$ and $\mathfrak{Y}_{0}$ satisfy
(4.1) $\|z_{2}^{*}yz_{1}\|_{\mathfrak{Y}}\leq\|z_{2}\|_{\mathfrak{Y}_{0}}\|z_{1}\|_{\mathfrak{Y}_{0}}\|y\|_{\mathfrak{Y}},\quad\forall y\in\mathfrak{Y},z_{1},z_{2}\in\mathfrak{Y}_{0}.$
Examples of such $\mathfrak{Y}$ and $\mathfrak{Y}_{0}$ can be found in Example 2.2 (where $L^{\infty}(\rho)$ has to be considered with the operator norm).
Let $\mathfrak{X}$ and $\mathfrak{Y}$ be vector spaces. We will denote by $\mathcal{S}_{\mathfrak{Y}}(\mathfrak{X})$ the space of all $\mathfrak{Y}$-valued sesquilinear maps on $\mathfrak{X}$.
Definition 4.1.
Let $\mathfrak{X}$ be a normed vector space, $\mathfrak{Y}$ an ordered Banach module over the *-algebra $\mathfrak{Y}_{0}$, with positive cone $\mathfrak{K}$ and $(\mathfrak{A},\mathfrak{A}_{0})$ a normed quasi *-algebra with unit. The sesquilinear map $\Phi:\mathfrak{A}\times\mathfrak{A}\to\mathcal{S}_{\mathfrak{Y}}(\mathfrak{X})$ is called
  • bounded if there exists a constant $M>0$ such that
    $\|\Phi(a,b)(x_{1},x_{2})\|_{\mathfrak{Y}}\leq M\|a\|\|b\|\|x_{1}\|\|x_{2}\|,\quad\forall a,b\in\mathfrak{A},x_{1},x_{2}\in\mathfrak{X};$
  • left-invariant if $\Phi(ac,d)=\Phi(c,a^{*}d),\quad\forall a\in\mathfrak{A},c,d\in\mathfrak{A}_{0}$;
  • completely positive if for every $N\in{\mathbb{N}}$, $a_{1},\dots,a_{N}\in\mathfrak{A}$,
    $x_{1},\dots,x_{N}\in{\mathfrak{X}}$,
    $\sum_{i,j=1}^{N}\Phi(a_{i},a_{j})(x_{i},x_{j})\in\mathfrak{K}.$
The Stinespring Theorem is the main result on completely positive linear maps. In the following we consider certain completely positive sesquilinear maps taking values on the space of all $\mathfrak{Y}$-valued sesquilinear maps on a vector space, see Theorem 4.3 below; the proof goes along the lines of that of [4] where more traditional sesquilinear forms were considered, however we give the proof for convenience of the reader.
Let $(\mathfrak{A},\mathfrak{A}_{0})$ be a normed quasi *-algebra with unit ${\sf e}$, $\mathfrak{X}$ be a normed complex vector space, $\mathfrak{Y}$ an ordered Banach bimodule over the *-algebra $\mathfrak{Y}_{0}$, with positive closed cone $\mathfrak{K}$ and satisfying $(D1)-(D2)$ and $\Phi:\mathfrak{A}\times\mathfrak{A}\to\mathcal{S}_{\mathfrak{Y}}({\mathfrak{X}})$ be a left-invariant positive sesquilinear map. Consider the algebraic tensor product $\mathfrak{A}\otimes{\mathfrak{X}}$ and its subset
$\mathcal{N}_{\Phi}=\left\{\sum_{i=1}^{n}a_{i}\otimes x_{i}\in\mathfrak{A}\otimes\mathfrak{X}|\,\sum_{i=1}^{n}\sum_{j=1}^{n}\Phi(a_{i},a_{j})(x_{i},x_{j})=0_{\mathfrak{Y}}\right\}.$
Lemma 4.2.
$\mathcal{N}_{\Phi}$ is a subspace of $\mathfrak{A}\otimes{\mathfrak{X}}$.
Proof.
Consider the map $\langle{\cdot}|{\cdot}\rangle:(\mathfrak{A}\otimes{\mathfrak{X}})\times(\mathfrak{A}\otimes{\mathfrak{X}})\to\mathfrak{Y}$ defined by
$\langle{\sum_{i=1}^{n}a_{i}\otimes x_{i}}|{\sum_{j=1}^{n}b_{j}\otimes x^{\prime}_{j}}\rangle=\sum_{i=1}^{n}\sum_{j=1}^{n}\Phi(a_{i},b_{j})(x_{i},x^{\prime}_{j}),$
$a_{i},b_{j}\in\mathfrak{A},x_{i},x^{\prime}_{j}\in{\mathfrak{X}}$; it is a positive $\mathfrak{Y}$-valued sesquilinear map. Then, the set $\mathcal{N}_{\Phi}$ is a subspace of $\mathfrak{A}\otimes{\mathfrak{X}}$ by Proposition 3.1. ∎
Hence, it is easy to check that the quotient space $(\mathfrak{A}\otimes{\mathfrak{X}})/\mathcal{N}_{\Phi}$ is a quasi normed space.
Theorem 4.3.
Let $\mathfrak{X}$ be a normed complex vector space, $\mathfrak{Y}$ an ordered Banach module over the *-algebra $\mathfrak{Y}_{0}$, with positive cone $\mathfrak{K}$ and satisfying $(D1)-(D2)$. Let $(\mathfrak{A},\mathfrak{A}_{0})$ be a normed quasi *-algebra with unit ${\sf e}$ and $\Phi:\mathfrak{A}\times\mathfrak{A}\to\mathcal{S}_{\mathfrak{Y}}({\mathfrak{X}})$ a left-invariant sesquilinear map. Let $\Phi$ be bounded with bound $M>0$. Then, $\Phi$ is completely positive if and only if there exist a quasi Banach space $\mathcal{K}_{\Phi}$, whose quasi norm is induced by an $\mathfrak{Y}$-valued quasi inner product $\langle{\cdot}|{\cdot}\rangle_{\Phi}$, a dense subspace $\mathcal{D}_{\Phi}$ of $\mathcal{K}_{\Phi}$, a closed *-representation $\pi$ of $\mathfrak{A}$ in ${\mathcal{L}}^{\mbox{\scriptsize${\dagger}$}}(\mathcal{D}_{\Phi},\mathcal{K}_{\Phi})$ and a bounded linear operator $V:\mathfrak{X}\to\mathcal{D}_{\Phi}$ such that $\pi(\mathfrak{A})V{\mathfrak{X}}=(\mathfrak{A}\otimes{\mathfrak{X}})/\mathcal{N}_{\Phi}$ and, for all $a,b\in\mathfrak{A}$ and $x_{1},x_{2}\in\mathfrak{X}$, it holds that
$\Phi(a,b)(x_{1},x_{2})=\langle{\pi(a)Vx_{1}}|{\pi(b)Vx_{2}}\rangle_{\mathcal{K}_{\Phi}}.$
In this case $\|V\|^{2}\leq M\|{\sf e}\|^{2}$. Moreover, the triple $(\pi,V,\mathcal{D}_{\Phi})$ is such that $\overline{\pi(\mathfrak{A})V{\mathfrak{X}}}=\mathcal{K}_{\Phi}$.
Proof.
Suppose that $\Phi$ is completely positive. Consider the quotient $(\mathfrak{A}\otimes\mathfrak{X})/\mathcal{N}_{\Phi}$. To keep the notation lighter, when there is no fear of confusion, we will identify the coset $(\sum_{i=1}^{n}a_{i}\otimes x_{i})+\mathcal{N}_{\Phi}$ for each $n\in\mathbb{N}$ and $a_{1},\dots,a_{n}\in\mathfrak{A}$, $x_{1},\dots,x_{n}\in\mathfrak{X}$ with its representative $\sum_{i=1}^{n}a_{i}\otimes x_{i}$. Define the sesquilinear map on $(\mathfrak{A}\otimes\mathfrak{X})/\mathcal{N}_{\Phi}$ by
$\sum_{i=1}^{n}a_{i}\otimes x_{i}\Bigg{|}\sum_{j=1}^{m}b_{j}\otimes x^{\prime}_{j}\right>=\sum_{i=1}^{n}\sum_{j=1}^{m}\Phi(a_{i},b_{j})(x_{i},x^{\prime}_{j}),$
$\sum_{i=1}^{n}a_{i}\otimes x_{i},\sum_{j=1}^{m}b_{j}\otimes x^{\prime}_{j}\in(\mathfrak{A}\otimes\mathfrak{X})/\mathcal{N}_{\Phi}$; it is positive since for each $N\in\mathbb{N}$ and every $a_{1},\dots,a_{N}\in\mathfrak{A}$, $x_{1},\dots,x_{N}\in\mathfrak{X}$ it is
$\sum_{i=1}^{N}a_{i}\otimes x_{i}\Bigg{|}\sum_{j=1}^{N}a_{j}\otimes x_{j}\right>=\sum_{i,j=1}^{N}\Phi(a_{i},a_{j})(x_{i},x_{j})\in\mathfrak{K}.$
[…]
$\pi_{0}(a)\left(\sum_{i=1}^{n}a_{i}\otimes x_{i}\right):=\sum_{i=1}^{n}aa_{i}\otimes x_{i},\quad a_{i}\in\mathfrak{A},x_{i}\in{\mathfrak{X}},n\in\mathbb{N}.$
Then it is well-defined. By the left-invariance of $\Phi$, $\pi_{0}(a)^{*}=\pi_{0}(a^{*})$. Hence, if $\sum_{i}a_{i}\otimes x_{i}\in\mathcal{N}_{\Phi}$, then
$\pi_{0}(a)\left(\sum_{i=1}^{n}a_{i}\otimes x_{i}\right)\Bigg{|}\sum_{j=1}^{m}c_{j}\otimes x^{\prime}_{j}\right>\right\|_{\mathfrak{Y}}\\ \leq 4\left\|\sum_{i=1}^{n}a_{i}\otimes x_{i}\right\|_{\mathcal{K}_{\Phi}}\left\|\pi_{0}(a^{*})\left(\sum_{j=1}^{m}c_{j}\otimes x^{\prime}_{j}\right)\right\|_{\mathcal{K}_{\Phi}}=0,$
for every $\sum_{j=1}^{m}c_{j}\otimes x^{\prime}_{j}\in(\mathfrak{A}_{0}\otimes{\mathfrak{X}})/\mathcal{N}_{\Phi}$; by the density of $(\mathfrak{A}_{0}\otimes{\mathfrak{X}})/\mathcal{N}_{\Phi}$ in $\mathcal{K}_{\Phi}$ and by the joint continuity of the quasi inner product, it follows that $\left\|\pi_{0}(a)\left(\sum_{i=1}^{n}a_{i}\otimes x_{i}\right)\right\|_{\mathcal{K}_{\Phi}}=0$, thus $\pi_{0}(a)\left(\sum_{i=1}^{n}a_{i}\otimes x_{i}\right)\in\mathcal{N}_{\Phi}$. Therefore, $\pi_{0}(a)$ maps $\mathcal{N}_{\Phi}$ into itself, so $\pi_{0}(a)$ is a well-defined linear map on $(\mathfrak{A}_{0}\otimes{\mathfrak{X}})/\mathcal{N}_{\Phi}$.
The map $\pi_{0}$ is a *-representation of $\mathfrak{A}$ in ${\mathcal{L}}^{\mbox{\scriptsize${\dagger}$}}((\mathfrak{A}_{0}\otimes{\mathfrak{X}})/\mathcal{N}_{\Phi},\mathcal{K}_{\Phi})$. By the density of $(\mathfrak{A}_{0}\otimes{\mathfrak{X}})/\mathcal{N}_{\Phi}$ and Remark 3.9, $\pi_{0}(a)$ is closable. Let $\pi$ be the closure of $\pi_{0}$ and let $\mathcal{D}_{\Phi}$ be its domain. Let now $V:\mathfrak{X}\to(\mathfrak{A}_{0}\otimes\mathfrak{X})/\mathcal{N}_{\Phi}\subset\mathcal{D}_{\Phi}$ be the operator given by
$Vx:={\sf e}\otimes x+\mathcal{N}_{\Phi},\quad x\in\mathfrak{X}.$
Then $\pi(\mathfrak{A})V{\mathfrak{X}}=(\mathfrak{A}\otimes{\mathfrak{X}})/\mathcal{N}_{\Phi}$; indeed, from $\pi(a)Vx=a\otimes x+\mathcal{N}_{\Phi}$, $a\in\mathfrak{A}$, $x\in{\mathfrak{X}}$. Then,
$\|Vx\|_{\mathcal{K}_{\Phi}}^{2}=\|\Phi({\sf e},{\sf e})(x,x)\|_{\mathfrak{Y}}\leq M\|{\sf e}\|^{2}\|x\|^{2},\quad\forall x\in\mathfrak{X};$
thus $V$ is bounded with $\|V\|^{2}\leq M\|{\sf e}\|^{2}$.
Finally, for all $a,b\in\mathfrak{A}$ and $x_{1},x_{2}\in\mathfrak{X}$, it holds that
$\Phi(a,b)(x_{1},x_{2})=\langle{\pi(a)Vx_{1}}|{\pi(b)Vx_{2}}\rangle.$
Conversely, suppose that for all $a,b\in\mathfrak{A}$ and $x_{1},x_{2}\in{\mathfrak{X}}$, we have that $\Phi(a,b)(x_{1},x_{2})=\langle{\pi(a)Vx_{1}}|{\pi(b)Vx_{2}}\rangle$. Then, for every $N\in\mathbb{N}$ and $a_{1},\dots,a_{N}\in\mathfrak{A}$, $x_{1},\dots,x_{N}\in{\mathfrak{X}}$,
$\begin{split}\sum_{i,j=1}^{N}\Phi(a_{i},a_{j})(x_{i},x_{j})&=\sum_{i,j=1}^{N}\langle{\pi(a_{i})Vx_{i}}|{\pi(a_{j})Vx_{j}}\rangle\\ &=\left<\sum_{i=1}^{N}\pi(a_{i})Vx_{i}\Bigg{|}\sum_{j=1}^{N}\pi(a_{j})Vx_{j}\right>\in\mathfrak{K}.\end{split}$
Finally, the triple $(\pi,V,\mathcal{D}_{\Phi})$ is such that $\overline{\pi(\mathfrak{A})V{\mathfrak{X}}}=\mathcal{K}_{\Phi}$. ∎
In the following we will refer to a triple $(\pi,V,\mathcal{D}_{\Phi})$ as in Theorem 4.3 as a Stinespring triple decomposing $\Phi$.
Definition 4.4.
Let ${\mathfrak{X}}_{1},{\mathfrak{X}}_{2}$ be two quasi $B_{\mathfrak{Y}}$-spaces with respect to the quasi norms $\|\cdot\|_{\Phi}$ and $\|\cdot\|_{\Psi}$, respectively, induced by two $\mathfrak{Y}$-valued quasi inner products $\Phi:{\mathfrak{X}}_{1}\times{\mathfrak{X}}_{1}\to\mathfrak{Y}$ and $\Psi:{\mathfrak{X}}_{2}\times{\mathfrak{X}}_{2}\to\mathfrak{Y}$. A surjective operator $U:{\mathfrak{X}}_{1}\to{\mathfrak{X}}_{2}$ is said unitary if
$\langle{U\xi}|{U\eta}\rangle_{\Psi}=\langle{\xi}|{\eta}\rangle_{\Phi},\quad\forall\xi,\eta\in{\mathfrak{X}}_{1}.$
In analogy to Corollary 3.10 in [4], we can prove the following
Proposition 4.5.
Let $\mathfrak{X}$ be a normed complex vector space, $\mathfrak{Y}$ an ordered Banach module over the *-algebra $\mathfrak{Y}_{0}$, with positive cone $\mathfrak{K}$ and satisfying $(D1)-(D2)$. Let $(\mathfrak{A},\mathfrak{A}_{0})$ be a normed quasi *-algebra with unit ${\sf e}$ and $\Phi:\mathfrak{A}\times\mathfrak{A}\to\mathcal{S}_{\mathfrak{Y}}({\mathfrak{X}})$ be a bounded left-invariant sesquilinear map. If $\Phi$ is completely positive then $\pi$ and $V$ in the triple $(\pi,V,\mathcal{D}_{\Phi})$ are uniquely determined by $\Phi$ up to unitary equivalence, i.e., if $(\pi,V,\mathcal{D}_{\Phi})$ and $(\pi_{1},V_{1},\mathcal{E}_{\Phi})$ are two Stinespring triples decomposing $\Phi$, there exists a unitary operator $U$ such that $UV=V_{1}$, $U\mathcal{D}_{\Phi}=\mathcal{E}_{\Phi}$ and $\pi(a)=U^{-1}\pi_{1}(a)U$, for all $a\in\mathfrak{A}$.
Remark 4.6.
By Proposition 4.5 all the Stinespring triples $(\pi,V,\mathcal{D}_{\Phi})$ decomposing $\Phi$ are such that $\overline{\pi(\mathfrak{A})V{\mathfrak{X}}}=\mathcal{K}_{\Phi}$.
In the next results we will need the additional assumption that the norm on the ordered Banach bimodule $\mathfrak{Y}$ is order-preserving. Examples of Banach bimodules with an order-preserving norm are $M([0,1])$ and $\ell_{2}(C(\Omega))$, with $\Omega$ a compact Hausdorff space.
Example 4.7.
Let $\rho$ be a finite trace on a von Neumann algebra ${\mathfrak{M}}$. The norm in $L^{2}(\rho)$ preserves the order. Indeed, if $A,B\in L^{\infty}(\rho)$ and $A,B\geq 0$, then
$\rho(AB)=\rho((AB)^{1/2}B^{1/2})=\rho(B^{1/2}AB^{1/2})\geq 0.$
Now, take some general $A,B\in L^{2}(\rho)$ with $A,B\geq 0$. Then there exist sequences $\{A_{n}\}_{n},\{B_{n}\}_{n}\subset L^{\infty}(\rho)$ such that $A_{n}\to A$, $B_{n}\to B$ as $n\to\infty$ and, moreover, $A_{n},B_{n}\geq 0$ and therefore $\rho(A_{n}B_{n})\geq 0$, for all $n\in\mathbb{N}$; from this it follows that $\rho(A_{n}B_{n})\to\rho(AB)$ as $n\to\infty$, hence $\rho(AB)\geq 0$ since $\rho(A_{n}B_{n})\geq 0$ for all $n\in\mathbb{N}$.
Finally, let $A,B\in L^{2}(\rho)$ with $A,B,A-B\geq 0$; then
$\rho(A^{2})-\rho(B^{2})=\rho(A(A-B))+\rho((A-B)B)\geq 0,$
so $\|A\|_{2}^{2}\geq\|B\|_{2}^{2}$, thus $\|A\|_{2}\geq\|B\|_{2}$.
We notice that the norm in $L^{2}(\rho)$ is order-preserving even if $\rho$ is not finite. Indeed, if $A,B\in{\mathfrak{M}}\cap L^{2}(\rho)$ with $A,B>0$, then by Lemma 4.8 in [6], it can be shown that $\rho(AB)\geq 0$ and the previous argument then applies.
Following [17, Section 3.5] we now provide a generalization of the Radon-Nikodym theorem for completely positive sesquilinear maps with values in the space $\mathcal{S}_{\mathfrak{Y}}({\mathfrak{X}})$ of $\mathfrak{Y}$-valued sesquilinear maps on ${\mathfrak{X}}\times{\mathfrak{X}}$.
Proposition 4.8.
Let $\mathfrak{X}$ be a normed complex vector space, $\mathfrak{Y}$ an ordered Banach module over the *-algebra $\mathfrak{Y}_{0}$, with positive cone $\mathfrak{K}$ and satisfying $(D1)-(D2)$. Let $(\mathfrak{A},\mathfrak{A}_{0})$ be a normed quasi *-algebra with unit ${\sf e}$ and $\Psi,\Phi:\mathfrak{A}\times\mathfrak{A}\to\mathcal{S}_{\mathfrak{Y}}({\mathfrak{X}})$ be bounded, left-invariant, completely positive sesquilinear maps. Suppose that there exists $\gamma>0$ such that, for every $N\in\mathbb{N}$, for all $a_{1},\dots,a_{N}\in\mathfrak{A},x_{1},\dots,x_{N}\in{\mathfrak{X}}$
$\sum_{i,j=1}^{N}\Psi(a_{i},a_{j})(x_{i},x_{j})\leq\gamma\sum_{i,j=1}^{N}\Phi(a_{i},a_{j})(x_{i},x_{j}).$
Let $(\pi_{\Psi},V_{\Psi},\mathcal{D}_{\Psi})$ and $(\pi_{\Phi},V_{\Phi},\mathcal{D}_{\Phi})$ be two Stinespring triples decomposing $\Psi$ and $\Phi$, respectively. Then, there exists an operator $T:\mathcal{K}_{\Phi}\to\mathcal{K}_{\Psi}$ such that:
  • (i)
    $TV_{\Phi}=V_{\Psi}$;
  • (ii)
    $T\pi_{\Phi}(a)=\pi_{\Psi}(a)\,{\scriptstyle\square}\,T$ on $\pi_{\Phi}(\mathfrak{A}_{0})V_{\Phi}{\mathfrak{X}}$, for every $a\in\mathfrak{A}$,
and for all $a,b\in\mathfrak{A}$, $x_{1},x_{2}\in{\mathfrak{X}}$,
$\Psi(a,b)(x_{1},x_{2})=\langle{T\pi_{\Phi}(a)V_{\Phi}x_{1}}|{T\pi_{\Phi}(b)V_{\Phi}x_{2}}\rangle_{\mathcal{K}_{\Psi}}.$
Moreover, if the norm $\|\cdot\|_{\mathfrak{Y}}$ in $\mathfrak{Y}$ is order-preserving, then $T$ is bounded and $\|T\|\leq\sqrt{\gamma}$.
Proof.
Let $N\in\mathbb{N}$ and $a_{1},\dots,a_{N}\in\mathfrak{A}$ and $x_{1},\dots,x_{N}\in{\mathfrak{X}}$, then
(4.2) $\sum_{i=1}^{N}\pi_{\Psi}(a_{i})V_{\Psi}x_{i}\Bigg{|}\sum_{j=1}^{N}\pi_{\Psi}(a_{j})V_{\Psi}x_{j}\right>_{\mathcal{K}_{\Psi}}=\sum_{i,j=1}^{N}\Psi(a_{i},a_{j})(x_{i},x_{j})$
$\hskip 28.45274pt\leq\gamma\sum_{i,j=1}^{N}\Phi(a_{i},a_{j})(x_{i},x_{j})$
$\hskip 28.45274pt=\gamma\left<\sum_{i=1}^{N}\pi_{\Phi}(a_{i})V_{\Phi}x_{i}\Bigg{|}\sum_{j=1}^{N}\pi_{\Phi}(a_{j})V_{\Phi}x_{j}\right>_{\mathcal{K}_{\Phi}}.$
If $\sum_{i=1}^{N}\pi_{\Phi}(a_{i})V_{\Phi}x_{i}=0$, then also $\sum_{i=1}^{N}\pi_{\Psi}(a_{i})V_{\Psi}x_{i}=0$, hence, we can define a bounded linear operator $T$ by
(4.3) $T\sum_{i=1}^{N}\pi_{\Phi}(a_{i})V_{\Phi}x_{i}=\sum_{i=1}^{N}\pi_{\Psi}(a_{i})V_{\Psi}x_{i},$
for all $a_{i}\in\mathfrak{A},\,x_{i}\in{\mathfrak{X}},i\in\{1,\dots,N\}$ and extend it to the closure $\mathcal{K}_{\Phi}=\overline{\pi_{\Phi}(\mathfrak{A})V_{\Phi}{\mathfrak{X}}}$ into $\mathcal{K}_{\Psi}=\overline{\pi_{\Psi}(\mathfrak{A})V_{\Psi}{\mathfrak{X}}}$. Moreover, by taking $a={\sf e}\in\mathfrak{A}_{0}$ then $TV_{\Phi}=V_{\Psi}$, since $\pi_{\Phi}({\sf e})={\mathbb{I}}_{\mathcal{D}_{\Phi}}$ and $\pi_{\Psi}({\sf e})={\mathbb{I}}_{\mathcal{D}_{\Psi}}$.
Then $T\,{\scriptstyle\square}\,\pi_{\Phi}(a)=\pi_{\Psi}(a)\,{\scriptstyle\square}\,T$, for every $a\in\mathfrak{A}$. Indeed, if now $a\in\mathfrak{A}$, $c\in\mathfrak{A}_{0}$, $x\in{\mathfrak{X}}$, by (4.3)
$\begin{split}T(\pi_{\Phi}(a)\,{\scriptstyle\square}\,\pi_{\Phi}(c))V_{\Phi}x&=T({\pi_{\Phi}(a)^{\dagger}}^{*}\pi_{\Phi}(c))V_{\Phi}x=(T{\pi_{\Phi}(a)^{\dagger}}^{*})\pi_{\Phi}(c)V_{\Phi}x,\end{split}$
on the other hand,
$\begin{split}T(\pi_{\Phi}(a)\,{\scriptstyle\square}\,\pi_{\Phi}(c))V_{\Phi}x&=T\pi_{\Phi}(ac)V_{\Phi}x=\pi_{\Psi}(ac)V_{\Psi}x\\ &={\pi_{\Psi}(a)^{\dagger}}^{*}\pi_{\Psi}(c)V_{\Psi}x={\pi_{\Psi}(a)^{\dagger}}^{*}(T\pi_{\Phi}(c)V_{\Phi}x)\\ &=({\pi_{\Psi}(a)^{\dagger}}^{*}T)\pi_{\Phi}(c)V_{\Phi}x=(\pi_{\Psi}(a)\,{\scriptstyle\square}\,T)\pi_{\Phi}(c)V_{\Phi}x,\end{split}$
and since the set $\pi_{\Phi}(\mathfrak{A}_{0})V_{\Phi}{\mathfrak{X}}=(\mathfrak{A}_{0}\otimes{\mathfrak{X}})/\mathcal{N}_{\Phi}$ is dense in $\mathcal{K}_{\Phi}$ (see the proof of Theorem 4.3), we conclude that $T\pi_{\Phi}(a)=\pi_{\Psi}(a)\,{\scriptstyle\square}\,T$ on $\pi_{\Phi}(\mathfrak{A}_{0})V_{\Phi}{\mathfrak{X}}$, for every $a\in\mathfrak{A}$.
Finally, if $y_{1}\leq y_{2}$, with $y_{1},y_{2}\in\mathfrak{Y}$ implies that also $\|y_{1}\|_{\mathfrak{Y}}\leq\|y_{2}\|_{\mathfrak{Y}}$ then, by (4.2), we have that $\|T\|\leq\sqrt{\gamma}$. The last statement of the proposition follows by the very construction of $T$.∎
A Radon Nikodym-like theorem holds for positive sesquilinear maps $\Phi,\Psi$ and not just for completely positive sesquilinear maps.
Proposition 4.9.
Let $\mathfrak{X}$ be a normed vector space, $\mathfrak{Y}$ an ordered Banach module over the *-algebra $\mathfrak{Y}_{0}$ with norm, with positive cone $\mathfrak{K}$ and satisfying $(D1)-(D2)$. Let the norm $\|\cdot\|_{\mathfrak{Y}}$ preserve the order in $\mathfrak{Y}$. Let $(\mathfrak{A},\mathfrak{A}_{0})$ be a normed quasi *-algebra with unit ${\sf e}$ and $\Phi,\Psi:\mathfrak{A}\times\mathfrak{A}\to\mathcal{S}_{\mathfrak{Y}}({\mathfrak{X}})$ be bounded left-invariant positive sesquilinear maps such that
$\Psi(a,a)(x,x)\leq\gamma\,\Phi(a,a)(x,x),\quad\forall a\in\mathfrak{A},x\in{\mathfrak{X}}$
for some $\gamma>0$. Let $\mathcal{K}_{\Phi}$ and $\mathcal{K}_{\Psi}$ be quasi $B_{\mathfrak{Y}}$-spaces, $\mathcal{D}_{\Phi}$ and $\mathcal{D}_{\Psi}$ dense subspaces of $\mathcal{K}_{\Phi}$ and $\mathcal{K}_{\Psi}$, respectively, $\pi_{\Phi}:\mathfrak{A}\to{\mathcal{L}}^{\dagger}(\mathcal{D}_{\Phi},\mathcal{K}_{\Phi})$ and $\pi_{\Psi}:\mathfrak{A}\to{\mathcal{L}}^{\dagger}(\mathcal{D}_{\Psi},\mathcal{K}_{\Psi})$ closed cyclic *-representations of $\Phi$ and $\Psi$, respectively, with cyclic vectors $\xi_{\Phi}$ and $\xi_{\Psi}$, respectively, as in Proposition 3.12. Then there exists a unique operator $T:\mathcal{K}_{\Phi}\to\mathcal{K}_{\Psi}$, with $\|T\|\leq\sqrt{\gamma}$ such that for all $a,b\in\mathfrak{A}$,
$\Psi(a,b)=\langle{T\pi_{\Phi}(a)\xi_{\Phi}}|{T\pi_{\Phi}(b)\xi_{\Phi}}\rangle_{\mathcal{K}_{\Psi}}.$
Moreover, $T\pi_{\Phi}(a)=\pi_{\Psi}(a)\,{\scriptstyle\square}\,T$ on $\pi_{\Phi}({\mathfrak{A}}_{0})\xi_{\Phi}$, for all $a\in\mathfrak{A}$.
Proof.
The proof goes along the lines of that of Proposition 4.8, by applying the representations of $\Phi$ and $\Psi$ of Proposition 3.12 and by letting $T\pi_{\Phi}(a)\xi_{\Phi}=\pi_{\Psi}(a)\xi_{\Psi}$ for all $a\in\mathfrak{A}$, see also Remark 3.13. ∎
4.1. Examples
We discuss here some concrete realizations.
Example 4.10.
Let $\mathfrak{Y}$ be either a $C^{*}$-algebra or an ordered Banach module over a *-algebra, satisfying $(D1)-(D2)$ and let $\widetilde{\Phi}:\mathfrak{A}\times\mathfrak{A}\to\mathfrak{Y}$ be a bounded left-invariant positive sesquilinear map. Define the map $\Phi:\mathfrak{A}\times\mathfrak{A}\to\mathcal{S}_{\mathfrak{Y}}(\mathfrak{A}_{0})$ by
$\Phi(a,b)(c,d)=\widetilde{\Phi}(ac,bd),\quad\forall a,b\in\mathfrak{A},c,d\in\mathfrak{A}_{0}.$
Then $\Phi$ is bounded, left-invariant, completely positive and sesquilinear.
Example 4.11.
Let $\mathfrak{Y}=L^{2}(\rho)$ where $\rho$ is a finite trace on a von Neumann algebra ${\mathfrak{M}}=L^{\infty}(\rho)$. Let $W_{1}\in L^{\infty}(\rho).$ Define the map
$\Phi_{1}:L^{4}(\rho)\times L^{4}(\rho)\to\mathcal{S}_{L^{2}(\rho)}(L^{\infty}(\rho))$
by
$\Phi_{1}(A,B)(X,Y)=Y^{*}W_{1}^{*}B^{*}AW_{1}X,\quad A,B\in L^{4}(\rho),X,Y\in L^{\infty}(\rho).$
Then $\Phi_{1}$ is a bounded, left-invariant, completely positive sesquilinear map.
Similarly, let $W_{2}\in L^{4}(\rho)$ and define the map
$\Phi_{2}:L^{\infty}(\rho)\times L^{\infty}(\rho)\to\mathcal{S}_{L^{2}(\rho)}(L^{\infty}(\rho))$
by
$\Phi_{2}(A,B)(X,Y)=Y^{*}W_{2}^{*}B^{*}AW_{2}X,\quad A,B,X,Y\in L^{\infty}(\rho).$
Then $\Phi_{2}$ is a bounded, left-invariant, completely positive sesquilinear map.
Example 4.12.
Let $\rho$ be a finite trace on a von Neumann algebra ${\mathfrak{M}}$ and let ${\mathfrak{M}}^{*}$ be the dual of ${\mathfrak{M}}$. Let $W\in{\mathfrak{M}}$ and
$\Phi:L^{2}(\rho)\times L^{2}(\rho)\to\mathcal{S}_{{\mathfrak{M}}^{*}}(L^{\infty}({\mathfrak{M}}))$
be given by
$(\Phi(A,B)(X,Y))(T)=\rho(AXWTW^{*}Y^{*}B^{*}),$
$T\in{\mathfrak{M}}$, $X,Y\in L^{\infty}(\rho)$, $A,B\in L^{2}(\rho)$. Then $\Phi$ is a bounded, left-invariant, completely positive sesquilinear map.
4.2. Applications to Banach bimodule-valued maps
Before introducing the applications we need the following definition.
Definition 4.13.
Let $\mathfrak{Y}$, $\mathfrak{Z}$ two Banach modules over the *-algebras with norm $\mathfrak{Y}_{0}$ and $\mathfrak{Z}_{0}$, respectively. Let $\mathfrak{Z}$ be an ordered space and be $\mathfrak{J}$ the positive cone of $\mathfrak{Z}_{0}$. Let $\omega:\mathfrak{Y}\to\mathfrak{Z}$ a linear map and $(\mathfrak{A},\mathfrak{A}_{0})$ be a quasi *-algebra. A sesquilinear map $\Phi:\mathfrak{A}\times\mathfrak{A}\to\mathfrak{Y}$ will be called completely positive with respect to $\omega$ if, for every $N\in\mathbb{N}$ and each $a_{j}\in\mathfrak{A}$, $y_{i}\in\mathfrak{Y}_{0}$, $i,j\in\{1,...,N\}$, it is
$\omega\left(\sum_{i,j=1}^{N}y_{j}^{*}\Phi(a_{i},a_{j})y_{i}\right)\in\mathfrak{J}.$
Proposition 4.14.
Let $(\mathfrak{A},\mathfrak{A}_{0})$ be a quasi *-algebra with unit ${\sf e}$, $\mathfrak{Y}$, $\mathfrak{Z}$ be two Banach modules over the *-algebras with norm $\mathfrak{Y}_{0}$ and $\mathfrak{Z}_{0}$, respectively. Let $\mathfrak{Z}$ satisfy $(D1)-(D2)$. Let $\omega:\mathfrak{Y}\to\mathfrak{Z}$ be a bounded linear map and $\Phi:\mathfrak{A}\times\mathfrak{A}\to\mathfrak{Y}$ be a bounded completely positive sesquilinear map with respect to $\omega$. Then, there exist a quasi $B_{\mathfrak{Z}}$-space $\mathcal{K}_{\Phi}$, a closed *-representation $\pi$ of $\mathfrak{A}$ defined on a dense domain $\mathcal{D}_{\Phi}\subseteq\mathcal{K}_{\Phi}$, and a bounded linear operator $V:\mathfrak{Y}_{0}\to\mathcal{D}_{\Phi}$ such that for all $a,b\in\mathfrak{A}$ and $y_{1},y_{2}\in\mathfrak{Y}_{0}$ we have
$\omega(y_{2}^{*}\Phi(a,b)y_{1})=\langle{\pi(a)Vy_{1}}|{\pi(b)Vy_{2}}\rangle_{\mathcal{K}_{\Phi}},$
and if $\|\omega\|\leq 1$, then $\|V\|^{2}\leq\|\Phi({\sf e},{\sf e})\|_{\mathfrak{Y}}$. Moreover, for all $a,b\in\mathfrak{A}$ and $y_{1},y_{2}\in\mathfrak{Y}_{0}$, we have
$\|\omega(y_{2}^{*}\Phi(a,b)y_{1})\|_{\mathfrak{Z}}\leq 4\|\omega(y_{1}^{*}\Phi(a,a)y_{1})\|_{\mathfrak{Z}}^{1/2}\|\omega(y_{2}^{*}\Phi(b,b)y_{2})\|_{\mathfrak{Z}}^{1/2}.$
Proof.
The map $\Phi$ induces $\widetilde{\Phi}:\mathfrak{A}\times\mathfrak{A}\to\mathcal{S}_{\mathfrak{Z}}(\mathfrak{Y}_{0})$, given by:
$\widetilde{\Phi}(a,b)(y_{1},y_{2})=\omega(y_{2}^{*}\Phi(a,b)y_{1}),\quad a,b\in\mathfrak{A},y_{1},y_{2}\in\mathfrak{Y}_{0}$
which is clearly completely positive in the sense of Definition 4.1. If $\Phi$ is positive and left-invariant, so is $\widetilde{\Phi}$.
By the boundedness of $\Phi$ and (4.1), it is straightforward to check that $\widetilde{\Phi}$ is bounded, hence, the first statement follows from Theorem 4.3. By the very construction of $V$ it is
$\begin{split}\|Vz\|^{2}&=\widetilde{\Phi}({\sf e},{\sf e})(z,z)=\omega(z^{*}\Phi({\sf e},{\sf e})z)\\ &\leq\|\Phi({\sf e},{\sf e})\|_{\mathfrak{Y}}\|z\|_{\mathfrak{Y}_{0}}^{2}\|\omega\|,\qquad\qquad\forall z\in\mathfrak{Y}_{0}.\end{split}$
If moreover $\|\omega\|\leq 1$, it is $\|V\|^{2}\leq\|\Phi({\sf e},{\sf e})\|_{\mathfrak{Y}}$. The last inequality of the proposition follows from Proposition 3.1 applied to the obtained representation. ∎
Example 4.15.
Consider $L^{2}([0,1])$ and $M([0,1])$. Let $h$ be a positive continuous function on $[0,1]$, $\alpha$ be a homeomorphism of $[0,1]$, $\mu$ be the Lebesgue measure on $[0,1]$ and define $\Phi:L^{2}([0,1])\times L^{2}([0,1])\to M([0,1])$ by
[…]
i.e., $d\Phi(f,g)=h((\overline{g}f)\circ\alpha)d\mu$, for all $f,g\in C([0,1])$. Then $\Phi$ is a bounded, left-invariant, completely positive sesquilinear map with respect to the identity map on $M([0,1])$.
Moreover, if $\mathfrak{Z}$ is another ordered Banach bimodule over a *-algebra $\mathfrak{Z}_{0}$ with norm, such that $\mathfrak{Z}$ satisfies $(D1)-(D2)$ and $\omega:M([0,1])\to\mathfrak{Z}$ is a bounded, positive linear map, then $\Phi$ will be completely positive w.r. to $\omega$. As a concrete example, let $\rho$ be a finite trace on a von Neumann algebra ${\mathfrak{M}}$. Let $W\in{\mathfrak{M}}$ , with $W\geq 0$ and $\|W\|=1$, and consider the function $\eta_{x}$ defined as in Example 3.14. For every $\mu\in M([0,1])$ and $T\in{\mathfrak{M}}$, let
[…]
Then, $\omega(\mu)\in{\mathfrak{M}}^{*}$ for every $\mu\in M([0,1])$. It is straightforward to check that $\omega:M([0,1])\to{\mathfrak{M}}^{*}$ is a bounded positive linear map.
4.3. Applications to operator-valued maps
In the following we will denote by ${\mathfrak{B}}({\mathfrak{X}},\mathfrak{Y})$ the space of bounded linear operators from ${\mathfrak{X}}$ into $\mathfrak{Y}$.
Definition 4.16.
Let ${\mathfrak{X}}$ be a Banach space, $\Gamma:{\mathfrak{X}}\times{\mathfrak{X}}\to{\mathfrak{X}}$ a bounded sesquilinear map and $\mathfrak{Y}$ an ordered Banach bimodule over the *-algebra $\mathfrak{Y}_{0}$ with norm, with positive cone $\mathfrak{K}$ and satisfying $(D1)-(D2)$. Let $(\mathfrak{A},\mathfrak{A}_{0})$ be a quasi *-algebra with unit ${\sf e}$. A sesquilinear map $\Phi:\mathfrak{A}\times\mathfrak{A}\to{\mathfrak{B}}({\mathfrak{X}},\mathfrak{Y})$ will be called completely positive w.r. to $\Gamma$ if for each $N\in\mathbb{N}$ and for every $a_{1},\dots,a_{N}\in\mathfrak{A}$, $x_{1},\dots,x_{N}\in{\mathfrak{X}}$ it is
$\sum_{i,j=1}^{N}\Phi(a_{i},a_{j})(\Gamma(x_{i},x_{j}))\in\mathfrak{K}.$
Proposition 4.17.
Let ${\mathfrak{X}}$ be a Banach space, $\Gamma:{\mathfrak{X}}\times{\mathfrak{X}}\to{\mathfrak{X}}$ a bounded sesquilinear map and $\mathfrak{Y}$ an ordered Banach bimodule over the *-algebra $\mathfrak{Y}_{0}$ with norm satisfying $(D1)-(D2)$. Let $(\mathfrak{A},\mathfrak{A}_{0})$ be a quasi *-algebra with unit ${\sf e}$ and $\Phi:\mathfrak{A}\times\mathfrak{A}\to{\mathfrak{B}}({\mathfrak{X}},\mathfrak{Y})$ be a bounded, left-invariant, completely positive sesquilinear map. Then, there exist a dense subspace $\mathcal{D}_{\Phi}$ of a quasi $B_{\mathfrak{Y}}$-space $\mathcal{K}_{\Phi}$, a closed *-representation $\pi$ of $\mathfrak{A}$ in ${\mathcal{L}}^{\mbox{\scriptsize${\dagger}$}}(\mathcal{D}_{\Phi},\mathcal{K}_{\Phi})$ and a bounded linear operator $V:{\mathfrak{X}}\to\mathcal{D}_{\Phi}$ with $\|V\|^{2}\leq\|\Phi({\sf e},{\sf e})\|\|\Gamma\|$ such that for all $a,b\in\mathfrak{A}$ we have
$\Phi(a,b)(\Gamma(x_{1},x_{2}))=\langle{\pi(a)Vx_{1}}|{\pi(b)Vx_{2}}\rangle_{\mathcal{K}_{\Phi}},\quad\forall a,b\in\mathfrak{A},\,\,x_{1},x_{2}\in{\mathfrak{X}}.$
Moreover, for all $a,b\in\mathfrak{A}$ and $x_{1},x_{2}\in{\mathfrak{X}}$, we have
$\|\Phi(a,b)(\Gamma(x_{1},x_{2}))\|_{\mathfrak{Y}}\leq 4\|\Phi(a,a)(\Gamma(x_{1},x_{1}))\|_{\mathfrak{Y}}^{1/2}\|\Phi(b,b)(\Gamma(x_{2},x_{2}))\|_{\mathfrak{Y}}^{1/2}.$
Proof.
Let $\mathcal{S}_{\mathfrak{Y}}({\mathfrak{X}})$ be the space of all $\mathfrak{Y}$-valued sesquilinear maps on ${\mathfrak{X}}$ and $\widetilde{\Phi}:\mathfrak{A}\times\mathfrak{A}\to\mathcal{S}_{\mathfrak{Y}}({\mathfrak{X}})$ be given by:
$\widetilde{\Phi}(a,b)(x_{1},x_{2}):=\Phi(a,b)(\Gamma(x_{1},x_{2})),\quad\forall a,b\in\mathfrak{A},\,\,x_{1},x_{2}\in{\mathfrak{X}}.$
Then, $\widetilde{\Phi}$ is a left-invariant, completely positive sesquilinear map on $\mathfrak{Y}$ since it inherits these properties from $\Phi$. Moreover $\widetilde{\Phi}$ is bounded. By Theorem 4.3, the statement follows. By the definition of the operator $V$ in the proof of Theorem 4.3, for every $x\in{\mathfrak{X}}$ we have that
$\begin{split}\|\langle{Vx}|{Vx}\rangle_{\mathcal{K}_{\Phi}}\|_{\mathfrak{Y}}&=\|\langle{{\sf e}\otimes x}|{{\sf e}\otimes x}\rangle_{\mathcal{K}_{\Phi}}\|_{\mathfrak{Y}}=\|(\Phi({\sf e},{\sf e}))(\Gamma(x,x))\|_{\mathfrak{Y}}\\ &\leq\|\Phi({\sf e},{\sf e})\|\|\Gamma(x,x)\|\leq\|\Phi({\sf e},{\sf e})\|\|\Gamma\|\|x\|^{2}.\end{split}$
The last inequality of the proposition follows from Proposition 3.1 applied to the obtained representation. ∎
The following example shows a map $\Phi$ satisfying the hypotheses of Proposition 4.17.
Example 4.18.
Take ${\mathfrak{X}}={\mathfrak{M}}$ a von Neumann algebra and $\Gamma(T_{1},T_{2})=T_{1}T_{2}^{*}$, for all $T_{1},T_{2}\in{\mathfrak{M}}$. Let $W$ and $\eta_{x}(W)$ be as in Example 3.14. For every $T\in{\mathfrak{M}}$ and $x\in[0,\|W\|]$, let
$(\varphi(A,B)(T))(x)=\rho(A\,\eta_{x}(W)\,T\,\eta_{x}(W)\,B^{*}),\quad A,B\in L^{2}(\rho).$
By some calculations it is not hard to see that $\varphi(A,B)(T)(\cdot)$ is a continuous function on $[0,\|W\|]$ and that $\varphi$ is a bounded, left-invariant, sesquilinear map from $L^{2}(\rho)\times L^{2}(\rho)$ into $C([0,\|W\|])$ such that for every $N\in\mathbb{N}$ and $A_{1},\dots,A_{N}\in L^{2}(\rho)$, $T_{1},\dots,T_{N}\in{\mathfrak{M}}$, we have that
$\sum_{i,j=1}^{N}\varphi(A_{i},A_{j})(T_{i}T_{j}^{*})(\cdot)$
is a nonnegative continuous function on $[0,\|W\|]$. Let $M([0,\|W\|])$ be the Banach space of all complex positive Radon measures on $[0,\|W\|]$, $\Phi:L^{2}(\rho)\times L^{2}(\rho)\to{\mathfrak{B}}({\mathfrak{M}},M([0,\|W\|]))$ be given by
$d\Phi(A,B)(T)=\varphi(A,B)(T)(\cdot)d\mu,\quad\mu\in M([0,\|W\|]).$
Then, $\Phi$ is also a bounded, left-invariant, completely positive sesquilinear map. Finally, Let $\widetilde{{\mathfrak{M}}}$ be a von Neumann algebra with finite trace $\widetilde{\rho}$, and $\widetilde{{\mathfrak{M}}}^{*}$ be the dual of $\widetilde{{\mathfrak{M}}}$. Let $\widetilde{\varphi}:M([0,\|W\|])\to\widetilde{{\mathfrak{M}}}^{*}$ be the a bounded positive linear map constructed as in Example 4.15, then the composition
$\widetilde{\varphi}\circ\Phi:L^{2}(\rho)\times L^{2}(\rho)\to{\mathfrak{B}}({\mathfrak{M}},\widetilde{{\mathfrak{M}}}^{*})$
is a left-invariant, bounded, completely positive sesquilinear map.
All the results in Section 4 remain valid if we replace $\mathfrak{Y}$ by an arbitrary $C^{*}$-algebra due to the Cauchy-Schwarz-like inequality for $C^{*}$-valued positive sesquilinear maps in [5, Lemma 2,3] and due to the fact that the norm on a $C^{*}$-algebra is order-preserving.
Acknowledgements: GB and CT acknowledge that this work has been carried out within the activities of Gruppo UMI Teoria dell’Appros-simazione e Applicazioni and of GNAMPA of the INdAM. SI is supported by the Ministry of Science, Technological Development and Innovations, Republic of Serbia, grant no. 451-03-66/2024-03/200029.