### v5

# BANACH BIMODULE-VALUED POSITIVE MAPS: INEQUALITIES AND INDUCED REPRESENTATIONS

Abstract.
Representations induced by general positive sesquilinear maps with values in ordered Banach bimodules such as commutative and non-commutative $L^{1}$-spaces and the spaces of bounded linear operators from a $C^{*}$-algebra into the dual of another $C^{*}$-algebra are considered. As a starting point, a generalized Cauchy-Schwarz inequality is proved for these maps and a representation of bounded positive maps from a (quasi) *-algebra into such an ordered Banach bimodule is derived and some more inequalities for these maps are deduced. In particular, an extension of Paulsen’s modified Kadison-Schwarz inequality for 2-positive maps to the case of general positive maps from a unital *-algebra into the space of trace-class operators on a separable Hilbert space and into the duals of von-Neumann algebras is obtained. Also, representations for completely positive maps with values in an ordered Banach bimodule and Cauchy-Schwarz inequality for infinite sums of such maps are provided. Concrete examples illustrate the results.
Key words and phrases:
Representations, modules, positive sesquilinear maps, completely positive sesquilinear maps, normed spaces.
1991 Mathematics Subject Classification:
Primary 46K10, 47A07, 16D10.
1. Introduction
Positive and completely positive maps play an important role in the theory of operator algebras and quantum information, see for instance [9]. This motivate a study of representations involving these maps. A classical result in representation theory is the famous Stinespring theorem [23] (see also [2, 20, 24] which gives a representation of completely positive maps from a $C^{*}$-algebra into the space of bounded operators in Hilbert space. Quite recently, in [5, Corollary 3.10] and [6, Corollary 3.3], a representation of general positive $C^{*}$-valued maps on arbitrary *-algebras has also been provided. Now, $C^{*}$-algebras are just a special case of Banach quasi *-algebras and, moreover, every Banach quasi *-algebra is a Banach bimodule over a *-algebra. Several beautiful and important mathematical structures (such as commutative and noncommutative $L^{p}$-spaces) are examples of Banach quasi *-algebras. Further, Hilbert $C^{*}$-modules [17, 18] and the spaces of bounded linear operators from a $C^{*}$-algebra into the dual of another $C^{*}$-algebra are other examples of (ordered) Banach bimodules over a *-algebra which are of interest for applications. All these facts motivated us to study positive and completely positive maps from a Banach quasi *-algebra into an (ordered) Banach bimodule over a *-algebra, and the main aim of this paper is therefore obtaining representations and related inequalities of such maps. In a recent paper [5], we considered a GNS-like construction defined by positive sesquilinear maps on a quasi *-algebra taking their values in a $C^{*}$-algebra. In this paper we go some steps further and consider the possibility of replacing sesquilinear maps taking values in a $C^{*}$-algebra with sesquilinear maps with values in an ordered Banach bimodule $\mathfrak{Y}$ over some *-algebra $\mathfrak{Y}_{0}$. In this situation, an extension of the result of [5] is possible, provided that the sesquilinear form $\Phi$, where we start from, satisfies a generalized Cauchy-Schwarz inequality. This discussion is developed in Proposition 3.1 in Section 3 where we characterize a class of ordered Banach bimodules $\mathfrak{Y}$ with the property that every $\mathfrak{Y}$-valued positive sesquilinear map satisfies a generalized Cauchy-Schwarz inequality. Concrete examples of such bimodules are, for instance, non-commutative $L^{1}$-spaces and the space of bounded linear operators from a von Neumann algebra into the dual of another von Neumann algebra. As a corollary, we obtain an extension of Paulsen’s modified Kadison-Schwarz inequality for 2-positive maps between $C^{*}$-algebras to the case of general positive maps from a unital *-algebra into any of the above mentioned ordered Banach bimodules (Corollary 3.2) as well as Cauchy-Schwarz inequality for infinite sums of ordered Banach bimodule-valued positive sesquilinear maps (Proposition 3.3). Then we provide a representation of positive sesquilinear maps from a quasi *-algebra into such an ordered Banach bimodule. Moreover, we give examples of such maps. In the last part of Section 3, we obtain also a representation of positive ordered Banach bimodule-valued linear maps on a unital *-algebra (Corollary 3.2). An important difference between representations of ordered Banach bimodule-valued positive maps given in Section 3 and the representations of $C^{*}$-valued positive maps given in [5] is that the ones from Section 3 are given in a proper Banach space and not just in a quasi-Banach space (as in [5]). This is due to the Cauchy-Schwarz inequality for positive Banach bimodule-valued sesquilinear maps given in Proposition 3.1. Positive maps from a unital $C^{*}$-algebra into the space of trace-class operators are a special case of the positive maps treated in Section 3. Since the space of trace-class operators is the subspace of the $C^{*}$-algebra of bounded linear operators on a Hilbert space, it follows that our results have an application in the theory of positive maps on $C^{*}$-algebras. Section 4 is devoted to the *-representations induced by completely positive sesquilinear maps. This is actually a subject for which several approaches have been proposed in view of a possible generalization of the Stinespring theorem to different environments (see, for instance, [3, 14, 12, 15, 21, 11, 19]). Motivated by [21, Proposition 3.1], following an idea proposed in [22, Ch. 11] and adopted in [4] for partial *-algebras, we consider completely positive sesquilinear maps from a normed quasi *-algebra $\mathfrak{A}$ into a set of sesquilinear maps on a vector space which take values in an ordered Banach bimodule $\mathfrak{Y}$ and provide a generalization of the classical Stinespring theorem in this context, as well as applications of this result to Cauchy-Schwarz inequality for infinite sums of such maps (Corollary 4.3). Further, we provide a version of Radon-Nikodym theorem for these maps as well as for ordered Banach bimodule-valued bounded positive sesquilinear maps on a quasi *-algebra, thus generalizing, in this setting, the results from [24, Section 3.5] and give examples of such maps. At the end of the section, we apply our results to obtain representations of bounded completely positive sesquilinear maps on a quasi *-algebra with values in the space of ordered Banach bimodule-valued, bounded, linear operators. As a consequence of these representations, in Proposition 4.12 we obtain Cauchy-Schwarz inequalities for such maps. Also, we provide some examples as well as some other applications.
2. Preliminaries
A quasi *-algebra $(\mathfrak{A},\mathfrak{A}_{0})$ is a pair consisting of a vector space $\mathfrak{A}$ and a *-algebra $\mathfrak{A}_{0}$ contained in $\mathfrak{A}$ as a subspace and such that
  • $\mathfrak{A}$ carries an involution $a\mapsto a^{*}$ extending the involution of $\mathfrak{A}_{0}$;
  • $\mathfrak{A}$ is a bimodule over $\mathfrak{A}_{0}$ and the module multiplications extend the multiplication of $\mathfrak{A}_{0}$. In particular, the following associative laws hold:
    $(ca)d=c(ad);\ \ a(cd)=(ac)d,\quad\forall\ a\in\mathfrak{A},\ c,d\in\mathfrak{A}_{0};$
  • $(ac)^{*}=c^{*}a^{*}$, for every $a\in\mathfrak{A}$ and $c\in\mathfrak{A}_{0}$.
The identity or unit element of $(\mathfrak{A},\mathfrak{A}_{0})$, if any, is a necessarily unique element ${\sf e}\in\mathfrak{A}_{0}$, such that $a{\sf e}=a={\sf e}a$, for all $a\in\mathfrak{A}$.
We will always suppose that
$ac=0,\;\forall c\in\mathfrak{A}_{0}\Rightarrow a=0$
$ac=0,\;\forall a\in\mathfrak{A}\Rightarrow c=0.$
Clearly, both these conditions are automatically satisfied if $(\mathfrak{A},\mathfrak{A}_{0})$ has an identity ${\sf e}$.
[…]
  • there exists $\gamma>0$ such that for every $a\in\mathfrak{A}$
    $\max\{\|ac\|,\|ca\|\}\leq\gamma\|a\|,\quad\forall c\in\mathfrak{A}_{0};$
  • $\|a^{*}\|=\|a\|,\;\forall a\in\mathfrak{A};$
  • $\mathfrak{A}_{0}$ is dense in $\mathfrak{A}[\|\cdot\|]$.
If the normed vector space $\mathfrak{A}[\|\cdot\|]$ is complete, then $(\mathfrak{A},\mathfrak{A}_{0})$ is called a Banach quasi *-algebra. We refer to [10] for further details.
Definition 2.1.
Let $\mathfrak{Y}$ be a Banach bimodule over the *-algebra $\mathfrak{Y}_{0}$. We say that $\mathfrak{Y}$ is an ordered Banach bimodule over $\mathfrak{Y}_{0}$ if
  • (i)
    $\mathfrak{Y}$ is ordered as a vector space; that is, $\mathfrak{Y}$ contains a (positive) closed cone $\mathfrak{K}$, i.e., $\mathfrak{K}\subset\mathfrak{Y}$ is such that $\mathfrak{K}+\mathfrak{K}\subset\mathfrak{K}$, $\lambda\mathfrak{K}\subset\mathfrak{K}$ for $\lambda\geq 0$ and $\mathfrak{K}\cap(-\mathfrak{K})=\{0\}$;
  • (ii)
    $z^{*}\mathfrak{K}z\subset\mathfrak{K},\quad\forall z\in\mathfrak{Y}_{0}.$
As usual, we will write $y_{1}\leq y_{2}$ whenever $y_{2}-y_{1}\in\mathfrak{K}$, with $y_{1},y_{2}\in\mathfrak{Y}$ and we will sometimes suppose that $\mathfrak{Y}$ has an order-preserving norm in the sense that if $y_{1}\leq y_{2}$ with $y_{1},y_{2}\in\mathfrak{Y}$, then also $\|y_{1}\|_{\mathfrak{Y}}\leq\|y_{2}\|_{\mathfrak{Y}}$. Throughout the paper, $\Omega$ will denote a (locally) compact Hausdorff space and $C(\Omega)$ and $M(\Omega)$ will denote the space of continuous functions on $\Omega$ and the Banach space of all complex Radon measures on $\Omega$ equipped with the total variation norm, respectively.
Example 2.2.
Examples of ordered Banach bimodules over a Banach *-algebra are provided by:
  • if $\mathcal{H}$ is a separable Hilbert space, and ${\mathfrak{B}}_{1}(\mathcal{H})$ denotes the space of all trace-class operators, then $\mathfrak{Y}=({\mathfrak{B}}_{1}(\mathcal{H}),\|\cdot\|_{1})$ is a Banach bimodule over the von Neumann algebra $\mathfrak{Y}_{0}=({\mathfrak{B}}(\mathcal{H}),\|\cdot\|)$ of bounded operators on $\mathcal{H}$;
  • the non commutative spaces $\mathfrak{Y}=L^{p}(\rho)$, over the Banach *-algebra $\mathfrak{Y}_{0}=L^{\infty}(\rho)$, with $\rho$ a faithful finite trace on a von Neumann algebra ${\mathfrak{M}}$, see [10, Section 5.6.1];
  • $\mathfrak{Y}={\mathfrak{B}}(\mathfrak{C}_{1},\mathfrak{C}_{2})$, with $\mathfrak{C}_{1}$ and $\mathfrak{C}_{2}$ being $C^{*}$-algebras with unit, $\mathfrak{Y}_{0}=\mathfrak{C}_{1}$. If $A\in{\mathfrak{B}}(\mathfrak{C}_{1},\mathfrak{C}_{2})$ and $c\in\mathfrak{C}_{1}$, the module multiplication can be defined as
    $(A\cdot c)(d)=A(cd)\mbox{ and }(c\cdot A)(d)=A(dc),\quad d\in\mathfrak{C}_{1}.$
    Similar construction applies if we instead of $\mathfrak{C}_{2}$ consider the dual of $\mathfrak{C}_{2}$;
  • $\mathfrak{Y}=\mathfrak{Y}_{0}=\ell_{2}(C(\Omega))$, with $\ell_{2}(C(\Omega))$ the space of squared summable sequences of functions in $C(\Omega)$ with $\Omega$ a compact Hausdorff space. It is a Banach bimodule over itself because it is a (non-unital) normed *-algebra.
The definitions of the positive cones are obvious in all these cases.
In the next results we will need the additional assumption that the norm on the ordered Banach bimodule $\mathfrak{Y}$ is order-preserving. Examples of Banach bimodules with an order-preserving norm are commutative $L^{p}$-spaces and $\ell_{2}(C(\Omega))$, with $\Omega$ a compact Hausdorff space. Furthermore, if $\mathfrak{C}_{1}$ and $\mathfrak{C}_{2}$ are $C^{*}$-algebras with unit $e_{\mathfrak{C}_{1}}$ and $e_{\mathfrak{C}_{2}}$, respectively, the Banach ordered bimodule ${\mathfrak{B}}(\mathfrak{C}_{1},\mathfrak{C}_{2})$ is order-preserving (this follows from the fact that $\|\varphi\|=\|\varphi(e_{\mathfrak{C}_{1}})\|_{\mathfrak{C}_{2}}\leq\|\psi(e_{\mathfrak{C}_{1}})\|_{\mathfrak{C}_{2}}=\|\psi\|$, with $\varphi,\psi\in{\mathfrak{B}}(\mathfrak{C}_{1},\mathfrak{C}_{2})^{+}$ and $\varphi\leq\psi$, see e.g. [24]). Moreover, if $\mathcal{H}$ is a separable Hilbert space, the norm in ${\mathfrak{B}}_{1}(\mathcal{H})$ is clearly order-preserving.
Let ${\mathfrak{X}}$ be a vector space and $\mathfrak{Y}$ an ordered Banach bimodule over the *-algebra $\mathfrak{Y}_{0}$, with positive cone $\mathfrak{K}$. Let $\Phi$ be a $\mathfrak{Y}$-valued positive sesquilinear map on ${\mathfrak{X}}\times{\mathfrak{X}}$
$\Phi:(x_{1},x_{2})\in{\mathfrak{X}}\times{\mathfrak{X}}\to\Phi(x_{1},x_{2})\in\mathfrak{Y}$
i.e., a map with the properties
  • $i)$
    $\Phi(x_{1},x_{1})\in\mathfrak{K}$,
  • $ii)$
    $\Phi(\alpha x_{1}+\beta x_{2},\gamma x_{3})=\overline{\gamma}[\alpha\Phi(x_{1},x_{3})+\beta\Phi(x_{2},x_{3})]$,
with $x_{1},x_{2},x_{3}\in{\mathfrak{X}}$ and $\alpha,\beta,\gamma\in\mathbb{C}$.
The $\mathfrak{Y}$-valued positive sesquilinear map $\Phi$ is called faithful if
$\Phi(x,x)=0_{\mathfrak{Y}}\;\Rightarrow\;x=0.$
3. Representations induced by positive maps
It is quite apparent that the cornerstone of every extension of the GNS-construction is a Cauchy-Schwarz inequality. We begin with presenting some results in this direction.
Throughout this section we will assume that $\mathfrak{C}$ is a $C^{*}$-algebra with positive convex cone $\mathfrak{C}^{+}$ and norm $\|\cdot\|_{\mathfrak{C}}$, $\mathfrak{Y}$ is (an ordered) Banach bimodule over a *-algebra $\mathfrak{Y}_{0}$ (with $\mathfrak{Y}_{0}$ equipped with a not necessarily sub-multiplicative norm $\|\cdot\|_{\mathfrak{Y}_{0}}$) with the respective cones $\mathfrak{K}$ and $\mathfrak{K}_{0}$, with $\mathfrak{K}_{0}=\{\sum_{i=1}^{N}z_{i}^{*}z_{i},\,z_{i}\in\mathfrak{Y}_{0},i=1,\dots,N;N\in\mathbb{N}\}$.
From now on, by $\rho$ we will denote a faithful semifinite trace on a von Neumann algebra ${\mathfrak{M}}$.
Proposition 3.1.
Let ${\mathfrak{X}}$ be a complex vector space. Assume that $\mathfrak{Y}$, $\mathfrak{Y}_{0}$, $\mathfrak{K}$ and $\mathfrak{C}$ are as above. Let $\Phi:{\mathfrak{X}}\times{\mathfrak{X}}\to\mathfrak{Y}$ be a positive sesquilinear map. Let $\Omega,\Upsilon$ be locally compact Hausdorff spaces, ${\mathfrak{M}}$ a von Neumann algebra and $\rho$ a faithful trace on ${\mathfrak{M}}$. Then, the Cauchy-Schwarz inequality holds in each one of the following cases
  1. (1)
    $\mathfrak{Y}=L^{2}(\Omega)$,
  2. (2)
    $\mathfrak{Y}={\mathfrak{B}}({\mathfrak{M}},{\mathfrak{B}}_{1}(\mathcal{H}))$, $\mathfrak{Y}_{0}={\mathfrak{M}}$,
  3. (3)
    $\mathfrak{Y}={\mathfrak{M}}^{*}$, the dual of ${\mathfrak{M}}$,
  4. (4)
    $\mathfrak{Y}=M(\Omega)$ and $\mathfrak{Y}_{0}=C(\Omega)$,
  5. (5)
    $\mathfrak{Y}={\mathfrak{B}}({\mathfrak{M}},M(\Omega))$, $\mathfrak{Y}_{0}={\mathfrak{M}}$,
  6. (6)
    $\mathfrak{Y}={\mathfrak{B}}(C(\Omega),C(\Upsilon))$, or $\mathfrak{Y}={\mathfrak{B}}({\mathfrak{M}},C(\Omega))$,
  7. (7)
    $\mathfrak{Y}=L^{1}(\rho)$ or $\mathfrak{Y}=L^{1}(\Omega)$,
  8. (8)
    $\mathfrak{Y}={\mathfrak{B}}({\mathfrak{M}},\widetilde{{\mathfrak{M}}}^{*})$ and $\mathfrak{Y}_{0}={\mathfrak{M}}$, with $\widetilde{{\mathfrak{M}}}$ another von Neumann algebra,
  9. (9)
    $\mathfrak{Y}=\mathfrak{Y}_{0}=\ell_{2}(C(\Omega))$.
Proof.
We just give a sketch of the proof of each case.
[…]
Then consider a sequence $\{S_{n}\}_{n}$ of simple functions such that $S_{n}\to\overline{\Phi(x_{1},x_{2})}$ in $L^{2}(\Omega,\mu)$ as $n\to\infty$. By taking limits in the above inequality, we get the Cauchy-Schwarz inequality.
(2) Let $P\in{\mathfrak{M}}$ and $Q\in{\mathfrak{B}}(\mathcal{H})$ be orthogonal projections. Define $\varphi_{P,Q}:{\mathfrak{X}}\times{\mathfrak{X}}\to\mathbb{C}$ by
(3.1) $\varphi_{P,Q}(x_{1},x_{2})=tr(Q(\Phi(x_{1},x_{2})(P))),\quad\forall x_{1},x_{2}\in{\mathfrak{X}}.$
Then $\varphi_{P,Q}$ is a positive sesquilinear form; hence, using the Cauchy-Schwarz inequality, we get
(3.2) $|tr(Q(\Phi(x_{1},x_{2})(P)))|\\ \leq(tr(Q(\Phi(x_{1},x_{1})(P))))^{1/2}(tr(Q(\Phi(x_{2},x_{2})(P))))^{1/2}$
for all $x_{1},x_{2}\in{\mathfrak{X}}$. Let now $S\in{\mathfrak{M}}$ and $T\in{\mathfrak{B}}(\mathcal{H})$ be two simple operator functions, so $S=\sum_{i=1}^{N}\alpha_{i}P_{i}$ and $T=\sum_{j=1}^{M}\beta_{j}Q_{j}$ with $P_{i_{1}}P_{i_{2}}=Q_{j_{1}}Q_{j_{2}}=0$ whenever $i_{1}\neq i_{2}$ and $j_{1}\neq j_{2}$, $\sum_{i=1}^{N}P_{i}={{\mathbb{I}}_{{\mathfrak{M}}}}$ and $\sum_{j=1}^{M}Q_{j}={{\mathbb{I}}_{\mathcal{H}}}$. By (3.2) and some computations we get that
(3.3) ||| $tr(T(\Phi(x_{1},x_{2})(S)))|\leq\sum_{i,j}|\alpha_{i}||\beta_{j}||tr(Q_{j}(\Phi(x_{1},x_{2})(P_{i})))|$
$\qquad\leq\|T\|\|S\|(tr(\Phi(x_{1},x_{1})({\mathbb{I}}_{{\mathfrak{M}}})))^{1/2}(tr(\Phi(x_{2},x_{2})({\mathbb{I}}_{{\mathfrak{M}}})))^{1/2}$
$\qquad\leq\|T\|\|S\|\|\Phi(x_{1},x_{1})\|^{1/2}\|\Phi(x_{2},x_{2})\|^{1/2},\quad\forall x_{1},x_{2}\in{\mathfrak{X}}.$
Let now $U\in{\mathfrak{M}}$ and $W\in{\mathfrak{B}}(\mathcal{H})$ be unitary. Then, by the spectral theorem, we can find sequences $\{S_{n}\}_{n}\subset{\mathfrak{M}}$ and $\{T_{n}\}_{n}\subset{\mathfrak{B}}(\mathcal{H})$ that converge in the operator norm to $U$ and $W$, respectively.
Then, since
(3.4) ||| $tr(B(\Phi(x_{1},x_{2})(A)))|\leq tr(|B(\Phi(x_{1},x_{2})(A))|)$
$\qquad\leq\|B\|\|\Phi(x_{1},x_{2})(A)\|_{1}\leq\|B\|\|A\|\|\Phi(x_{1},x_{2})\|,$
for all $A\in{\mathfrak{M}}$, $B\in{\mathfrak{B}}(\mathcal{H})$, $x_{1},x_{2}\in{\mathfrak{X}}$, it follows that
$tr(T_{n}(\Phi(x_{1},x_{2})(S_{n})))\to tr(W(\Phi(x_{1},x_{2})(U))),\,\mbox{ as}n\to\infty.$
Also, $\|T_{n}\|\|S_{n}\|\to\|W\|\|U\|=1$ as $n\to\infty$, hence, by using inequality (3.3), passing to the limits, we deduce that
(3.5) $|tr(W(\Phi(x_{1},x_{2})(U)))|\leq\|\Phi(x_{1},x_{1})\|^{1/2}\|\Phi(x_{2},x_{2})\|^{1/2},$
for all $x_{1},x_{2}\in{\mathfrak{X}}$. This inequality extends easily to convex combinations $\sum_{i=1}^{N}\lambda_{i}U_{i}$, $\sum_{j=1}^{M}\gamma_{j}W_{j}$.
The Russo-Dye theorem then allows to get the following inequality for $A\in{\mathfrak{M}}$ and $B\in{\mathfrak{B}}(\mathcal{H})$ with $\|A\|,\|B\|<1$,
(3.6) $|tr(B(\Phi(x_{1},x_{2})(A)))|\\ \leq\|\Phi(x_{1},x_{1})\|^{1/2}\|\Phi(x_{2},x_{2})\|^{1/2},\quad\forall x_{1},x_{2}\in{\mathfrak{X}}.$
Finally, let $A\in{\mathfrak{M}}$ with $\|A\|=1$, and $V\in{\mathfrak{B}}(\mathcal{H})$ be the partial isometry with $V(\Phi(x_{1},x_{2})(A))=|\Phi(x_{1},x_{2})(A)|$ and choose sequences $\{A_{n}\}_{n}\subset{\mathfrak{M}}$ and $\{V_{n}\}_{n}\subset{\mathfrak{B}}(\mathcal{H})$, with $\|A_{n}\|,\|V_{n}\|<1$ for all $n$, such that $A_{n}\to A$ and $V_{n}\to V$ in the operator norm as $n\to\infty$. Once again, by using (3.4) and a limit procedure we get
$\|\Phi(x_{1},x_{2})(A)\|_{1}=tr(V(\Phi(x_{1},x_{2})(A)))\leq\|\Phi(x_{1},x_{1})\|^{1/2}\|\Phi(x_{2},x_{2})\|^{1/2},$
for all $x_{1},x_{2}\in{\mathfrak{X}}$. Finally, by taking the supremum over the unit ball in ${\mathfrak{M}}$, we conclude that
$\|\Phi(x_{1},x_{2})\|\leq\|\Phi(x_{1},x_{1})\|^{1/2}\|\Phi(x_{2},x_{2})\|^{1/2},\quad\forall x_{1},x_{2}\in{\mathfrak{X}}.$
(3) It is a special case of $(2)$: if $\mathcal{H}$ is one dimensional, then ${\mathfrak{B}}_{1}(\mathcal{H})=\mathbb{C}$.
(4) We can apply the Cauchy-Schwarz inequality to every positive sesquilinear form $\varphi_{E}(x_{1},x_{2})=\Phi(x_{1},x_{2})(E)$, with $E$ a given Borel subset of $\Omega$, and therefore by considering a partition $\{E_{n}\}_{n=1}^{N}$ of $\Omega$ we obtain
$\sum_{n=1}^{N}|\Phi(x_{1},x_{2})(E_{n})|$ $\leq\sum_{n=1}^{N}(\Phi(x_{1},x_{1})(E_{n}))^{1/2}(\Phi(x_{2},x_{2})(E_{n}))^{1/2}$
$\leq\left(\sum_{n=1}^{N}\Phi(x_{1},x_{1})(E_{n})\right)^{\frac{1}{2}}\left(\sum_{n=1}^{N}\Phi(x_{2},x_{2})(E_{n})\right)^{\frac{1}{2}}$
hence, by taking the supremum of both sides, we get the Cauchy-Schwarz inequality.
(5) Let $P\in{\mathfrak{M}}$ be an orthogonal projection and $E\subset$ be a Borel subset of $\Omega$. The map $\varphi_{P,E}:{\mathfrak{X}}\times{\mathfrak{X}}\to\mathbb{C}$ given by
$\varphi_{P,E}(x_{1},x_{2})=\Phi(x_{1},x_{2})(P)(E),\quad\forall x_{1},x_{2}\in{\mathfrak{X}}$
is a positive sesquilinear form, hence, by similar arguments as in the proof of $(4)$ we can deduce that
$\|\Phi(x_{1},x_{2})(P)\|\leq(\Phi(x_{1},x_{1})(P))^{1/2}(\Phi(x_{2},x_{2})(P))^{1/2},\quad\forall x_{1},x_{2}\in{\mathfrak{X}}.$
Since this holds for an arbitrary orthogonal projection $P\in{\mathfrak{M}}$, we can then proceed further in a similar way as in the proof of $(2)$ and deduce Cauchy-Schwarz inequality for $\Phi$. We leave the details to the reader.
(6) The map $\widetilde{\varphi}_{t}:{\mathfrak{X}}\times{\mathfrak{X}}\to(C(\Omega))^{*}\cong M(\Omega)$ given by
$\widetilde{\varphi}_{t}(x_{1},x_{2})(f)=(\Phi(x_{1},x_{2})(f))(t),\quad\forall f\in C(\Omega),x_{1},x_{2}\in{\mathfrak{X}}$
is sesquilinear and positive, hence, the Cauchy-Schwarz inequality holds for $\widetilde{\varphi}_{t}$ by case (4). Therefore, for every $f\in C(\Omega)$, with $\|f\|\leq 1$ and every $t\in\Upsilon$, we get
$\left|(\Phi(x_{1},x_{2})(f))(t)\right|\leq\|\widetilde{\varphi}_{t}(x_{1},x_{2})\|\leq\|\widetilde{\varphi}_{t}(x_{1},x_{1})\|^{1/2}\|\widetilde{\varphi}_{t}(x_{2},x_{2})\|^{1/2}$
$\qquad\quad\leq\left(\sup_{\begin{subarray}{c}g\in C(\Omega),\\ \|g\|_{\infty}\leq 1\end{subarray}}\|\Phi(x_{1},x_{1})(g)\|_{\infty}\right)^{1/2}\left(\sup_{\begin{subarray}{c}h\in C(\Omega),\\ \|h\|_{\infty}\leq 1\end{subarray}}\|\Phi(x_{2},x_{2})(g)\|_{\infty}\right)^{1/2}$
$\qquad\quad\leq\|\Phi(x_{1},x_{1})\|^{1/2}\|\Phi(x_{2},x_{2})\|^{1/2}.$
The statement follows by taking supremums.
The case $\mathfrak{Y}={\mathfrak{B}}({\mathfrak{M}},C(\Omega))$ can be treated similarly by using (3).
(7) Consider the map $\iota:L^{1}(\rho)\to\widetilde{{\mathfrak{M}}}^{*}$ given by
$\iota(T)(S)=\rho(ST),\quad\forall\,T\in L^{1}(\rho),S\in{\mathfrak{M}}.$
It is not hard to check that $\iota$ is an isometry. Moreover, since $\iota(T)(S)=\rho(S^{1/2}TS^{1/2})\geq 0$ whenever $S$ and $T$ are positive, it follows that $\iota$ preserves positivity. Hence, if $\Phi:{\mathfrak{X}}\times{\mathfrak{X}}\to L^{1}(\rho)$ is a positive sesquilinear map, then $\iota\circ\Phi:{\mathfrak{X}}\times{\mathfrak{X}}\to{\mathfrak{M}}^{*}$ is also a positive sesquilinear map. By (3) and since $\iota$ is an isometry, we get that
$\|\Phi(x_{1},x_{2})\|_{1}$ $=\|\iota\circ\Phi(x_{1},x_{2})\|\leq\|\iota\circ\Phi(x_{1},x_{1})\|^{1/2}\|\iota\circ\Phi(x_{2},x_{2})\|^{1/2}$
$=\|\Phi(x_{1},x_{2})\|_{1}^{1/2}\|\Phi(x_{1},x_{2})\|_{1}^{1/2},\quad\forall x_{1},x_{2}\in{\mathfrak{X}}.$
Of course, an analogous result holds in the commutative case (i.e. $\mathfrak{Y}=L^{1}(\Omega)$).
(8) In the proof of $(2)$, given $P\in{\mathfrak{M}}$ and $Q\in{\mathfrak{B}}(\mathcal{H})$, we have considered the sesquilinear form $\varphi_{P,Q}$ in (3.1). Now, given $S\in{\mathfrak{M}}$ and $T\in\widetilde{{\mathfrak{M}}}$, we can consider, instead, the sesquilinear form $\varphi_{S,T}:{\mathfrak{X}}\times{\mathfrak{X}}\to\mathbb{C}$ given by
$\varphi_{S,T}(x_{1},x_{2})=\Phi(x_{1},x_{2})(S)(T),\quad\forall x_{1},x_{2}\in{\mathfrak{X}}$
and proceed analogously to the proof of $(2)$.
(9) We can apply the Cauchy-Schwarz inequality to all the positive sesquilinear forms $\varphi_{n,t}(x_{1},x_{2})=(\Phi(x_{1},x_{2}))_{n}(t)$ with $(\Phi(x_{1},x_{2}))_{n}$ the n-th component of the sequence $\Phi(x_{1},x_{2})$ and then applying the Cauchy-Schwarz inequality for the inner product in $\ell_{2}(C(\Omega))$. ∎
From now on $\mathfrak{Y}$ will denote any Banach bimodule considered in Proposition 3.1. The next corollary is motivated by modified Schwarz inequality for 2-positive maps (see [20, Ch. 3, Ex. 3.4, p.40] and [8, Lemma 2.6])
Corollary 3.2.
If $\mathfrak{A}$ is a unital *-algebra and $\omega:\mathfrak{A}\to\mathfrak{Y}$ is a positive linear map, then
$\|\omega(b^{*}a)\|_{\mathfrak{Y}}\leq\|\omega(a^{*}a)\|_{\mathfrak{Y}}^{1/2}\|\omega(b^{*}b)\|_{\mathfrak{Y}}^{1/2},\quad\forall a,b\in\mathfrak{A}.$
Proof.
This follows from Proposition 3.1 by considering the sesquilinear map $\Phi:\mathfrak{A}\times\mathfrak{A}\to\mathfrak{Y}$ given by $\Phi(a,b)=\omega(b^{*}a)$, for all $a,b\in\mathfrak{A}$. ∎
Now we give an extension of Proposition 3.1.
Proposition 3.3.
Let ${\mathfrak{X}}$ be a vector space and assume that the norm on $\mathfrak{Y}$ is order-preserving. If $\{\Phi_{n}\}_{n}$ is a sequence of positive sesquilinear maps $\Phi_{n}:{\mathfrak{X}}\times{\mathfrak{X}}\to\mathfrak{Y}$ and $\{x_{n}\}_{n}$, $\{\widetilde{x}_{n}\}_{n}$ are sequences in ${\mathfrak{X}}$ such that the series $\sum_{n=1}^{\infty}\Phi_{n}(x_{n},x_{n})$ and $\sum_{n=1}^{\infty}\Phi_{n}(\widetilde{x}_{n},\widetilde{x}_{n})$ are convergent in $\mathfrak{Y}$ then
$\left\|\sum_{n=1}^{\infty}\Phi_{n}(x_{n},\widetilde{x}_{n})\right\|_{\mathfrak{Y}}\leq\left\|\sum_{n=1}^{\infty}\Phi_{n}(x_{n},x_{n})\right\|_{\mathfrak{Y}}^{1/2}\left\|\sum_{n=1}^{\infty}\Phi_{n}(\widetilde{x}_{n},\widetilde{x}_{n})\right\|_{\mathfrak{Y}}^{1/2}.$
Proof.
The statement can be proved by arguments similar to the first part of [18, Example 1.3.5] due to Proposition 3.1 and the fact that the norm of $\mathfrak{Y}$ is order-preserving. ∎
For a detailed overview of some other operator-inequalities for positive maps we refer to [7, 16].
Definition 3.4.
Let ${\mathfrak{X}}$ be a vector space. A $\mathfrak{Y}$-valued faithful positive sesquilinear map $\Phi$ on ${\mathfrak{X}}\times{\mathfrak{X}}$ is said to be a $\mathfrak{Y}$-valued inner product on ${\mathfrak{X}}$ and we often write $\langle{x_{1}}|{x_{2}}\rangle_{\Phi}:=\Phi(x_{1},x_{2})$, $x_{1},x_{2}\in{\mathfrak{X}}$.
A $\mathfrak{Y}$-valued inner product on ${\mathfrak{X}}$
$\Phi:{\mathfrak{X}}\times{\mathfrak{X}}\to\mathfrak{Y}$
induces a norm $\|\cdot\|_{\Phi}$ on ${\mathfrak{X}}$:
$\|x\|_{\Phi}:=\sqrt{\|\langle{x}|{x}\rangle_{\Phi}\|_{\mathfrak{Y}}}=\sqrt{\|\Phi(x,x)\|_{\mathfrak{Y}}},\quad x\in{\mathfrak{X}},$
since
$\|x_{1}+x_{2}\|_{\Phi}\leq\|x_{1}\|_{\Phi}+\|x_{2}\|_{\Phi},\qquad\forall x_{1},x_{2}\in{\mathfrak{X}};$
this can be shown similarly to what done in [5].
The space ${\mathfrak{X}}$ is then a normed space w.r.to the norm $\|\cdot\|_{\Phi}$.
Definition 3.5.
Let ${\mathfrak{X}}$ be a complex vector space and $\Phi$ be a $\mathfrak{Y}$-valued inner product on ${\mathfrak{X}}$. If ${\mathfrak{X}}$ is complete w.r. to the norm $\|\cdot\|_{\Phi}$, then ${\mathfrak{X}}$ is called a Banach space with $\mathfrak{Y}$-valued inner product or for short a $B_{\mathfrak{Y}}$-space.
If $\Phi$ is not faithful, we can consider the subspace of ${\mathfrak{X}}$
$\mathfrak{N}_{\Phi}=\{x_{1}\in{\mathfrak{X}}:\,\Phi(x_{1},x_{2})=0_{\mathfrak{Y}},\forall x_{2}\in{\mathfrak{X}}\}.$
By Proposition 3.1, then
$\mathfrak{N}_{\Phi}=\{x\in{\mathfrak{X}}:\,\Phi(x,x)=0_{\mathfrak{Y}}\}.$
We denote by $\Lambda_{\Phi}(x)$ the coset of ${\mathfrak{X}}/\mathfrak{N}_{\Phi}$ containing $x\in{\mathfrak{X}}$; i.e., $\Lambda_{\Phi}(x)=x+\mathfrak{N}_{\Phi}$ and define a $\mathfrak{Y}$-valued inner product on ${\mathfrak{X}}/\mathfrak{N}_{\Phi}$ as follows:
$\langle{\Lambda_{\Phi}(x_{1})}|{\Lambda_{\Phi}(x_{2})}\rangle_{\Phi}:=\Phi(x_{1},x_{2}),\quad x_{1},x_{2}\in{\mathfrak{X}}.$
The associated norm is:
$\|\Lambda_{\Phi}(x)\|_{\Phi}:=\sqrt{\|\Phi(x,x)\|_{\mathfrak{Y}}},\quad x\in{\mathfrak{X}}.$
The quotient space ${\mathfrak{X}}/\mathfrak{N}_{\Phi}=\Lambda_{\Phi}({\mathfrak{X}})$ is a normed space (see [5]).
Let $\mathcal{K}$ be a $B_{\mathfrak{Y}}$-space and $D(T)$ be a dense subspace of $\mathcal{K}$. A linear map $T:D(T)\to\mathcal{K}$ is said $\Phi$-adjointable if there exists a linear map $T^{*}$ defined on a subspace $D(T^{*})\subset\mathcal{K}$ such that
$\Phi(T\xi,\eta)=\Phi(\xi,T^{*}\eta),\quad\forall\xi\in D(T),\eta\in D(T^{*}).$
Let $\mathcal{D}$ be a dense subspace of $\mathcal{K}$ and let us consider the following families of linear operators acting on $\mathcal{D}$:
${{\mathcal{L}}^{\mbox{\scriptsize${\dagger}$}}(\mathcal{D},\mathcal{K})}$ =$=\{T\mbox{ $\Phi$-adjointable},D(T)=\mathcal{D};\;D(T^{*})\supset\mathcal{D}\}$
${{\mathcal{L}}^{\dagger}(\mathcal{D})}$ =$=\{T\in{\mathcal{L}}^{\mbox{\scriptsize${\dagger}$}}(\mathcal{D},\mathcal{K}):T\mathcal{D}\subset\mathcal{D};\;T^{*}\mathcal{D}\subset\mathcal{D}\}.$
The involution in ${\mathcal{L}}^{\mbox{\scriptsize${\dagger}$}}(\mathcal{D},\mathcal{K})$ is defined by $T^{\dagger}:=T^{*}\upharpoonright\mathcal{D}$, the restriction of $T^{*}$, the $\Phi$-adjoint of $T$, to $\mathcal{D}$. The set ${\mathcal{L}}^{\dagger}(\mathcal{D})$ is a *-algebra.
Remark 3.6.
[…]
$\Phi(T_{2}\xi,T_{1}^{\dagger}\eta)=\Phi(W\xi,\eta),\quad\forall\xi,\eta\in\mathcal{D}.$
Due to the density of $\mathcal{D}$ in $\mathcal{K}$, the element $W$, if it exists, is unique. We put $W=T_{1}\,{\scriptstyle\square}\,T_{2}$.
Definition 3.7.
Let $(\mathfrak{A},\mathfrak{A}_{0})$ be a quasi *-algebra with unit ${\sf e}$. Let $\mathcal{D}$ be a dense subspace of a certain $B_{\mathfrak{Y}}$-space $\mathcal{K}$ with $\mathfrak{Y}$-valued inner product $\langle{\cdot}|{\cdot}\rangle_{\mathcal{K}}$. A linear map $\pi$ from $\mathfrak{A}$ into ${\mathcal{L}}^{\mbox{\scriptsize${\dagger}$}}(\mathcal{D},\mathcal{K})$ is called a *-representation of $(\mathfrak{A},\mathfrak{A}_{0})$, if the following properties are fulfilled:
  • (i)
    $\pi(a^{*})=\pi(a)^{\dagger}:=\pi(a)^{*}\upharpoonright\mathcal{D},\quad\forall\ a\in\mathfrak{A}$;
  • (ii)
    for $a\in\mathfrak{A}$ and $c\in\mathfrak{A}_{0}$, $\pi(a)\,{\scriptstyle\square}\,\pi(c)$ is well-defined and $\pi(a)\,{\scriptstyle\square}\,\pi(c)=\pi(ac)$.
We assume that for every *-representation $\pi$ of $(\mathfrak{A},\mathfrak{A}_{0})$, $\pi({\sf e})={{\mathbb{I}}_{\mathcal{D}}}$, the identity operator on the space $\mathcal{D}$.
The *-representation $\pi$ is said to be
  • closable if there exists $\widetilde{\pi}$ the closure of $\pi$, defined as $\widetilde{\pi}(a)=\overline{\pi(a)}\upharpoonright{\widetilde{\mathcal{D}}}$ where $\widetilde{\mathcal{D}}$ is the completion under the graph topology $t_{\pi}$ defined by the seminorms $\xi\in\mathcal{D}\to\|\xi\|_{\mathcal{K}}+\|\pi(a)\xi\|_{\mathcal{K}}$, $a\in\mathfrak{A}$, with $\|\cdot\|_{\mathcal{K}}$ the norm induced by the inner product on $\mathcal{K}$;
  • closed if $\mathcal{D}[t_{\pi}]$ is complete;
  • cyclic if there exists $\xi\in\mathcal{D}$ such that $\pi(\mathfrak{A}_{0})\xi$ is dense in $\mathcal{K}$ in its norm topology. In this case $\xi$ is called cyclic vector.
Definition 3.8.
Let $(\mathfrak{A},\mathfrak{A}_{0})$ be a quasi *-algebra. We denote by ${\mathcal{I}}_{{\mathfrak{A}}_{0}}^{\,\mathfrak{Y}}(\mathfrak{A})$ the set of all $\mathfrak{Y}$-valued positive sesquilinear maps on $\mathfrak{A}\times\mathfrak{A}$ with the following properties:
  • (i)
    $\Lambda_{\Phi}(\mathfrak{A}_{0})=\mathfrak{A}_{0}/\mathfrak{N}_{\Phi}$ is dense in the completion $\widetilde{\mathfrak{A}}$ of $\mathfrak{A}$ w.r. to the norm $\|\cdot\|_{\Phi}$;
  • (ii)
    $\Phi(ac,d)=\Phi(c,a^{*}d),\quad\forall\ a\in\mathfrak{A},\ c,d\in\mathfrak{A}_{0}$ (left-invariant).
If $\mathfrak{A}=\mathfrak{A}_{0}$ and $\omega:\mathfrak{A}\to\mathfrak{Y}$ is a positive linear map, then $\Phi:\mathfrak{A}\times\mathfrak{A}\to\mathfrak{Y}$ given by
$\Phi(a,b)=\omega(b^{*}a),\quad\forall a,b\in\mathfrak{A}$
satisfies the conditions of Definition 3.8, i.e. $\Phi\in{\mathcal{I}}_{{\mathfrak{A}}_{0}}^{\,\mathfrak{Y}}(\mathfrak{A})$.
Proposition 3.9.
Let $(\mathfrak{A},\mathfrak{A}_{0})$ be a quasi *-algebra with unit ${\sf e}$ and $\Phi$ be a $\mathfrak{Y}$-valued left-invariant positive sesquilinear map on $\mathfrak{A}\times\mathfrak{A}$. The following statements are equivalent:
  • (i)
    $\Phi\in{\mathcal{I}}_{{\mathfrak{A}}_{0}}^{\,\mathfrak{Y}}(\mathfrak{A})$;
  • (ii)
    there exist a $B_{\mathfrak{Y}}$-space $\mathcal{K}_{\Phi}$ with $\mathfrak{Y}$-valued inner product $\langle{\cdot}|{\cdot}\rangle_{\mathcal{K}_{\Phi}}$, a dense subspace $\mathcal{D}_{\Phi}\subseteq\mathcal{K}_{\Phi}$ and a closed cyclic *-representation $\pi:\mathfrak{A}\to{\mathcal{L}}^{\dagger}(\mathcal{D}_{\Phi},\mathcal{K}_{\Phi})$ with cyclic vector $\xi_{\Phi}$ such that
    $\langle{\pi(a)\xi}|{\eta}\rangle_{\mathcal{K}_{\Phi}}=\langle{\xi}|{\pi(a^{*})\eta}\rangle_{\mathcal{K}_{\Phi}},\quad\forall\xi,\eta\in\mathcal{D}_{\Phi},a\in\mathfrak{A}$
    and such that
    $\Phi(a,b)=\langle{\pi(a)\xi_{\Phi}}|{\pi(b)\xi_{\Phi}}\rangle_{\mathcal{K}_{\Phi}},\quad\forall a,b\in\mathfrak{A}.$
Proof.
The proof proceeds along the lines of that one of [5, Theorem 3.2], due to the Cauchy-Schwarz inequality for positive $\mathfrak{Y}$-valued sesquilinear maps (Proposition 3.1) and due to Remark 3.6. ∎
Remark 3.10.
By the same arguments as in [5, Corollary 3.5] one can show that every $\mathfrak{Y}$-valued, bounded, left-invariant, positive sesquilinear map on a unital normed quasi *-algebra belongs to ${\mathcal{I}}_{{\mathfrak{A}}_{0}}^{\,\mathfrak{Y}}(\mathfrak{A})$.
In the following example we construct some $\mathfrak{Y}$-valued, bounded, left-invariant, positive sesquilinear maps.
Example 3.11.
Let $\rho$ be a finite trace on a von Neumann algebra ${\mathfrak{M}}$. Let $W\in L^{\infty}(\rho)$, with $W\geq 0$. Let $k\in C([0,\|W\|]\times[0,\|W\|])$ be such that $k\geq 0$. Then, for each $x\in[0,\|W\|]$, the function $\eta_{x}:[0,\|W\|]\to\mathbb{C}$ defined by $\eta_{x}(t)=k(x,t)$ is a continuous, positive function on $[0,\|W\|]$. Therefore, by the functional calculus, $\eta_{x}(W)$ defines a positive operator in $L^{\infty}(\rho)$. Let us define, for every $x\in[0,\|W\|]$ and $X,Y\in L^{2}(\rho)$,
$\varphi(X,Y)(x)=\rho(X\eta_{x}(W)Y^{*}).$
Then, $\varphi(X,Y)\in C([0,\|W\|])$ for all $X,Y\in L^{2}(\rho)$. Moreover, $\varphi:L^{2}(\rho)\times L^{2}(\rho)\to C([0,\|W\|])$ is a bounded, left-invariant, positive sesquilinear map (see [5]).
Let $f:[0,1]\to[0,\|W\|]$ be measurable. For every $X,Y\in L^{2}(\rho)$, let $\psi:L^{2}(\rho)\times L^{2}(\rho)\to L^{2}([0,1])$ defined by $\psi(X,Y)=\phi(X,Y)\circ f$. Then one can check that $\psi$ is a bounded, left-invariant, positive sesquilinear map.
Finally, if $(\mathfrak{A},\mathfrak{A}_{0})$ is a unital quasi *-algebra and $\Psi:\mathfrak{A}\times\mathfrak{A}\to L^{1}(\Omega)$ is a bounded left-invariant positive sesquilinear map, then the induced map $\widetilde{\Psi}:\mathfrak{A}\times\mathfrak{A}\to M(\Omega)$ given by $d\widetilde{\Psi}(a,b)=\Psi(a,b)d\mu$, for all $a,b\in\mathfrak{A}$ and some fixed positive measure $\mu\in M(\Omega)$, is also a bounded left-invariant positive sesquilinear map.
Corollary 3.12.
Let $\mathfrak{A}$ be a *–algebra with unit ${\sf e}$ and let $\omega$ be a positive linear $\mathfrak{Y}$-valued map on $\mathfrak{A}$. Then, there exists a $B_{\mathfrak{Y}}$-space $\mathcal{K}_{\Phi}$ whose norm is induced by a $\mathfrak{Y}$-valued inner product $\langle{\cdot}|{\cdot}\rangle_{\mathcal{K}_{\Phi}}$, a dense subspace $\mathcal{D}_{\omega}\subseteq\mathcal{K}_{\Phi}$ and a closed cyclic *–representation $\Pi_{\omega}$ of $\mathfrak{A}$ with domain $\mathcal{D}_{\omega}$, such that
$\omega(b^{*}ac)=\langle{\Pi_{\omega}(a)\Lambda_{\omega}(c)}|{\Lambda_{\omega}(b)}\rangle_{\mathcal{K}_{\Phi}},\quad\forall a,b,c\in\mathfrak{A}.$
Moreover, there exists a cyclic vector $\eta_{\omega}$, such that
$\omega(a)=\langle{{\Pi}_{\omega}(a)\eta_{\omega}}|{\eta_{\omega}}\rangle_{\mathcal{K}_{\Phi}},\quad\forall\ a\in\mathfrak{A}.$
The representation is unique up to unitary equivalence.
Proof.
The statement can be proved analogously to [5, Corollary 3.10], due to Proposition 3.9. ∎
4. Representations induced by completely positive sesquilinear maps
In this section, we extend the results of Section 3 to the case of $\mathfrak{Y}$-valued completely positive sesquilinear maps.
Let $\mathfrak{X}$ and $\mathfrak{Y}$ be vector spaces. We will denote by $\mathcal{S}_{\mathfrak{Y}}(\mathfrak{X})$ the space of all $\mathfrak{Y}$-valued sesquilinear maps on $\mathfrak{X}$.
Definition 4.1.
Let $\mathfrak{X}$ be a normed vector space, $\mathfrak{Y}$ an ordered Banach module over the *-algebra $\mathfrak{Y}_{0}$, with positive cone $\mathfrak{K}$ and $(\mathfrak{A},\mathfrak{A}_{0})$ a normed quasi *-algebra with unit. The sesquilinear map $\Phi:\mathfrak{A}\times\mathfrak{A}\to\mathcal{S}_{\mathfrak{Y}}(\mathfrak{X})$ is called
  • bounded if there exists a constant $M>0$ such that
    $\|\Phi(a,b)(x_{1},x_{2})\|_{\mathfrak{Y}}\leq M\|a\|\|b\|\|x_{1}\|\|x_{2}\|,\quad\forall a,b\in\mathfrak{A},x_{1},x_{2}\in\mathfrak{X};$
  • left-invariant if $\Phi(ac,d)=\Phi(c,a^{*}d),\quad\forall a\in\mathfrak{A},c,d\in\mathfrak{A}_{0}$;
  • completely positive if for every $N\in{\mathbb{N}}$, $a_{1},\dots,a_{N}\in\mathfrak{A}$,
    $x_{1},\dots,x_{N}\in{\mathfrak{X}}$,
    $\sum_{i,j=1}^{N}\Phi(a_{i},a_{j})(x_{i},x_{j})\in\mathfrak{K}.$
The Stinespring Theorem is the main result on completely positive linear maps. In the following we consider certain completely positive sesquilinear maps taking values on the space of all $\mathfrak{Y}$-valued sesquilinear maps on a vector space, see Theorem 4.2 below.
Let $(\mathfrak{A},\mathfrak{A}_{0})$ be a normed quasi *-algebra with unit ${\sf e}$, $\mathfrak{X}$ be a normed complex vector space. Let the cone $\mathfrak{K}$ in $\mathfrak{Y}$ be closed and $\Phi:\mathfrak{A}\times\mathfrak{A}\to\mathcal{S}_{\mathfrak{Y}}({\mathfrak{X}})$ be a left-invariant positive sesquilinear map. Consider the algebraic tensor product $\mathfrak{A}\otimes{\mathfrak{X}}$ and its subset
$\mathcal{N}_{\Phi}=\left\{\sum_{i=1}^{n}a_{i}\otimes x_{i}\in\mathfrak{A}\otimes\mathfrak{X}|\,\sum_{i=1}^{n}\sum_{j=1}^{n}\Phi(a_{i},a_{j})(x_{i},x_{j})=0_{\mathfrak{Y}}\right\}.$
Hence, it is easy to check that the quotient space $(\mathfrak{A}\otimes{\mathfrak{X}})/\mathcal{N}_{\Phi}$ is a normed space.
Theorem 4.2.
Let $\mathfrak{X}$ be a normed complex vector space. Let $(\mathfrak{A},\mathfrak{A}_{0})$ be a normed quasi *-algebra with unit ${\sf e}$ and $\Phi:\mathfrak{A}\times\mathfrak{A}\to\mathcal{S}_{\mathfrak{Y}}({\mathfrak{X}})$ a left-invariant sesquilinear map. Let $\Phi$ be bounded with bound $M>0$. Then, $\Phi$ is completely positive if and only if there exist a $B_{\mathfrak{Y}}$-space $\mathcal{K}_{\Phi}$, a dense subspace $\mathcal{D}_{\Phi}$ of $\mathcal{K}_{\Phi}$, a closed *-representation $\pi$ of $\mathfrak{A}$ in ${\mathcal{L}}^{\mbox{\scriptsize${\dagger}$}}(\mathcal{D}_{\Phi},\mathcal{K}_{\Phi})$ and a bounded linear operator $V:\mathfrak{X}\to\mathcal{D}_{\Phi}$ such that $\pi(\mathfrak{A})V{\mathfrak{X}}=(\mathfrak{A}\otimes{\mathfrak{X}})/\mathcal{N}_{\Phi}$ and, for all $a,b\in\mathfrak{A}$ and $x_{1},x_{2}\in\mathfrak{X}$, it holds that
$\Phi(a,b)(x_{1},x_{2})=\langle{\pi(a)Vx_{1}}|{\pi(b)Vx_{2}}\rangle_{\mathcal{K}_{\Phi}}.$
In this case $\|V\|^{2}\leq M\|{\sf e}\|^{2}$. Moreover, the triple $(\pi,V,\mathcal{D}_{\Phi})$ is such that $\overline{\pi(\mathfrak{A})V{\mathfrak{X}}}=\mathcal{K}_{\Phi}$.
Proof.
The proof goes along the lines of that of [4] where more traditional sesquilinear forms were considered thanks to Proposition 3.1 and Remark 3.6. We omit the details. ∎
By Theorem 4.2, and similar arguments to those in the proof of Proposition 3.3, we deduce the following
Corollary 4.3.
Let $\mathfrak{X}$ be a normed complex vector space. Let $(\mathfrak{A},\mathfrak{A}_{0})$ be a normed quasi *-algebra with unit ${\sf e}$ and $\{\Phi_{n}\}_{n}$ be a sequence of bounded, left-invariant, completely positive sesquilinear maps $\Phi_{n}:\mathfrak{A}\times\mathfrak{A}\to\mathcal{S}_{\mathfrak{Y}}({\mathfrak{X}})$. Let $\{a_{n}\}_{n}$, $\{\widetilde{a}_{n}\}_{n}$ be sequences in $\mathfrak{A}$ and $\{x_{n}\}_{n}$, $\{\widetilde{x}_{n}\}_{n}$ be sequences in ${\mathfrak{X}}$ such that the series $\sum_{n=1}^{\infty}\Phi_{n}(a_{n},a_{n})(x_{n},x_{n})$ and $\sum_{n=1}^{\infty}\Phi_{n}(\widetilde{a}_{n},\widetilde{a}_{n})(\widetilde{x}_{n},\widetilde{x}_{n})$ are convergent in $\mathfrak{Y}$. Then the series $\sum_{n=1}^{\infty}\Phi_{n}(a_{n},\widetilde{a}_{n})(x_{n},\widetilde{x}_{n})$ is convergent and
(4.1) $\hskip 8.53581pt\left\|\sum_{n=1}^{\infty}\Phi_{n}(a_{n},\widetilde{a}_{n})(x_{n},\widetilde{x}_{n})\right\|_{\mathfrak{Y}}$
$\quad\quad\leq\left\|\sum_{n=1}^{\infty}\Phi_{n}(a_{n},a_{n})(x_{n},x_{n})\right\|_{\mathfrak{Y}}^{1/2}\left\|\sum_{n=1}^{\infty}\Phi_{n}(\widetilde{a}_{n},\widetilde{a}_{n})(\widetilde{x}_{n},\widetilde{x}_{n})\right\|_{\mathfrak{Y}}^{1/2}.$
In the following we will refer to a triple $(\pi,V,\mathcal{D}_{\Phi})$ as in Theorem 4.2 as a Stinespring triple decomposing $\Phi$.
Definition 4.4.
Let ${\mathfrak{X}}_{1},{\mathfrak{X}}_{2}$ be two $B_{\mathfrak{Y}}$-spaces with respect to the norms $\|\cdot\|_{\Phi}$ and $\|\cdot\|_{\Psi}$, respectively, induced by two $\mathfrak{Y}$-valued inner products $\Phi:{\mathfrak{X}}_{1}\times{\mathfrak{X}}_{1}\to\mathfrak{Y}$ and $\Psi:{\mathfrak{X}}_{2}\times{\mathfrak{X}}_{2}\to\mathfrak{Y}$. A surjective operator $U:{\mathfrak{X}}_{1}\to{\mathfrak{X}}_{2}$ is said unitary if
$\langle{U\xi}|{U\eta}\rangle_{\Psi}=\langle{\xi}|{\eta}\rangle_{\Phi},\quad\forall\xi,\eta\in{\mathfrak{X}}_{1}.$
In analogy to Corollary 3.10 in [4], we can prove the following
Proposition 4.5.
Let $\mathfrak{X}$ be a normed complex vector space. Let $(\mathfrak{A},\mathfrak{A}_{0})$ be a normed quasi *-algebra with unit ${\sf e}$ and $\Phi:\mathfrak{A}\times\mathfrak{A}\to\mathcal{S}_{\mathfrak{Y}}({\mathfrak{X}})$ be a bounded left-invariant sesquilinear map. If $\Phi$ is completely positive then $\pi$ and $V$ in the triple $(\pi,V,\mathcal{D}_{\Phi})$ are uniquely determined by $\Phi$ up to unitary equivalence, i.e., if $(\pi,V,\mathcal{D}_{\Phi})$ and $(\pi_{1},V_{1},\mathcal{E}_{\Phi})$ are two Stinespring triples decomposing $\Phi$, there exists a unitary operator $U$ such that $UV=V_{1}$, $U\mathcal{D}_{\Phi}=\mathcal{E}_{\Phi}$ and $\pi(a)=U^{-1}\pi_{1}(a)U$, for all $a\in\mathfrak{A}$.
Remark 4.6.
By Proposition 4.5 all the Stinespring triples $(\pi,V,\mathcal{D}_{\Phi})$ decomposing $\Phi$ are such that $\overline{\pi(\mathfrak{A})V{\mathfrak{X}}}=\mathcal{K}_{\Phi}$.
Following [24, Section 3.5] we now provide a generalization of the Radon-Nikodym theorem for completely positive sesquilinear maps with values in the space $\mathcal{S}_{\mathfrak{Y}}({\mathfrak{X}})$ of $\mathfrak{Y}$-valued sesquilinear maps on ${\mathfrak{X}}\times{\mathfrak{X}}$.
Proposition 4.7.
Let $\mathfrak{X}$ be a normed complex vector space, $(\mathfrak{A},\mathfrak{A}_{0})$ be a normed quasi *-algebra with unit ${\sf e}$ and $\Psi,\Phi:\mathfrak{A}\times\mathfrak{A}\to\mathcal{S}_{\mathfrak{Y}}({\mathfrak{X}})$ be bounded, left-invariant, completely positive sesquilinear maps. Suppose that there exists $\gamma>0$ such that, for every $N\in\mathbb{N}$, for all $a_{1},\dots,a_{N}\in\mathfrak{A},x_{1},\dots,x_{N}\in{\mathfrak{X}}$
$\sum_{i,j=1}^{N}\Psi(a_{i},a_{j})(x_{i},x_{j})\leq\gamma\sum_{i,j=1}^{N}\Phi(a_{i},a_{j})(x_{i},x_{j}).$
Let $(\pi_{\Psi},V_{\Psi},\mathcal{D}_{\Psi})$ and $(\pi_{\Phi},V_{\Phi},\mathcal{D}_{\Phi})$ be two Stinespring triples decomposing $\Psi$ and $\Phi$, respectively. Then, there exists a linear operator $T:\pi_{\Phi}(\mathfrak{A})V_{\Phi}{\mathfrak{X}}\to\mathcal{K}_{\Psi}$ such that:
  • (i)
    $TV_{\Phi}=V_{\Psi}$;
  • (ii)
    $T\pi_{\Phi}(a)=\pi_{\Psi}(a)\,{\scriptstyle\square}\,T$ on $\pi_{\Phi}(\mathfrak{A}_{0})V_{\Phi}{\mathfrak{X}}$, for every $a\in\mathfrak{A}$,
and for all $a,b\in\mathfrak{A}$, $x_{1},x_{2}\in{\mathfrak{X}}$,
$\Psi(a,b)(x_{1},x_{2})=\langle{T\pi_{\Phi}(a)V_{\Phi}x_{1}}|{T\pi_{\Phi}(b)V_{\Phi}x_{2}}\rangle_{\mathcal{K}_{\Psi}}.$
Moreover, if the norm $\|\cdot\|_{\mathfrak{Y}}$ in $\mathfrak{Y}$ is order-preserving, then $T$ is bounded, defined on $\mathcal{K}_{\Phi}=\overline{\pi_{\Phi}(\mathfrak{A})V_{\Phi}{\mathfrak{X}}}$ and $\|T\|\leq\sqrt{\gamma}$. In this case, for every $a\in\mathfrak{A}$, $T\pi_{\Phi}(a)=\pi_{\Psi}(a)\,{\scriptstyle\square}\,T$ on the whole $\mathcal{K}_{\Phi}$.
Proof.
Let $N\in\mathbb{N}$ and $a_{1},\dots,a_{N}\in\mathfrak{A}$ and $x_{1},\dots,x_{N}\in{\mathfrak{X}}$, then
(4.2) $\sum_{i=1}^{N}\pi_{\Psi}(a_{i})V_{\Psi}x_{i}\Bigg{|}\sum_{j=1}^{N}\pi_{\Psi}(a_{j})V_{\Psi}x_{j}\right>_{\mathcal{K}_{\Psi}}=\sum_{i,j=1}^{N}\Psi(a_{i},a_{j})(x_{i},x_{j})$
$\hskip 28.45274pt\leq\gamma\sum_{i,j=1}^{N}\Phi(a_{i},a_{j})(x_{i},x_{j})$
$\hskip 28.45274pt=\gamma\left<\sum_{i=1}^{N}\pi_{\Phi}(a_{i})V_{\Phi}x_{i}\Bigg{|}\sum_{j=1}^{N}\pi_{\Phi}(a_{j})V_{\Phi}x_{j}\right>_{\mathcal{K}_{\Phi}}.$
If $\sum_{i=1}^{N}\pi_{\Phi}(a_{i})V_{\Phi}x_{i}=0$, then also $\sum_{i=1}^{N}\pi_{\Psi}(a_{i})V_{\Psi}x_{i}=0$, hence, we can define a linear operator $T$ by
(4.3) $T\sum_{i=1}^{N}\pi_{\Phi}(a_{i})V_{\Phi}x_{i}=\sum_{i=1}^{N}\pi_{\Psi}(a_{i})V_{\Psi}x_{i},$
for all $a_{i}\in\mathfrak{A},\,x_{i}\in{\mathfrak{X}},i\in\{1,\dots,N\}$. Moreover, by taking $a={\sf e}\in\mathfrak{A}_{0}$ then $TV_{\Phi}=V_{\Psi}$, since $\pi_{\Phi}({\sf e})={\mathbb{I}}_{\mathcal{D}_{\Phi}}$ and $\pi_{\Psi}({\sf e})={\mathbb{I}}_{\mathcal{D}_{\Psi}}$.
Then $T\,{\scriptstyle\square}\,\pi_{\Phi}(a)=\pi_{\Psi}(a)\,{\scriptstyle\square}\,T$, for every $a\in\mathfrak{A}$. Indeed, if now $a\in\mathfrak{A}$, $c\in\mathfrak{A}_{0}$, $x\in{\mathfrak{X}}$, by (4.3)
$\begin{split}T(\pi_{\Phi}(a)\,{\scriptstyle\square}\,\pi_{\Phi}(c))V_{\Phi}x&=T({\pi_{\Phi}(a)^{\dagger}}^{*}\pi_{\Phi}(c))V_{\Phi}x=(T{\pi_{\Phi}(a)^{\dagger}}^{*})\pi_{\Phi}(c)V_{\Phi}x,\end{split}$
on the other hand,
$\begin{split}T(\pi_{\Phi}(a)\,{\scriptstyle\square}\,\pi_{\Phi}(c))V_{\Phi}x&=T\pi_{\Phi}(ac)V_{\Phi}x=\pi_{\Psi}(ac)V_{\Psi}x\\ &={\pi_{\Psi}(a)^{\dagger}}^{*}\pi_{\Psi}(c)V_{\Psi}x={\pi_{\Psi}(a)^{\dagger}}^{*}(T\pi_{\Phi}(c)V_{\Phi}x)\\ &=({\pi_{\Psi}(a)^{\dagger}}^{*}T)\pi_{\Phi}(c)V_{\Phi}x=(\pi_{\Psi}(a)\,{\scriptstyle\square}\,T)\pi_{\Phi}(c)V_{\Phi}x.\end{split}$
Finally, if $y_{1}\leq y_{2}$, with $y_{1},y_{2}\in\mathfrak{K}$ implies that also $\|y_{1}\|_{\mathfrak{Y}}\leq\|y_{2}\|_{\mathfrak{Y}}$ then, by (4.2), we have that $T$ is bounded, hence it extends to the closure $\mathcal{K}_{\Phi}=\overline{\pi_{\Phi}(\mathfrak{A})V_{\Phi}{\mathfrak{X}}}$ and $\|T\|\leq\sqrt{\gamma}$. Since the set $\pi_{\Phi}(\mathfrak{A}_{0})V_{\Phi}{\mathfrak{X}}=(\mathfrak{A}_{0}\otimes{\mathfrak{X}})/\mathcal{N}_{\Phi}$ is dense in $\mathcal{K}_{\Phi}$, we conclude that $T\pi_{\Phi}(a)=\pi_{\Psi}(a)\,{\scriptstyle\square}\,T$ on $\mathcal{K}_{\Phi}$, for every $a\in\mathfrak{A}$. The last statement follows by the very construction of $T$.∎
A Radon Nikodym-like theorem holds for positive sesquilinear maps $\Phi,\Psi$ and not just for completely positive sesquilinear maps.
Proposition 4.8.
Let $\mathfrak{X}$ be a normed vector space and let the norm $\|\cdot\|_{\mathfrak{Y}}$ preserve the order in $\mathfrak{Y}$. Let $(\mathfrak{A},\mathfrak{A}_{0})$ be a normed quasi *-algebra with unit ${\sf e}$ and $\Phi,\Psi:\mathfrak{A}\times\mathfrak{A}\to\mathcal{S}_{\mathfrak{Y}}({\mathfrak{X}})$ be bounded left-invariant positive sesquilinear maps such that
$\Psi(a,a)(x,x)\leq\gamma\,\Phi(a,a)(x,x),\quad\forall a\in\mathfrak{A},x\in{\mathfrak{X}}$
for some $\gamma>0$. Let $\mathcal{K}_{\Phi}$ and $\mathcal{K}_{\Psi}$ be $B_{\mathfrak{Y}}$-spaces, $\mathcal{D}_{\Phi}$ and $\mathcal{D}_{\Psi}$ dense subspaces of $\mathcal{K}_{\Phi}$ and $\mathcal{K}_{\Psi}$, respectively, $\pi_{\Phi}:\mathfrak{A}\to{\mathcal{L}}^{\dagger}(\mathcal{D}_{\Phi},\mathcal{K}_{\Phi})$ and $\pi_{\Psi}:\mathfrak{A}\to{\mathcal{L}}^{\dagger}(\mathcal{D}_{\Psi},\mathcal{K}_{\Psi})$ closed cyclic *-representations of $\Phi$ and $\Psi$, respectively, with cyclic vectors $\xi_{\Phi}$ and $\xi_{\Psi}$, respectively, as in Proposition 3.9. Then there exists a unique operator $T:\mathcal{K}_{\Phi}\to\mathcal{K}_{\Psi}$, with $\|T\|\leq\sqrt{\gamma}$ such that for all $a,b\in\mathfrak{A}$,
$\Psi(a,b)=\langle{T\pi_{\Phi}(a)\xi_{\Phi}}|{T\pi_{\Phi}(b)\xi_{\Phi}}\rangle_{\mathcal{K}_{\Psi}}.$
Moreover, $T\pi_{\Phi}(a)=\pi_{\Psi}(a)\,{\scriptstyle\square}\,T$ on $\pi_{\Phi}({\mathfrak{A}}_{0})\xi_{\Phi}$, for all $a\in\mathfrak{A}$.
Proof.
The proof goes along the lines of that of Proposition 4.7, by applying the representations of $\Phi$ and $\Psi$ of Proposition 3.9 and by letting $T\pi_{\Phi}(a)\xi_{\Phi}=\pi_{\Psi}(a)\xi_{\Psi}$ for all $a\in\mathfrak{A}$, see also Remark 3.10. ∎
We give now some examples.
Example 4.9.
Let $\mathfrak{Y}$ be either a $C^{*}$-algebra or an ordered Banach module over a *-algebra, satisfying $(D1)-(D2)$ and let $\widetilde{\Phi}:\mathfrak{A}\times\mathfrak{A}\to\mathfrak{Y}$ be a bounded left-invariant positive sesquilinear map. Define the map $\Phi:\mathfrak{A}\times\mathfrak{A}\to\mathcal{S}_{\mathfrak{Y}}(\mathfrak{A}_{0})$ by
$\Phi(a,b)(c,d)=\widetilde{\Phi}(ac,bd),\quad\forall a,b\in\mathfrak{A},c,d\in\mathfrak{A}_{0}.$
Then $\Phi$ is bounded, left-invariant, completely positive and sesquilinear.
Example 4.10.
Let ${\mathfrak{M}}$ be a von Neumann algebra with a finite trace $\rho$ and let $W$ and $\eta_{x}(W)$ be as in Example 3.11. Consider a separable Hilbert space $\mathcal{H}$ and the space $L^{2}([0,\|W\|],{\mathfrak{B}}(\mathcal{H}))$ w.r. to the Gel’fand-Pettis integral (see [13] and [5, Example 2.26]). Fix $F\in L^{2}([0,\|W\|],{\mathfrak{B}}(\mathcal{H}))$ with $F(t)\geq 0$ for a.e. $t\in[0,\|W\|]$ and $T\in{\mathfrak{B}}_{2}(\mathcal{H})$ and consider $\Phi:L^{2}(\rho)\times L^{2}(\rho)\to\mathcal{S}_{{\mathfrak{B}}_{1}(\mathcal{H})}(L^{\infty}(\rho))$ given by
$\Phi(A,B)(X_{1},X_{2})=T^{*}\left(\int_{0}^{\|W\|}\rho(X_{2}^{*}B^{*}AX_{1}\eta_{x}(W)))F(x)dx\right)T,$
for every $A,B\in L^{2}(\rho)$, $X_{1},X_{2}\in L^{2}(\rho).$ Then $\Phi$ is a bounded, left-invariant, completely positive sesquilinear map.
4.1. Applications to operator-valued maps
In the following we will denote by ${\mathfrak{B}}({\mathfrak{X}},\mathfrak{Y})$ the space of bounded linear operators from ${\mathfrak{X}}$ into $\mathfrak{Y}$, where ${\mathfrak{X}}$ is a general Banach space.
Definition 4.11.
Let ${\mathfrak{X}}$ be a Banach space, $\Gamma:{\mathfrak{X}}\times{\mathfrak{X}}\to{\mathfrak{X}}$ a bounded sesquilinear map. Let $(\mathfrak{A},\mathfrak{A}_{0})$ be a quasi *-algebra with unit ${\sf e}$. A sesquilinear map $\Phi:\mathfrak{A}\times\mathfrak{A}\to{\mathfrak{B}}({\mathfrak{X}},\mathfrak{Y})$ will be called completely positive w.r. to $\Gamma$ if for each $N\in\mathbb{N}$ and for every $a_{1},\dots,a_{N}\in\mathfrak{A}$, $x_{1},\dots,x_{N}\in{\mathfrak{X}}$ it is
$\sum_{i,j=1}^{N}\Phi(a_{i},a_{j})(\Gamma(x_{i},x_{j}))\in\mathfrak{K}.$
Proposition 4.12.
Let ${\mathfrak{X}}$ be a Banach space, $\Gamma:{\mathfrak{X}}\times{\mathfrak{X}}\to{\mathfrak{X}}$ a bounded sesquilinear map. Let $(\mathfrak{A},\mathfrak{A}_{0})$ be a normed quasi *-algebra with unit ${\sf e}$ and $\Phi:\mathfrak{A}\times\mathfrak{A}\to{\mathfrak{B}}({\mathfrak{X}},\mathfrak{Y})$ be a bounded, left-invariant, completely positive sesquilinear map. Then, there exist a dense subspace $\mathcal{D}_{\Phi}$ of a $B_{\mathfrak{Y}}$-space $\mathcal{K}_{\Phi}$, a closed *-representation $\pi$ of $\mathfrak{A}$ in ${\mathcal{L}}^{\mbox{\scriptsize${\dagger}$}}(\mathcal{D}_{\Phi},\mathcal{K}_{\Phi})$ and a bounded linear operator $V:{\mathfrak{X}}\to\mathcal{D}_{\Phi}$ with $\|V\|^{2}\leq\|\Phi({\sf e},{\sf e})\|\|\Gamma\|$ such that we have
$\Phi(a,b)(\Gamma(x_{1},x_{2}))=\langle{\pi(a)Vx_{1}}|{\pi(b)Vx_{2}}\rangle_{\mathcal{K}_{\Phi}},\quad\forall a,b\in\mathfrak{A},\,\,x_{1},x_{2}\in{\mathfrak{X}}.$
Moreover, for all $a,b\in\mathfrak{A}$ and $x_{1},x_{2}\in{\mathfrak{X}}$, we have
(4.4) $\|\Phi(a,b)(\Gamma(x_{1},x_{2}))\|_{\mathfrak{Y}}\\ \leq\|\Phi(a,a)(\Gamma(x_{1},x_{1}))\|_{\mathfrak{Y}}^{\frac{1}{2}}\|\Phi(b,b)(\Gamma(x_{2},x_{2}))\|_{\mathfrak{Y}}^{\frac{1}{2}}.$
If $B_{1}$ denotes the unit ball in ${\mathfrak{X}}$ and $\Gamma(B_{1}\times B_{1})=B_{1}$, then for all $a,b\in\mathfrak{A}$ we have
(4.5) $\|\Phi(a,b)\|\leq\|\Phi(a,a)\|^{1/2}\|\Phi(b,b)\|^{1/2}.$
Proof.
Let $\mathcal{S}_{\mathfrak{Y}}({\mathfrak{X}})$ be the space of all $\mathfrak{Y}$-valued sesquilinear maps on ${\mathfrak{X}}$ and $\widetilde{\Phi}:\mathfrak{A}\times\mathfrak{A}\to\mathcal{S}_{\mathfrak{Y}}({\mathfrak{X}})$ be given by:
$\widetilde{\Phi}(a,b)(x_{1},x_{2}):=\Phi(a,b)(\Gamma(x_{1},x_{2})),\quad\forall a,b\in\mathfrak{A},\,\,x_{1},x_{2}\in{\mathfrak{X}}.$
Then, $\widetilde{\Phi}$ is a left-invariant, completely positive sesquilinear map on $\mathfrak{Y}$ since it inherits these properties from $\Phi$. Moreover $\widetilde{\Phi}$ is bounded. By Theorem 4.2, the statement follows. Since $\pi$ is a *-representation, $\pi({\sf e})={\mathbb{I}}_{\mathcal{K}_{\Phi}}$, the identity operator on the space $\mathcal{K}_{\Phi}$, hence, for every $x\in{\mathfrak{X}}$, we have that
$\begin{split}\|\langle{Vx}|{Vx}\rangle_{\mathcal{K}_{\Phi}}\|_{\mathfrak{Y}}&=\|\langle{\pi({\sf e})Vx}|{\pi({\sf e})Vx}\rangle_{\mathcal{K}_{\Phi}}\|_{\mathfrak{Y}}=\|\langle{{\sf e}\otimes x}|{{\sf e}\otimes x}\rangle_{\mathcal{K}_{\Phi}}\|_{\mathfrak{Y}}\\ &=\|(\Phi({\sf e},{\sf e}))(\Gamma(x,x))\|_{\mathfrak{Y}}\leq\|\Phi({\sf e},{\sf e})\|\|\Gamma(x,x)\|\\ &\leq\|\Phi({\sf e},{\sf e})\|\|\Gamma\|\|x\|^{2}.\end{split}$
The inequality (4.4) follows from Proposition 3.1 applied to the obtained representation, whereas (4.5) follows by taking supremums over $B_{1}$ in (4.4) and using that $\Gamma(B_{1}\times B_{1})=B_{1}$. ∎
The following example shows maps $\Phi$ and $\Gamma$ satisfying the hypotheses of Proposition 4.12.
Example 4.13.
Take ${\mathfrak{X}}={\mathfrak{M}}$ a von Neumann algebra with a finite trace $\rho$ and $\Gamma(T_{1},T_{2})=T_{1}T_{2}^{*}$, for all $T_{1},T_{2}\in{\mathfrak{M}}$. Since ${\mathfrak{M}}$ is unital, it follows that $\Gamma(B_{1}\times B_{1})=B_{1}$. Let $W$ and $\eta_{x}(W)$ be as in Example 3.11. For every $T\in{\mathfrak{M}}$ and $x\in[0,\|W\|]$, let
$(\varphi(A,B)(T))(x)=\rho(A\,\eta_{x}(W)\,T\,\eta_{x}(W)\,B^{*}),\quad A,B\in L^{2}(\rho).$
By some calculations it is not hard to see that $\varphi(A,B)(T)(\cdot)$ is a continuous function on $[0,\|W\|]$ and that $\varphi$ is a bounded, left-invariant, sesquilinear map from $L^{2}(\rho)\times L^{2}(\rho)$ into ${\mathfrak{B}}({\mathfrak{M}},C([0,\|W\|]))$, which is completely positive w.r. to $\Gamma$. Moreover, if $\widetilde{{\mathfrak{M}}}$ is another von Neumann algebra with a semi-finite trace $\widetilde{\rho}$ and if we choose some $\widetilde{W}\in\widetilde{{\mathfrak{M}}}$, with $\widetilde{W}\geq 0$ and $\|\widetilde{W}\|=\|W\|$, then by the functional calculus, $\varphi(A,B)(T)(\widetilde{W})\in\widetilde{{\mathfrak{M}}}$, for all $A,B\in L^{2}(\rho)$ and $T\in{\mathfrak{M}}$. Hence, the map $\Phi:L^{2}(\rho)\times L^{2}(\rho)\to{\mathfrak{B}}({\mathfrak{M}},\widetilde{{\mathfrak{M}}}^{*})$ given by
$(\Phi(A,B)(T))(S)=\tilde{\rho}(S(\varphi(A,B)(T))(\widetilde{W}))$
for all $T\in{\mathfrak{M}}$, $A,B\in L^{2}(\rho)$ and $S\in\widetilde{{\mathfrak{M}}}$ is another example of a bounded, left-invariant sesquilinear map which is completely positive w.r. to $\Gamma$. Further, if $G\in L^{1}(\widetilde{\rho})$, then the map $\widetilde{\Phi}:L^{2}(\rho)\times L^{2}(\rho)\to{\mathfrak{B}}({\mathfrak{M}},L^{1}(\widetilde{\rho}))$ given by
$\widetilde{\Phi}(A,B)(T)=G(\varphi(A,B)(T)(\widetilde{W}))G^{*}$
is also an example of bounded left-invariant sesquilinear map which is completely positive w.r. to $\Gamma$.
Motivated by the notion of the classical complete positivity, we introduce now the following generalization.
Definition 4.14.
Let ${\mathfrak{X}},{\mathfrak{Z}}$ be Banach spaces, $\Psi:{\mathfrak{Z}}\times{\mathfrak{X}}\to\mathfrak{Y}$ and $\Phi:\mathfrak{A}\times\mathfrak{A}\to{\mathfrak{B}}({\mathfrak{X}},{\mathfrak{Z}})$ be sesquilinear maps. The map $\Phi$ is said completely positive w.r. to $\Psi$ if for every $N\in\mathbb{N}$ and all $a_{1},\dots,a_{N}\in\mathfrak{A}$, $x_{1},x_{2},\dots,x_{N}\in{\mathfrak{X}}$ it holds that
$\sum_{i,j=1}^{N}\Psi(\Phi(a_{i},a_{j})x_{i},x_{j})\in\mathfrak{K}.$
Remark 4.15.
If ${\mathfrak{X}}={\mathfrak{Z}}=\mathcal{H}$ is a Hilbert space, $\mathfrak{Y}=\mathbb{C}$, $\mathfrak{A}$ is a unital $C^{*}$-algebra, $\Psi$ is inner product on $\mathcal{H}$, $\phi:\mathfrak{A}\to{\mathfrak{B}}(\mathcal{H})$ is a linear map and $\Phi(a,b)=\phi(b^{*}a)$, for all $a,b\in\mathfrak{A}$, then Definition 4.14 reduces to the definition of the classical complete positivity of $\phi$ and $\Phi$.
Proposition 4.16.
Let ${\mathfrak{X}},{\mathfrak{Z}}$ be Banach spaces, $\Psi:{\mathfrak{Z}}\times{\mathfrak{X}}\to\mathfrak{Y}$ and $\Phi:\mathfrak{A}\times\mathfrak{A}\to{\mathfrak{B}}({\mathfrak{X}},{\mathfrak{Z}})$ be sesquilinear maps. If $\Phi$ and $\Psi$ are bounded and $\Phi$ is left-invariant and completely positive w.r. to $\Psi$, then there exists a dense subspace $\mathcal{D}_{\Phi}$ of a $B_{\mathfrak{Y}}$-space $\mathcal{K}_{\Phi}$, a closed *-representation $\pi$ of $\mathfrak{A}$ in ${\mathcal{L}}^{\mbox{\scriptsize${\dagger}$}}(\mathcal{D}_{\Phi},\mathcal{K}_{\Phi})$ and a bounded linear operator $V:{\mathfrak{X}}\to\mathcal{D}_{\Phi}$ with $\|V\|^{2}\leq\|\Phi({\sf e},{\sf e})\|\|\Psi\|$ such that we have
$\Psi(\Phi(a,b)x_{1},x_{2})=\langle{\pi(a)Vx_{1}}|{\pi(b)Vx_{2}}\rangle_{\mathcal{K}_{\Phi}},\quad\forall a,b\in\mathfrak{A},\,\,x_{1},x_{2}\in{\mathfrak{X}}.$
Moreover,
$\|\Psi(\Phi(a,b)x_{1},x_{2})\|_{\mathfrak{Y}}\leq\|\Psi(\Phi(a,a)x_{1},x_{2})\|_{\mathfrak{Y}}^{1/2}\|\Psi(\Phi(b,b)x_{1},x_{2})\|_{\mathfrak{Y}}^{1/2},$
for all $a,b\in\mathfrak{A},\,\,x_{1},x_{2}\in{\mathfrak{X}}$.
If $\|\Psi\|=1$ and for all $z\in{\mathfrak{Z}}$ we have that $\|z\|_{{\mathfrak{Z}}}=\sup_{\begin{subarray}{c}x\in{\mathfrak{X}}\\ \|x\|_{{\mathfrak{X}}}\leq 1\end{subarray}}\|\Psi(z,x)\|_{\mathfrak{Y}}$, then for all $a,b\in\mathfrak{A}$, we have
$\|\Phi(a,b)\|\leq\|\Phi(a,a)\|^{1/2}\|\Phi(b,b)\|^{1/2}.$
Proof.
As in the proof of Proposition 4.12, we can consider the induced sesquilinear map $\widetilde{\Phi}:\mathfrak{A}\times\mathfrak{A}\to\mathcal{S}_{\mathfrak{Y}}({\mathfrak{X}})$ given by
$\widetilde{\Phi}(a,b)(x_{1},x_{2})=\Psi(\Phi(a,b)x_{1},x_{2}),\quad\forall x_{1},x_{2}\in{\mathfrak{X}}$
and deduce the first two statements by applying Theorem 4.2 and the obtained *-representation. The last inequality in the statements follows from the first one, by taking the supremums over the unit ball in ${\mathfrak{Z}}$. ∎
Example 4.17.
Let ${\mathfrak{M}}$ be a von Neumann algebra with a finite trace $\rho$ and let $W\in L^{\infty}(\rho)$. Put ${\mathfrak{Z}}=\mathfrak{Y}=L^{1}(\rho)$, ${\mathfrak{X}}=L^{\infty}(\rho)$. If $\Psi:{\mathfrak{Z}}\times{\mathfrak{X}}\to\mathfrak{Y}$ is given by $\Psi(Z,X)=X^{*}Z$, then $\Psi$ is bounded, $\|\Psi\|=1$ and $\|Z\|_{{\mathfrak{Z}}}=\sup_{\begin{subarray}{c}X\in{\mathfrak{X}}\\ \|X\|_{{\mathfrak{X}}}\leq 1\end{subarray}}\|\Psi(Z,X)\|_{\mathfrak{Y}}$, due to the fact that $L^{\infty}(\rho)\subset L^{1}(\rho)$ and that $L^{\infty}(\rho)={\mathfrak{M}}$ is unital. Let $\Phi:L^{2}(\rho)\times L^{2}(\rho)\to{\mathfrak{B}}({\mathfrak{X}},{\mathfrak{Z}})$ be given by
$\Phi(A,B)(X)=W^{*}B^{*}AWX,\quad\forall A,B\in L^{2}(\rho),\,X\in{\mathfrak{X}}=L^{\infty}(\rho).$
Then $\Phi$ is a bounded, left-invariant sesquilinear map which is completely positive w.r. to $\Psi$.
Acknowledgements: GB and CT acknowledge that this work has been carried out within the activities of Gruppo UMI Teoria dell’Appros-simazione e Applicazioni and of GNAMPA of the INdAM. SI is supported by the Ministry of Science, Technological Development and Innovations, Republic of Serbia, grant no. 451-03-66/2024-03/200029.